### v2

# Stabilizer Entanglement as a Magic Highway

Abstract
Non-stabilizerness is a key resource for fault-tolerant quantum computation, yet its interplay with entanglement in dynamical settings remains underexplored. We study a well-controlled, analytically tractable setup that isolates entanglement generation from magic injection. We analytically and numerically demonstrate that stabilizer entanglement functions as a highway that facilitates the spreading of locally injected magic throughout the entire system. Specifically, for an initial stabilizer state with bipartite entanglement $E$, the total magic growth, quantified by the linear stabilizer entropy $Y$, follows $\overline{Y}\propto 2^{-|A|-E}$ under a Haar random unitary on a local subregion $A$. Moreover, when applying a tensor product of local Haar random unitaries, the resulting state’s global magic approaches that of a genuine Haar random state if the initial stabilizer state is sufficiently entangled by a system-size-independent amount. Similar results are also obtained for tripartite stabilizer entanglement. We further extend our analysis to non-stabilizer entanglement and magic injection via a shallow-depth brickwork circuit, and find that the qualitative picture of our conclusion remains unchanged.
Introduction.- Understanding what distinguishes quantum systems from classical ones has been a challenging and multi-faceted endeavor. Much of the impact of quantumness in physics has been focusing on the idea of quantum entanglement, which played a central role in characterizing topological order [1, 2, 3], dynamically induced phases [4], quantum chaos, and even the emergence of spacetime [5]. However, entanglement being a quantum correlation, is just one facet of quantumness. Quantum advantage, for example, is concerned with the hardness of simulating quantum systems on a classical computer. This facet is distinct from entanglement as many highly entangled systems prepared using Clifford operations violate Bell’s inequalities but can be efficiently simulated classically [6, 7]. The notion of classical hardness therefore constitutes a second layer of quantumness and is intimately connected to non-stabilizerness or magic [8, 9, 10, 11, 12, 13].
Like entanglement, magic is an important but distinct quantum resource for realizing fault-tolerant universal quantum computation [13, 14]. Therefore, it is crucial to understand any physical consequences played by this orthogonal dimension to entanglement in quantum systems. The importance of non-stabilizerness has been largely overlooked in quantum many-body physics until recently, where a flurry of activities have been dedicated to study its total extent in various systems, which has implication for resource estimation in state preparation and quantum simulation [15, 16, 17, 18, 19, 20], in characterizing phase transitions [21, 22], in quantum dynamics [23, 24, 25, 26, 27], in improved quantum simulation with tensor networks [28, 29], and in quantum gravity [30, 31]. Arguably, the most fascinating perspective comes from the interplay between magic and entanglement [32, 31, 33, 34, 35, 36, 27] as both are needed for quantum advantage.
Recent work revealed that, unlike the spread of entanglement which takes time linear in system size in spatially local Haar random circuits, magic appears to saturate in $O(\log n)$ time [37] (whereas it spreads much more slowly in Hamiltonian systems [38]). In the long time limit, it is also understood that magic in these highly entangled system is similar to the scrambling/encoding of information in quantum error correcting codes where it quickly delocalizes and can only be found when one can access a large part of the system [39, 26, 40]. This asymptotic behavior is consistent with the observation that subsystem magic quickly rises and decays to almost zero as a state thermalizes [24].
Despite these preliminary findings, the fundamental mechanisms underlying magic dynamics, particularly the interplay between magic injection and entanglement, remain poorly understood. Specifically, the role of entanglement in facilitating the growth and spread of magic, and vice versa, is yet to be determined in a quantitative manner. In generic quantum circuits and Hamiltonian dynamics, this interplay is inherently complex, as each time step of evolution simultaneously injects magic and increases entanglement [41, 37]. To gain deeper insights into magic dynamics, a well-controlled setup that isolates different contributing factors is essential. Here we study precisely one such setting and show that entanglement, or at least stabilizer entanglement, acts as a conduit for magic, allowing it to spread much faster to parts of the system that are entangled. This would constitute an important first step in understanding how magic dynamics is distinct and yet interlinked with various types of entanglement dynamics.
In this work, we analytically and numerically demonstrate that, given an initial stabilizer state with bipartite entanglement $E$, the total growth of magic, as quantified by the linear stabilizer entropy $Y$ under magic injection via a Haar-random unitary acting on a local subregion $A$, is enhanced by the presence of entanglement: $\overline{Y}\propto 2^{-|A|-E}$. Furthermore, when a unitary of the form $U=U_{A}\otimes U_{B}$, consisting of independently Haar-random local unitaries on subsystems $A$ and $B$, is applied, the global magic of the resulting state becomes nearly indistinguishable from that of a fully Haar-random state, provided the initial state is sufficiently entangled. Interestingly, the amount of entanglement necessary for such indistinguishability from Haar random magic is independent of system size. Similar results are also obtained for tripartite stabilizer entanglement. We further extend our analysis to non-stabilizer entanglement and magic injection via a shallow-depth brickwork circuit, and find that the qualitative picture of our conclusion remains unchanged in these cases.
Quantum magic and its measure.- Quantum magic quantifies the amount of “non-stabilizer” or “non-Clifford” resources in a quantum state. A computable measure for quantum magic, which we will primarily focus on in this work, is given by the linear stabilizer entropy. Denote by $P_{N}$ the set of $N$-qubit Pauli strings. The linear stabilizer entropy for a pure state $\rho=|\psi\rangle\langle\psi|$ is defined as $Y(\rho):=2^{-N}\sum_{P\in P_{N}}{\rm tr}(P\rho)^{4}$. To gain some intuition about this quantity, first consider a stabilizer pure state. There are precisely $2^{N}$ Pauli strings corresponding to the stabilizer group elements of $\rho$ for which ${\rm tr}(P\rho)^{4}=1$, and ${\rm tr}(P\rho)^{4}=0$ for all other Pauli strings. Thus $Y(\rho)=1$ for stabilizer states. For a generic non-stabilizer pure state, the Pauli spectrum satisfies the normalization condition: $12^{-N}\sum_{P\in P_{N}}{\rm tr}(P\rho)^{2}=$ 1. Therefore, $Y(\rho)$ defined above can be thought of as the second moment of the Pauli spectrum of $\rho$, which characterizes how delocalized the expectation values of $P$ are over all possible Pauli strings. The linear stabilizer entropy satisfies the following properties: (i) faithfulness: let $\text{STAB}_{n}$ denote the set of all $n$-qubit pure stabilizer states, then $Y(\rho)=1$ iff $\rho\in\text{STAB}_{n}$, otherwise $Y(\rho)<1$; (ii) invariant under Clifford unitaries $U_{C}$: $Y(U_{C}\rho U^{\dagger}_{C})=Y(\rho)$; (iii) multiplicative: $Y(\rho_{A}\otimes\rho_{B})=Y(\rho_{A})Y(\rho_{B})$; (iv) the closely related quantity $M^{\rm lin}:=1-Y$ is a strong pure-state monotone under free operations [42]. Hence, $Y(\rho)$ serves as a good measure for quantum magic from the point of view of magic-state resource theory. Moreover, the logarithm of $Y(\rho)$ is related to the second stabilizer Rényi entropy $M_{2}(\rho)=-{\rm log}Y(\rho)$, another widely used magic measure [43].
Figure 1: Schematics of the setup used for analytical and numerical calculations. The initial states are bipartite or tripartite stabilizer pure states. (a) A Haar random unitary injects magic by acting on a subregion $A$. (b) A factorized unitary $U=U_{A}\otimes U_{B}$, where $U_{A}$ and $U_{B}$ are independently Haar-random, acts on the initial stabilizer state. (c) A factorized unitary $U=U_{A}\otimes U_{B}$ injects magic by acting on subregion $AB$ of a tripartite system. (d) Initial state (2) for numerical simulations of the setup in (a). The state consists of $f_{A}$ and $f_{B}$ qubits in |+ket|$\rangle$| + ⟩ state in subsystem $A$ and $B$, respectively, along with $E$ Bell pairs shared between $A$ and $B$. (e) Initial state for obtaining Eq. (5). The state initial state consists of $g$ GHZ states shared among $A$, $B$, and $C$; $b_{AB}$, $b_{AC}$, and $b_{BC}$ Bell pairs shared between subsystems $AB$, $AC$, and $BC$, respectively, and $f_{A}$, $f_{B}$, $f_{C}$ single-qubit states in each individual subsystem.
Stabilizer entanglement facilitates magic injection.-To illustrate how stabilizer entanglement can influence magic injection, we examine the effect of a single-qubit $T$ gate on two distinct stabilizer states: an unentangled product state, $|\psi_{1}\rangle=|+\rangle^{\otimes N}=\left((|0\rangle+|1\rangle)/\sqrt{2}\right)^{\otimes N}$, and a GHZ state with long-range entanglement, $|\psi_{2}\rangle=(|0\rangle^{\otimes N}+|1\rangle^{\otimes N})/\sqrt{2}$. In both cases, the initial states contain no magic, and the $T$ gate injects magic locally. However, the manner in which this injected magic manifests differs significantly between the two cases. In $T|\psi_{1}\rangle$, the injected magic remains locally accessible and can be directly probed at the specific qubit where the $T$ gate is applied. In contrast, for $T|\psi_{2}\rangle$, the state appears non-magical to any local observer with access only to a subregion of the system. This simple example illustrates that the propagation of locally injected magic throughout the system is intrinsically tied to the entanglement structure of the initial state.
To further quantify how stabilizer entanglement facilitates magic spreading, we consider the setup illustrated in Fig. 1(a), which enables an analytical calculation. Specifically, we begin with an initial stabilizer pure state $\rho=|\psi\rangle\langle\psi|$, partitioned into subregion $A$ and its complement $B$. A Haar random unitary $U_{A}$ is then applied to subregion $A$, and we are interested in computing the linear stabilizer entropy of the resulting state $\sigma=U_{A}|\psi\rangle\langle\psi|U_{A}^{\dagger}$, averaged over Haar random unitaries. We first make a few remarks on this particular choice of setup before showing the result. This construction is a natural generalization of the toy example discussed above, replacing the single-qubit unitary with a Haar random unitary on a finite subregion. While the Haar-random unitary may appear nonlocal within subregion $A$, it is local relative to the entire system if $|A|\ll|B|$. The use of Haar random unitaries primarily serves to render the analytical calculations more tractable. We will show below that key qualitative features of our results persist even when considering more structured unitaries within subregion $A$. Secondly, this setup clearly separates entanglement generation with respect to the bipartition and magic injection. The unitary acting on $A$ injects magic to the system while preserving the amount of entanglement between $A$ and $B$, including the full entanglement spectrum. As many, if not most, operationally convenient circuits apply Clifford gates and non-Clifford phase gates in separate stages, e.g. Clifford+T decomposition [13] and the PFC ensemble that prepare pseudorandom unitaries [44], our setup is to probe the internal dynamics that take place between the most elementary layers in those circuits. This is complementary to the random unitary circuit model used in Ref. [37] to study magic saturation. In that model, unitary gates simultaneously generate entanglement and inject magic, leading to universal late-time magic dynamics prior to saturation, regardless of the initial state.
We analytically compute the linear stabilizer entropy $Y(U_{A}|\psi\rangle\langle\psi|U_{A}^{\dagger})$, averaged over the ensemble of Haar random unitaries $U_{A}$. In the limit $|A|\gg 1$, we obtain, to leading order,
$\overline{Y}=\mathbb{E}_{U_{A}}\ Y\left(U_{A}|\psi\rangle\langle\psi|\ U_{A}^{\dagger}\right)=4\cdot{{2}^{-|A|-E}}\left[1+O({2}^{-{|A|-E}})\right],$ (1)
where $E=-\rm{tr}(\rho_{A}\text{log}_{2}\rho_{A})$ is the amount of entanglement between $A$ and $B$, which is an integer for a stabilizer state $|\psi\rangle$. An exact calculation of $\overline{Y}$ for arbitrary $|A|$ and $E$ is possible, and is shown in the Supplemental Material (SM) [45]. Eq. (1) establishes a quantitative relation between the efficiency of local magic injection and the amount of stabilizer entanglement contained in the target state. For a fixed subregion $A$, the average $Y(\rho)$ decays exponentially with the amount of entanglement between $A$ and its complement. Therefore, global magic injection on $AB$ is enhanced with entanglement, which acts as a throttle that controls how much magic can flow to region $B$ from $A$. A detailed derivation of Eq. (1) is provided in the SM [45], yet its qualitative features, particularly the factor $2^{-|A|-E}$ can be understood straightforwardly. Since the unitary acts only on subsystem $A$, stabilizers supported entirely on $B$ remains unchanged under $U_{A}$ and contribute to $Y(\rho)$. A simple counting shows that the number of surviving stabilizers is precisely $2^{|B|-E}$, leading to the factor $2^{-|A|-E}$ in Eq. (1). The prefactor 4 in front of $2^{-|A|-E}$ is, in some sense, non-universal. To see this, let us define the generalized linear stabilizer entropy $Y_{\alpha}(\rho):=2^{-N}\sum_{P\in P_{N}}{\rm tr}(P\rho)^{2\alpha}$. Eq. (1) thus corresponds to the specific case $\alpha=2$. Similarly, $Y_{\alpha}(\rho)$ is related to the $\alpha$-th stabilizer Rényi entropy: $M_{\alpha}(\rho)=\frac{1}{1-\alpha}{\rm log}Y_{\alpha}(\rho)$. In the limit $\alpha\rightarrow\infty$, only the surviving stabilizers supported on $B$ contribute, yielding $\overline{Y_{\infty}}=2^{-|A|-E}$. For arbitrary $\alpha$, we have $\overline{Y_{\alpha}}=[1+c_{\alpha}(|A|,E)]2^{-|A|-E}$, with the one coming from the surviving stabilizers, and $c_{\alpha}(|A|,E)$ coming from contributions of all other Pauli strings, which in general depends on both $|A|$ and $E$ [46]. Specifically, Eq. (1) implies that for $\alpha=2$, we have $c_{2}=3$.
Figure 2: (a) Numerical simulations of the setup depicted in Fig. 1(a). We consider initial states of the form Eq. (2) with varying $|A|$ and $E$. We compute the average value of $\overline{Y}$ (upper triangles), as well as $2^{-\overline{M_{2}}}$ (lower triangles), where $\overline{M_{2}}$ is the average value of the second stabilizer Rényi entropy. The analytical result Eq. (1) is shown as dashed lines. (b) Numerical simulations of the setup depicted in Fig. 1(b). Choice of the initial state is the same as in (a). The analytic result Eq. (LABEL:eq:UaUb) is shown as dashed lines. (c) Exact results (squares) and leading order contributions Eq. (5) of the setup depicted in Fig. 1(c). (d) Same as in (a) but for non-stabilizer initial states with $k$ imperfect Bell pairs $|\phi_{\theta}\rangle=\cos\theta|00\rangle+\sin\theta|11\rangle$ shared between $A$ and $B$. We compute the magic already existing in the initial state $\overline{Y}_{\text{ini}}$ (dashed lines) and the change after the action of $U_{A}$: $\overline{Y}_{\text{final}}/\overline{Y}_{\text{ini}}$ (solid lines). (e) Same as in (c) but initial states are now constructed with four different types of entanglement spectra (see main text). Error bars are smaller than the size of the data points and hence not visible. (f) Same as in (d)&(e), but now plotting $\overline{Y}_{\rm final}$ in each case considered.
To test Eq. (1), particularly the regime of validity of neglecting subleading corrections, we perform numerical simulations of the setup depicted in Fig. 1(a). Recall that any stabilizer state $|\psi\rangle$ can be brought to the form of Eq. (2) by local Clifford unitaries [47]: $|\Phi\rangle=U^{\rm Cliff}_{A}\otimes U^{\rm Cliff}_{B}|\psi\rangle$ such that
$|\Phi\rangle=|+\rangle^{\otimes f_{A}}\ |{\rm Bell}\rangle_{AB}^{\otimes E}\ |+\rangle^{\otimes f_{B}},$ (2)
which is illustrated in Fig. 1(d). The state consists of $f_{A}$ and $f_{B}$ qubits in |+ket|$\rangle$| + ⟩ state in subsystem $A$ and $B$, respectively, along with $E$ Bell pairs shared between $A$ and $B$, so that the entanglement entropy between $A$ and $B$ is $E$. Since $\overline{Y}$ is obtained by averaging over the Haar random ensemble, $U^{\rm Cliff}_{A}$ can be simply absorbed into the average. Moreover, $Y(\rho)$ is invariant under Clifford unitary $U_{B}^{\rm Cliff}$. Thus we have
$\mathbb{E}_{U_{A}}Y\left(U_{A}|\psi\rangle\langle\psi|U_{A}^{\dagger}\right)=\mathbb{E}_{U_{A}}Y\left(U_{A}|\Phi\rangle\langle\Phi|U_{A}^{\dagger}\right)$ (3)
for an arbitrary stabilizer state $|\psi\rangle$. We numerically average over $Y(U_{A}|\Phi\rangle\langle\Phi|U_{A}^{\dagger})$ with Haar random $U_{A}$ for various values of $|A|$ and $E$. The results are shown in Fig. 2(a), where we find excellent agreement with the analytical prediction of Eq. (1). In particular, the subleading corrections become negligible already at moderate subsystem sizes and entanglement $|A|+E\gtrsim 4$. We also simulate the Haar averaged second stabilizer Rényi entropy $\overline{M_{2}}$, which turns out to be extremely close to ${\rm log}\ {\overline{Y}}$. This implies that the average in Eq. (1) is also typical, as sample-to-sample fluctuations are extremely small. The same qualitative behavior holds if $U_{A}$ is given by a brickwork circuit instead of Haar random unitary, but the total injected magic is limited by the depth (see SM [45]).
One can further explore the most general approach to magic injection while preserving the entanglement spectrum. Starting from a stabilizer state, we now apply a unitary that is a tensor product of local unitaries acting on subsytems $A$ and $B$: $U=U_{A}\otimes U_{B}$, as illustrated in Fig. 1(b). We compute the linear stabilizer entropy of the resulting state, averaged over Haar random $U_{A}$ and $U_{B}$. In the limit where $|A|\gg 1$ and $|B|\gg 1$, we find [45]
$\overline{Y}$ $=$ $\mathbb{E}_{U_{A}}\mathbb{E}_{U_{B}}Y\left(U_{A}U_{B}|\psi\rangle\langle\psi|U_{A}^{\dagger}U_{B}^{\dagger}\right)$
$=$ $4\cdot{2}^{-N}\left[1+3\cdot 2^{-2E}+O({{2}^{-|A|-E}})+O({{2}^{-|B|-E}})\right],$
where $N=|A|+|B|$ is the total number of qubits. This result reveals an intriguing mechanism by which stabilizer entanglement facilitates magic saturation. Notice that the average value of $\overline{Y}$ for Haar random states is given by $4\cdot 2^{-N}$. Eq. (LABEL:eq:UaUb) shows that the gap between $\overline{Y}$ for a stabilizer state acted upon by $U_{A}\otimes U_{B}$ and that for a Haar random state decays exponentially with the amount of entanglement. In other words, despite the applied unitary being a tensor product of local unitaries on subsystems $A$ and $B$, from the perspective of magic, its effect on magic injection closely resembles that of a genuinely global Haar-random unitary, provided that the initial state is sufficiently entangled. Interestingly, the amount of entanglement necessary for such indistinguishability from Haar random magic is independent of system size. In Fig. 2(b), we numerically simulate this set up using the initial state (2) and averaging over Haar random unitaries $U_{A}$ and $U_{B}$. Again, we find excellent agreement with the prediction of Eq. (LABEL:eq:UaUb), with the subleading corrections negligible for moderate subsystem sizes and entanglement. Moreover, the average $\overline{Y}$ saturates to that of a Haar random state already at $E\approx 3$, which is independent of system size.
Tripartite entanglement.- The above results for bipartite systems can be naturally extended to tripartite systems. Consider a stabilizer state shared among three parties $A$, $B$ and $C$. Since any tripartite stabilizer state can be transformed via local unitaries acting on $A$, $B$, and $C$ into a collection of GHZ state, Bell pair states, and single-qubit states [48], we can, without loss of generality, consider initial states as depicted in Fig. 1(e). Specifically, the initial state consists of $g$ GHZ states shared among $A$, $B$, and $C$; $b_{AB}$, $b_{AC}$, and $b_{BC}$ Bell pairs shared between subsystems $AB$, $AC$, and $BC$, respectively, and $f_{A}$, $f_{B}$, $f_{C}$ single-qubit states in each individual subsystem. Since magic injection via unitary operations on any single subsystem in this setting mirrors the bipartite case discussed earlier, we first consider unitaries of the form $U=U_{A}\otimes U_{B}$ [Fig. 1(c)]. In the limit where $|A|\gg 1$ and $|B|\gg 1$, we find that the average $\overline{Y}$ after applying $U$ is given by [45]
$\overline{Y}$ $=$ $\mathbb{E}_{U_{A}}\mathbb{E}_{U_{B}}Y\left(U_{A}U_{B}|\psi\rangle\langle\psi|U_{A}^{\dagger}U_{B}^{\dagger}\right)$ (5)
$=$ $4\cdot{2}^{-|A|-|B|-g-{{b}_{AC}}-{{b}_{BC}}}\left[1+3\cdot{2}^{-2{{b}_{AB}}-g}\right],$
where subleading corrections are of order $O(2^{-|A|},2^{-|B|})$. Eq. (5) can be understood as follows. The average value $\overline{Y}$ is once again exponentially suppressed in both the total size of the subregion where the unitary is applied: $|A|+|B|$, and the amount of entanglement between subregion $AB$ and its complement: $g+b_{AC}+b_{BC}$. Hence the physical interpretation of the prefactor in Eq. (5) is identical to that of Eq. (1). The discrepancy between these two cases lies in the additional factor $3\cdot{2}^{-2{{b}_{AB}}-g}$, which becomes negligible once $A$ and $B$ are sufficiently entangled. This is precisely what we have seen in Eq. (LABEL:eq:UaUb): the effect of a factorized unitary $U_{A}\otimes U_{B}$ becomes essentially indistinguishable from that of a global unitary from the standpoint of magic, once $A$ and $B$ are sufficiently entangled. Thus, Eq. (5) encapsulates the combined effects of Eqs. (1) and (LABEL:eq:UaUb) in the tripartite setting. In the SM [45], we further consider a factorized unitary acting on all three subregions: $U=U_{A}\otimes U_{B}\otimes U_{C}$, where we confirm that in this case the effect of a factorized Haar random unitary on magic injection again becomes indistinguishable from that of a global Haar random unitary, provided that the subregions are sufficiently entangled with one another.
Non-stabilizer entanglement.- So far, we have focused on the propagation of magic on stabilizer states. However, realistic dynamics almost always contains magic in the initial state. Does non-stabilizerness of this kind impose any road blocks on the magic highway? Our findings strongly suggest that, to leading order, the answer is no. Again consider an entangled initial state on $AB$ where $U_{A}$ is applied to inject magic. Since any initial magic local to $A$ can be absorbed via the averaging of $U_{A}$, let us start with a state where non-stabilizerness only lives in the quantum correlation. Such a state is known to have non-local magic, which is non-trivial as long as the entanglement spectrum is non-flat [31]. The simplest example of such a state consists of $k$ imperfect Bell pairs shared between $A$ and $B$: $|\psi\rangle=|0\rangle^{\otimes f_{A}}|\phi_{\theta}\rangle_{AB}^{\otimes k}$, where $|\phi_{\theta}\rangle=\cos\theta|00\rangle+\sin\theta|11\rangle$. Since $k$ copies of $|\phi_{\theta}\rangle$ each with entanglement $S$ can be unitarily distilled into $E=Sk-O(\sqrt{k})$ number of perfect Bell states tensoring a magic state on the $O(\sqrt{k})$ remaining qubits to good approximation [49], our previous analysis with stabilizer initial state again applies to leading order. That is, magic injection by $U_{A}$ will increase in the same way as one increases the total entanglement $E$. Interestingly, this simple dependence also appears to extend to other states with non-local magic. For example, initial states with non-local magic can be parametrized in the form $|\psi\rangle=|0\rangle^{\otimes f_{A}}\sum_{i=1}^{2^{k}}\lambda_{i}|i\rangle_{A}\otimes|i\rangle_{B}$, where $|i\rangle$ denotes a computational basis state of $k$ qubits. We choose four arbitrary entanglement spectra: (i) $\lambda_{i}\propto e^{-i/2^{k}}$; (ii) $\lambda_{i}\propto i$; (iii) $\lambda_{i}\propto i^{2}$; (iv) $\lambda_{i}\propto i^{3}$ and examine the magic spread numerically. Magic propagation is again displaying an $E$ dependence similar to having a stabilizer state or imperfect Bell states as the initial state [Fig. 2.(f)].
Nonetheless, more care is needed going beyond leading order. It is clear that there is non-stabilizerness dependence even in the simplest example with imperfect Bell states at the order of $O(\sqrt{k})$. Furthermore, because initial states can contain magic, the remaining magic capacity for further injection is reduced, i.e., the change of total magic per magic state injection is smaller compared to stabilizer entanglement. We numerically calculate both the magic already existing in the initial states $\overline{Y}_{\text{ini}}$ and the change after the action of $U_{A}$: $\overline{Y}_{\text{final}}/\overline{Y}_{\text{ini}}$. As shown in Fig. 2(d,e), $\overline{Y}_{\rm ini}$ itself in general decays with the increase of entanglement. Nonetheless, we find that the ratio $\overline{Y}_{\text{final}}/\overline{Y}_{\text{ini}}$ also decays exponentially with the amount of entanglement. For the same $E$, the change in global magic becomes smaller as the initial state becomes more magical.
Discussion.- In this work, we demonstrate that stabilizer entanglement functions as a highway that facilitates the spread of locally injected magic throughout the entire system. In particular, we consider an analytically tractable setup where a local Haar random unitary acts on a subregion of a non-magical but potentially highly entangled state, thereby separating the process of magic injection from entanglement generation. We find that the average magic, as quantified by the linear stabilizer entropy, is enhanced by stabilizer entanglement. We further generalize our results to tripartite entanglement and non-stabilizer entanglement. A few interesting open questions remain. Our results for imperfect Bell-pair states suggest that the average amount of injected magic decreases as the amount of magic already existing in the target state increases. This implies that there may be a notion of intrinsic magic capacity of a given state that depends on both its entanglement and magic. Addressing this question would require a more systematic extension of our results to non-stabilizer states, which we leave for future work. Secondly, once we replace the Haar random $U_{A}$ with a more structured brickwork circuit, the question of what is the minimal circuit depth required for the effect of magic injection to coincide with that of a Haar random $U_{A}$ is worth exploring. The question is perhaps intimately related to the complexity growth of local random unitary circuits with circuit depth.
Acknowledgement.- We thank Cheng Wang, Yingfei Gu, and Yuzhen Zhang for helpful discussions, and Xhek Turkeshi for useful comment on the manuscript. We especially thank Daniele Iannotti, Gianluca Esposito, Lorenzo Campos Venuti, and Alioscia Hamma for sharing their note on obtaining the exact expression of average stabilizer purity at a fixed amount of entanglement. This work is supported by Grant No. 12375027 from the National Natural Science Foundation of China (Z.-C. Y.). Numerical simulations were performed on the High-performance Computing Platform of Peking University.
Supplemental Material for “Stabilizer entanglement as a magic highway”
Appendix A Analytical method
In this section, we present the analytical method used to derive the main results quoted in the main text. In Section A.1, we provide a brief introduction to a diagrammatic approach for computing polynomials of matrix elements of unitary matrices averaged over the Haar ensemble. Section A.2 introduces a key quantity used throughout our analysis and outlines approximations made to simplify both the calculation and the final results. Finally, in Section A.3, we detail the step-by-step derivation of the results presented in the main text. Exact results of all setups considered in the main text (i.e. without any approximation such as $|A|\gg 1$) can also be obtained using an alternative analytical method, as provided in Appendix E. This alternative approach, while more compatible with symbolic computations, offers less physical intuition.
A.1 A diagrammatic approach
In this subsection, we outline a diagrammatic method for systematically computing expectation values of polynomials in the matrix elements of unitary matrices under the Haar ensemble. Our presentation closely follows the original work of Brouwer (1996) [50]. Consider a polynomial function of the form
$f(U)=U_{a_{1}b_{1}}\dots U_{a_{n}b_{n}}U^{*}_{\alpha_{1}\beta_{2}}\dots U^{*}_{\alpha_{m}\beta_{m}}.$ (6)
where $U_{ij}$ and $U_{kl}^{*}$ denote matrix elements of a unitary matrix and its complex conjugate, respectively. The Haar measure expectation value of such functions, denoted by $\overline{f(U)}:=\mathbb{E}_{\rm Haar}[f(U)]$, is given by
$\overline{f(U)}=\delta_{nm}\sum_{\sigma,\pi}V_{\sigma,\pi}\prod\limits_{j=1}^{n}\delta_{a_{j}\alpha_{\sigma(j)}}\delta_{b_{j}\beta_{\pi(j)}}$ (7)
where the summation is over all permutations $P$ and $P^{\prime}$ of the integers $(1,\dots,n)$. The coefficients $V_{\sigma,\pi}$, also known as the Weingarten function, depend only on the cycle structure of the permutation $\sigma^{-1}\pi$ [51]. In other words, it only depends on the lengths $c_{1},\dots,c_{k}$ of the cycles in the factorization of $\sigma^{-1}\pi$. So below we write $V_{c_{1},\dots,c_{k}}$ instead of $V_{\sigma,\pi}$.
The diagrams consist of the building blocks shown in Fig. 3. The matrix elements $U_{ab}$ or $U^{*}_{\alpha\beta}$ are represented by thick dotted lines. The first index ($a$ or $\alpha$) is represented by a black dot, the second index ($b$ or $\beta$) is represented by a white dot. Matrix element $A_{ij}$ of a fixed matrix $A$ (i.e., a matrix that is not part of the average) is represented by a directed thick solid line, pointing from the first to the second index, without dots at the endpoints. The Kronecker delta is represented by an undirected thin solid line, without dots at the endpoints. Two dots connected by a solid line indicate contractions of their corresponding matrix indices. As an example, the functions $f(U)=\text{Tr}(AUBU^{\dagger})$ and $g(U)=\text{Tr}(AUBUCU^{\dagger}DU^{\dagger})$ are represented in Fig. 4.
Figure 3: Diagrammatic representations for the unitary matrix $U$ and $U^{*}$, the fixed matrix $A$ and the Kronecker delta.
Figure 4: Top: diagrams representing $f(U)=\text{Tr}(AUBU^{\dagger})$ and $g(U)=\text{Tr}(AUBUCU^{\dagger}DU^{\dagger})$, respectively. Bottom: diagrams that keep track of all terms after taking the average over Haar random $U$. The average of $\overline{f(U)}$ contains only one term, whereas $\overline{g(U)}$ contains four terms.
Since the average over Haar random ensemble leads to a summation of permutations, we need to keep track of each term in the summation in Eq. (7) by keeping track of $P$ and $P^{\prime}$ in the diagram. Each $P$ represents one way of pairing the first index of $U$ with that of $U^{*}$, and $P^{\prime}$ pairs the second index of $U$ with that of $U^{*}$. This pairing is represented by adding additional thin lines connecting the endpoints of $U$ and that of $U^{*}$ in the diagram, with black dots connected to black dots, and white dots connected to white dots, as depicted in Fig. 4. To associate each diagram with a specific value, we follow the following rules:
(i) Cycle structure: A closed circuit consisting of alternating dotted or thin lines correspond to a cycle in $P^{-1}P^{\prime}$. The length $c_{k}$ of it is half the number of dotted lines in the circuit. We call the circuit a $U$-cycle of length $c_{k}$.
(ii) Matrix trace: A closed circuit consisting of alternating thick or thin lines is called a $T$-cycle. A $T$-cycle containing matrices $A^{(1)},A^{(2)},\dots,A^{(k)}$ (ordered when going in one direction of the closed circuit) corresponds to $\text{Tr}A^{(1)}A^{(2)}\dots A^{(k)}$ (recall that thin lines represent $\delta_{ab}$ that locks the two indices they connect). If a thick line corresponding to matrix $A$ is traversed in the opposite direction, the matrix is replaced by its transpose $A^{T}$.
To illustrate this procedure, we consider the averages of the functions $\overline{f(U)}=\overline{\text{Tr}(AUBU^{\dagger})}$ and $\overline{g(U)}=\overline{\text{Tr}(AUBUCU^{\dagger}DU^{\dagger})}$. By connecting the dots with thin lines, we obtain the diagrams shown in Fig. 4. For $f$, there is only one diagram, which consists of a single $U$-cycle of length 1 (with weight $V_{1}$) and two $T$-cycles, which generate $\text{Tr}A$ and $\text{Tr}B$. Thus, the expected value is:
$\overline{f(U)}=V_{1}\text{Tr}A\ \text{Tr}B.$ (8)
As for $g$, there are four diagrams that contribute. The first diagram contains two $U$-cycles of length $1$, and three $T$-cycles. Its contribution is $V_{1,1}\text{Tr}A\ \text{Tr}(BD)\ \text{Tr}C$. The second diagram contains two $U$-cycles of length $1$ and a single $T$-cycle. Its contribution is $V_{1,1}\text{Tr}(ABCD)$. The third and fourth diagram each contain a single $U$-cycle of length $2$ and two $T$-cycles. Their contributions are $V_{2}\text{Tr}A\ \text{Tr}(BDC)$ and $V_{2}\text{Tr}(ADB)\ \text{Tr}C$. So we obtain:
$\overline{g(U)}=V_{1,1}\left[\text{Tr}(ABCD)+\text{Tr}A\ \text{Tr}(BD)\ \text{Tr}C\right]+V_{2}\left[\text{Tr}A\ \text{Tr}(BDC)+\text{Tr}(ADB)\ \text{Tr}C\right].$ (9)
A.2 A useful quantity and its approximations
In this subsection, we define a quantity that will be used intensively in evaluating $\overline{Y}$. We will calculate the value of this quantity in its most general form, to leading order in $2^{|A|+E}$, where $E$ is the entanglement entropy. Let $\mathbb{S}_{A}$ denote an $|A|$-qubit stabilizer group whose size is $2^{f_{A}}$ with $f_{A}<|A|$; then let $u_{i},i=1,2,3,4$ denote four Pauli strings supported on $A$ (where some of them can be identical) that commute with every element in $\mathbb{S}_{A}$ (i.e. $u_{i}$ is in the normalizer group of $\mathbb{S}_{A}$). Notice that we allow $u_{i}=I_{A}$, but otherwise $u_{i}\notin\mathbb{S}_{A}$. Since we assume $f_{A}<|A|$, one can always find a nontrivial $u_{i}$. And because $u_{i}\mathbb{S}_{A}=u_{i}g\mathbb{S}_{A},\forall g\in\mathbb{S}_{A}$, we can choose one specific $u_{i}$. Then we define the following function:
${{\alpha}_{{{P}_{A}}}}({{u}_{1}},{{u}_{2}},{{u}_{3}},{{u}_{4}}):=\overline{\prod\limits_{i=1}^{4}{\text{Tr}({{P}_{A}}{{U}_{A}}{{u}_{i}}\mathbb{S}_{A}{U}_{A}^{\dagger})}},$ (10)
where $P_{A}$ is a Pauli string supported on subsystem $A$. In the above definition, we also abuse notation by using $\mathbb{S}_{A}$ to represent an equal weight summation of all elements in $\mathbb{S}_{A}$: $\sum_{g\in\mathbb{S}_{A}}g$. The connection between the above quantity and $\overline{Y}$ will be explained in detail in the next subsection. Basically it naturally arises when computing ${\rm Tr}(PU_{A}\rho U_{A}^{\dagger})^{4}$, with $\rho$ represented as a projector onto the stabilizer subspace.
Eq. (10) can be represented using the diagrammatic method introduced in Sec. A.1 as follows:
Figure 5: (a) Leading order and the corresponding diagrams for $\alpha_{P_{A}}(u_{1},u_{2},u_{3},u_{4})$. There are $576{4!}^{2}=$ 576 diagrams in total. If $P_{A}\neq I_{A}$, then $\text{Tr}(P_{A})=\text{Tr}(P_{A}^{3})=0$, and there are 216 diagrams with non-zero contributions. (b) Subleading order and the corresponding diagrams for $\alpha_{P_{A}}(u_{1},u_{2},u_{3},u_{4})$.
First of all, the factors $V_{c_{1},c_{2},\dots}$ can be readily obtained using known results of Weingarten functions. Let $N:=2^{|A|}$, we have:
$V_{1,1,1,1}=\frac{N^{4}-8N^{2}+6}{N^{2}(N^{2}-1)(N^{2}-4)(N^{2}-9)}=2^{-4|A|}+6\cdot 2^{-6|A|}+O(2^{-8|A|}),$ (11)
$V_{2,1,1}=\frac{-N^{3}+4N}{N^{2}(N^{2}-1)(N^{2}-4)(N^{2}-9)}=-2^{-5|A|}+18\cdot 2^{-7|A|}+O(2^{-9|A|}),$
$V_{2,2}=\frac{N^{2}+6}{N^{2}(N^{2}-1)(N^{2}-4)(N^{2}-9)}=2^{-6|A|}+20\cdot 2^{-8|A|}+O(2^{-10|A|}),$
$V_{3,1}=\frac{2N^{2}-6}{N^{2}(N^{2}-1)(N^{2}-4)(N^{2}-9)}=2\cdot 2^{-6|A|}+8\cdot 2^{-8|A|}+O(2^{-10|A|}),$
$V_{4}=\frac{-5N}{N^{2}(N^{2}-1)(N^{2}-4)(N^{2}-9)}=-5\cdot 2^{-7|A|}+O(2^{-9|A|}).$
Careful inspections of all diagrams reveals that factors involving $P_{A}$ are of the form $\text{Tr}^{4}(P_{A})$, $\text{Tr}(P_{A}^{3})\text{Tr}({P_{A}})$, $\text{Tr}(P_{A}^{2})\text{Tr}^{2}({P_{A}})$, $\text{Tr}^{2}(P_{A}^{2})$, and $\text{Tr}(P_{A}^{4})$. Notice that $P_{A}^{2}=P_{A}^{4}=I_{A}$, $P_{A}^{3}=P_{A}$. Also, if $P_{A}\neq I_{A}$, ${\rm Tr}(P_{A})=0$. Thus, in this case, the only non-zero contributions come from terms involving $\text{Tr}^{2}({P_{A}}^{2})=2^{2|A|}$ and $\text{Tr}({P_{A}}^{4})=2^{|A|}$. Similarly, terms involving $u_{i}$ and $\mathbb{S}_{A}$ are of the form $\text{Tr}(u_{1}u_{2}u_{3}u_{4}{\mathbb{S}}_{A}^{4})$, $\text{Tr}(u_{1}u_{2}u_{3}{\mathbb{S}}_{A}^{3})\text{Tr}(u_{4}{\mathbb{S}}_{A})$, $\text{Tr}(u_{1}u_{2}{\mathbb{S}}_{A}^{2})\text{Tr}(u_{3}u_{4}{\mathbb{S}}_{A}^{2})$, $\text{Tr}(u_{1}u_{2}{\mathbb{S}}_{A}^{2})\text{Tr}(u_{3}{\mathbb{S}}_{A})\text{Tr}(u_{4}{\mathbb{S}}_{A})$, and $\text{Tr}(u_{1}{\mathbb{S}}_{A})\text{Tr}(u_{2}{\mathbb{S}}_{A})\text{Tr}(u_{3}{\mathbb{S}}_{A})\text{Tr}(u_{4}{\mathbb{S}}_{A})$ (and also permutations of the $u_{i}$’s). It is also useful to remember ${\rm Tr}(\mathbb{S}_{A})=2^{|A|}$, and ${\rm Tr}(\mathbb{S}_{A}^{2})=2^{|A|+f_{A}}$.
When $P_{A}=I_{A}$, $\alpha_{I_{A}}$ trivially becomes $\prod_{i=1}^{4}{\rm Tr}(u_{i}\mathbb{S}_{A})$, which vanishes unless $u_{i}=I_{A}$, $\forall i$. We thus have:
${{\alpha}_{{{I}_{A}}}}({{u}_{1}},{{u}_{2}},{{u}_{3}},{{u}_{4}})=2^{4|A|}\delta_{1}(u_{1})\delta_{1}(u_{2})\delta_{1}(u_{3})\delta_{1}(u_{4}),$ (12)
where $\delta_{1}(u_{i})=1$ iff $u_{i}=I_{A}$ and zero otherwise.
When $P_{A}\neq I_{A}$, by carefully examining the structures of all diagrams, we find that the leading order contributions to $\alpha_{P_{A}}$ come from the diagrams shown in Fig. 5.
${{\alpha}_{{{P}_{A}}}}({{u}_{1}},{{u}_{2}},{{u}_{3}},{{u}_{4}})$ (13)
$=2^{2f_{A}}\left(\underbrace{{{\delta}_{2}}({{u}_{1}},{{u}_{2}}){{\delta}_{2}}({{u}_{3}},{{u}_{4}})+\text{permutations}}_{3\text{ terms}}\right)$
$\quad-2^{f_{A}}\left(\underbrace{{{\delta}_{2}}({{u}_{1}},{{u}_{2}}){{\delta}_{1}}({{u}_{3}}){{\delta}_{1}}({{u}_{4}})+\text{permutations}}_{\text{6 terms}}\right)$
$\quad+O\left(2^{3f_{A}-2|A|}\delta_{4}(u_{1},u_{2},u_{3},u_{4})\right),$
where $\delta_{2}(u_{1},u_{2})=1$ iff $u_{1}=u_{2}$, and zero otherwise; $\delta_{4}(u_{1},u_{2},u_{3},u_{4})=\frac{\text{Tr}(u_{1}u_{2}u_{3}u_{4})}{2^{|A|}}$. In the above equation, the first line comes from terms like $V_{1,1,1,1}{\rm Tr}^{2}(P_{A}^{2}){\rm Tr}(u_{1}u_{2}\mathbb{S}_{A}^{2}){\rm Tr}(u_{3}u_{4}\mathbb{S}_{A}^{2})$, and the 2nd line comes from $V_{2,1,1}{\rm Tr}^{2}(P_{A}^{2}){\rm Tr}(u_{1}u_{2}\mathbb{S}_{A}^{2}){\rm Tr}(u_{3}\mathbb{S}_{A}){\rm Tr}(u_{4}\mathbb{S}_{A})$.
A.3 Derivation of the key results presented in the main text
In this subsection, we provide a detailed derivation of Eqs. (1), (4) and (5) in the main text.
A.3.1 Eq. (1)
We start with Eq. (1). The key to our derivation is to represent the initial state as a projector onto the stabilizer subspace, and a decomposition of the stabilizer group of the full sytem in terms of the cosets of stabilizer subgroups of subsystems $A$ and $B$. Let $\mathbb{S}_{AB}$ denote the stabilizer group of the initial state. Then initial density matrix can be written as a projector
$\rho_{0}=\frac{1}{2^{N}}\sum_{g\in\mathbb{S}_{AB}}g.$ (14)
Let $\mathbb{S}_{A}$ and $\mathbb{S}_{B}$ denote the stabilizer subgroups of subsystem $A$ and $B$, respectively. Let $E$ denote the entanglement entropy between subsystem $A$ and $B$. We have $E=|A|-{\rm dim}(\mathbb{S}_{A})=|B|-{\rm dim}(\mathbb{S}_{B})$. $\mathbb{S}_{AB}$ can then be decomposed in terms of the cosets of $\mathbb{S}_{A}$ and $\mathbb{S}_{B}$ as follows:
$\mathbb{S}_{AB}=\underbrace{\mathbb{S}_{A}\otimes\mathbb{S}_{B}\ \cup\ a_{1}\mathbb{S}_{A}\otimes b_{1}\mathbb{S}_{B}\ \cup\ a_{2}\mathbb{S}_{A}\otimes b_{2}\mathbb{S}_{B}\ \cup\ldots}_{4^{E}\ \rm{terms}}$ (15)
where $a_{i}\notin\mathbb{S}_{A}$ and $b_{i}\notin\mathbb{S}_{B}$ are the logical operators of subsystems $A$ and $B$, respectively. To ensure the proper commutation relation, the logical operators must be paired up into $(a_{i},b_{i})$ such that: (i) if $[a_{i},a_{j}]=0$, then $[b_{i},b_{j}]=0$; or (ii) if $\{a_{i},a_{j}\}=0$, then $\{b_{i},b_{j}\}=0$. While it is not hard to convince oneself that the above decomposition is correct for pure states with ${\rm dim}(\mathbb{S}_{AB})=N$, let us check that the dimensions work out. First of all, since the entanglement between $A$ and $B$ is $E$, there are precisely $E$ logical qubits encoded in both subsystems, with $4^{E}$ logical operators in each subsytem, and hence $4^{E}$ terms in the decomposition. Secondly, we can count the total number of group elements in $\mathbb{S}_{AB}$ from the right-hand side of the decomposition. We have $4^{E}\times 2^{{\rm dim}(\mathbb{S}_{A})}\times 2^{{\rm dim}(\mathbb{S}_{B})}=2^{|A|+|B|}=2^{N}$, as it should. We will give a concrete example of such a decomposition in Appendix B.
We can now express the initial state using the above coset decomposition of $\mathbb{S}_{AB}$:
${{\rho}_{0}}=\frac{1}{{2}^{N}}{\mathbb{S}_{AB}}=\frac{1}{2^{N}}\left(\underbrace{\mathbb{S}_{A}\otimes\mathbb{S}_{B}\ +\ a_{1}\mathbb{S}_{A}\otimes b_{1}\mathbb{S}_{B}\ +\ a_{2}\mathbb{S}_{A}\otimes b_{2}\mathbb{S}_{B}\ +\ldots}_{4^{E}\ \rm{terms}}\right)$ (16)
Armed with the above representation of $\rho_{0}$, we can cast the average value $\overline{Y}$ as:
$\overline{Y}=\frac{1}{{{2}^{5N}}}\sum\limits_{P\in P_{N}}{\overline{{{\text{Tr}}^{4}}(P{{U}_{A}}{\mathbb{S}_{AB}}{{U}_{A}}^{\dagger})}}$ (17)
$=\frac{1}{{{2}^{5N}}}\sum_{P\in P_{N}}{\sum\limits_{k=0}^{{{2}^{2E}-1}}{{{\alpha}_{{{P}_{A}}}}({{a}_{k}},{{a}_{k}},{{a}_{k}},{{a}_{k}})\ {{\text{Tr}}^{4}}({{P}_{B}}{{b}_{k}}{{\mathbb{S}}_{B}})}}.$
In going from the first to the second line, notice that the Pauli string $P=P_{A}\otimes P_{B}$, and that the unitary $U_{A}$ is supported on subsystem $A$ and hence leaves $P_{B}$ invariant. Moreover, since different cosets $b_{j}\mathbb{S}_{B}$ do not overlap, for a given $P$ (and $P_{B}$), there is at most one coset such that ${\text{Tr}^{4}({{P}_{B}}{{b}_{k}}{{\mathbb{S}}_{B}})}$ does not vanish. The trace involving $P_{A}$ is now exactly of the form in Eq. (10), with $u_{1}=u_{2}=u_{3}=u_{4}=a_{k}$ for some $k$. We can now directly apply our general results in Eqs. (12) and (13).
Consider first a fixed Pauli string $P$ and $k$, which corresponds to one term of the summation in $\overline{Y}$. We have the following result:
${{{\alpha}_{{{P}_{A}}}}({{a}_{k}},{{a}_{k}},{{a}_{k}},{{a}_{k}})\ \text{Tr}^{4}({{P}_{B}}{{b}_{k}}{{\mathbb{S}}_{B}})}$ (18)
$=$ $\delta(P_{B},b_{k}\mathbb{S}_{B})\times\begin{cases}3\cdot{{2}^{2{|A|}+4{|B|}-2E}}\left[1-2\cdot{{2}^{E-{|A|}}}+O({{2}^{-E-{|A|}}})\right],&\quad{{P}_{A}}\neq{{I}_{A}},\quad k=0\\ 3\cdot{{2}^{2{|A|}+4{|B|}-2E}}\left[1+O({{2}^{-E-{|A|}}})\right],&\quad{{P}_{A}}\neq{{I}_{A}},\quad k\neq 0\\ {{2}^{4{|A|}+4{|B|}}},&\quad{{P}_{A}}={{I}_{A}},\quad k=0\\ 0,&\quad{{P}_{A}}={{I}_{A}},\quad k\neq 0\\ \end{cases}$
In the above equation, $k=0$ corresponds to the first term in the coset decomposition, with $a_{0}=b_{0}=I$; $\delta(P_{B},b_{k}\mathbb{S}_{B})=1$ iff $P_{B}\in b_{k}\mathbb{S}_{B}$, and zero otherwise. Summing over $P\in P_{N}$ and $k$, we have:
$\sum\limits_{P\in P_{N}}{\sum\limits_{k=0}^{{{2}^{2E}-1}}{{{\alpha}_{{{P}_{A}}}}({{a}_{k}},{{a}_{k}},{{a}_{k}},{{a}_{k}})\ {{\text{Tr}}^{4}}({{P}_{B}}{{b}_{k}}{{\mathbb{S}}_{B}})}}$
$=$ $\underbrace{{2}^{|B|-E}\cdot(2^{2|A|}-1)\cdot 3\cdot{{2}^{2{|A|}+4{|B|}-2E}}\left[1-2\cdot{{2}^{E-{|A|}}}O({{2}^{-E-{|A|}}})\right]}_{P_{A}\neq I_{A},\quad k=0}$
$+$ $\underbrace{{2}^{|B|-E}\cdot(2^{2E}-1)\cdot(2^{2|A|}-1)\cdot 3\cdot{{2}^{2{|A|}+4{|B|}-2E}}\left[1+O({{2}^{-E-{|A|}}})\right]}_{P_{A}\neq I_{A},\quad k\neq 0}$
$+$ $\underbrace{{2}^{|B|-E}\cdot{2}^{4|A|+4|B|}}_{P_{A}=I_{A},\quad k=0}$
$=$ $4\cdot{2}^{4|A|+5|B|-E}\left[1+O({2}^{-|A|-E})\right].$ (19)
In the above summation, the prefactor $(2^{2|A|}-1)$ counts the total number of $P_{A}\neq I_{A}$; $(2^{|B|-E})$ comes from the size of each coset $|b_{k}\mathbb{S}_{B}|$. The final result is precisely Eq. (1) quoted in the main text after dividing by $2^{5N}$.
A.3.2 Eq. (4)
We provide detailed calculations for Eq. (4) in the main text. Since the unitary acting on $\rho_{0}$ takes a factorized form $U=U_{A}\otimes U_{B}$, the average $\overline{Y}$ similarly factorizes into
$\overline{Y}=\frac{1}{{{2}^{5N}}}\sum\limits_{P\in P_{N}}{\sum\limits_{{{k}_{1}},{{k}_{2}},{{k}_{3}},{{k}_{4}}}{{{\alpha}_{{{P}_{A}}}}({{a}_{{{k}_{1}}}},{{a}_{{{k}_{2}}}},{{a}_{{{k}_{3}}}},{{a}_{{{k}_{4}}}}){{\beta}_{{{P}_{B}}}}({{b}_{{{k}_{1}}}},{{b}_{{{k}_{2}}}},{{b}_{{{k}_{3}}}},{{b}_{{{k}_{4}}}})}},$ (20)
where $\beta_{P_{B}}\left(b_{k_{1}},b_{k_{2}},b_{k_{3}},b_{k_{4}}\right)$ is defined analogously to Eq. (10).
For convenience, we define the following functions:
$f_{1}:=\underbrace{{{\delta}_{2}}({{a}_{k_{1}}},{{a}_{k_{2}}}){{\delta}_{2}}({{a}_{k_{3}}},{{a}_{k_{4}}})+\text{permutations}}_{3\text{ terms}},$
$f_{2}:=\underbrace{{{\delta}_{2}}({{a}_{k_{1}}},{{a}_{k_{2}}}){{\delta}_{1}}({{a}_{k_{3}}}){{\delta}_{1}}({{a}_{k_{4}}})+\text{permutations}}_{\text{6 terms}},$
$f_{3}:=\delta_{4}({a}_{k_{1}},{a}_{k_{2}},{a}_{k_{3}},{a}_{k_{4}}),$
where the definition of $\delta_{2}$ and $\delta_{4}$ was given below Eq. (13). The above functions $f_{1},f_{2},$ and $f_{3}$ correspond to the three leading order contributions in $\alpha_{P_{A}\neq I_{A}}$ as given in Eq. (13). We suppress the indices $k_{1},\ldots,k_{4}$ in the definition of $f_{i}$’s to simplify notation. We then have the following summations:
$\sum\limits_{{{k}_{1}},{{k}_{2}},{{k}_{3}},{{k}_{4}}}{{{f}_{1}}}=3\cdot 2^{4E},\quad\sum\limits_{{{k}_{1}},{{k}_{2}},{{k}_{3}},{{k}_{4}}}{{{f}_{2}}}=6\cdot 2^{2E},\quad\sum\limits_{{{k}_{1}},{{k}_{2}},{{k}_{3}},{{k}_{4}}}{{{f}_{3}}}\leq 2^{6E}.$
Now, Eq. (20) can be readily evaluated by summing over contributions from (i) $P_{A}\neq I_{A}$, $P_{B}\neq I_{B}$; (ii) $P_{A}\neq I_{A}$, $P_{B}=I_{B}$; (iii) $P_{A}=I_{A}$, $P_{B}\neq I_{B}$; (iv) $P_{A}=I_{A}$, $P_{B}=I_{B}$, using Eq. (13). Let us consider contributions from each case.
(i) $P_{A}\neq I_{A}$ and $P_{B}\neq I_{B}$, we have:
$\sum\limits_{{{k}_{1}},{{k}_{2}},{{k}_{3}},{{k}_{4}}}{{{\alpha}_{{{P}_{A}}}}({{a}_{{{k}_{1}}}},{{a}_{{{k}_{2}}}},{{a}_{{{k}_{3}}}},{{a}_{{{k}_{4}}}}){{\beta}_{{{P}_{B}}}}({{b}_{{{k}_{1}}}},{{b}_{{{k}_{2}}}},{{b}_{{{k}_{3}}}},{{b}_{{{k}_{4}}}})}$ (21)
$={{2}^{2{|A|}+2{|B|}-4E}}\sum\limits_{{{k}_{1}},{{k}_{2}},{{k}_{3}},{{k}_{4}}}{{{f}_{1}}{{f}_{1}}}-{{2}^{{|A|}+2{|B|}-3E}}\sum\limits_{{{k}_{1}},{{k}_{2}},{{k}_{3}},{{k}_{4}}}{{{f}_{1}}{{f}_{2}}}+$
$-{{2}^{2{|A|}+{|B|}-3E}}\sum\limits_{{{k}_{1}},{{k}_{2}},{{k}_{3}},{{k}_{4}}}{{{f}_{1}}{{f}_{2}}}+{{2}^{{|A|}+{|B|}-2E}}\sum\limits_{{{k}_{1}},{{k}_{2}},{{k}_{3}},{{k}_{4}}}{{{f}_{2}}{{f}_{2}}}+$
$+{{c}_{1}}\cdot{{2}^{{|A|}+2{|B|}-5E}}\sum\limits_{{{k}_{1}},{{k}_{2}},{{k}_{3}},{{k}_{4}}}{{{f}_{1}}{{f}_{3}}}+{{c}_{1}}\cdot{{2}^{2{|A|}+{|B|}-5E}}\sum\limits_{{{k}_{1}},{{k}_{2}},{{k}_{3}},{{k}_{4}}}{{{f}_{1}}{{f}_{3}}}+$
$+{{c}_{2}}\cdot{{2}^{{|A|}+{|B|}-6E}}\sum\limits_{{{k}_{1}},{{k}_{2}},{{k}_{3}},{{k}_{4}}}{{{f}_{3}}{{f}_{3}}}+{{c}_{3}}\cdot{{2}^{{|A|}+{|B|}-4E}}\sum\limits_{{{k}_{1}},{{k}_{2}},{{k}_{3}},{{k}_{4}}}{{{f}_{2}}{{f}_{3}}}$
We now analyze the order of magnitude of each term. Since $\sum\limits_{{{k}_{1}},{{k}_{2}},{{k}_{3}},{{k}_{4}}}f_{i}\cdot(...)\leq\sum\limits_{{{k}_{1}},{{k}_{2}},{{k}_{3}},{{k}_{4}}}f_{i}$, where ()(...)( … ) represents any product of $\delta_{1},\delta_{2}$, or $\delta_{4}$ (as additional $\delta$-functions further reduce the number of unconstrained summations over $k_{1},\ldots,k_{4}$), we have
$\sum_{k_{1},k_{2},k_{3},k_{4}}f_{i}^{2}=\sum_{k_{1},k_{2},k_{3},k_{4}}f_{i}+O(2^{-E}),\quad\sum_{k_{1},k_{2},k_{3},k_{4}}f_{i}f_{j}\sim O(2^{-E})\times{\rm min}(\sum f_{i},\sum f_{j}).$ (22)
A careful inspection of Eq. (21) reveals that, in the limit $|A|\gg 1,|B|\gg 1$, the leading-order contribution comes from the term involving $\sum f_{1}^{2}$. We then have
$\sum\limits_{{{k}_{1}},{{k}_{2}},{{k}_{3}},{{k}_{4}}}{{{\alpha}_{{{P}_{A}}}}({{a}_{{{k}_{1}}}},{{a}_{{{k}_{2}}}},{{a}_{{{k}_{3}}}},{{a}_{{{k}_{4}}}}){{\beta}_{{{P}_{B}}}}({{b}_{{{k}_{1}}}},{{b}_{{{k}_{2}}}},{{b}_{{{k}_{3}}}},{{b}_{{{k}_{4}}}})}$ (23)
$=3\cdot{{2}^{2{|A|}+2{|B|}}}\left[1+2\cdot 2^{-2E}+O(2^{-|A|-E})+O(2^{-|B|-E})\right].$
(ii) $P_{A}=I_{A}$ and $P_{B}\neq I_{B}$. Recall $\alpha_{I_{A}}(a_{k_{1}},a_{k_{2}},a_{k_{3}},a_{k_{4}})=2^{4|A|}\delta_{1}(a_{k_{1}})\delta_{1}(a_{k_{2}})\delta_{1}(a_{k_{3}})\delta_{1}(a_{k_{4}})$. We have:
$\sum\limits_{{{k}_{1}},{{k}_{2}},{{k}_{3}},{{k}_{4}}}{{{\alpha}_{{{I}_{A}}}}({{a}_{{{k}_{1}}}},{{a}_{{{k}_{2}}}},{{a}_{{{k}_{3}}}},{{a}_{{{k}_{4}}}}){{\beta}_{{{P}_{B}}}}({{b}_{{{k}_{1}}}},{{b}_{{{k}_{2}}}},{{b}_{{{k}_{3}}}},{{b}_{{{k}_{4}}}})}$ (24)
$=3\cdot 2^{4|A|+2|B|-2E}\left[1+O(2^{E-|B|})+O(2^{-E-|B|})\right]$
A similar result holds when $P_{A}\neq I_{A}$ and $P_{B}=I_{B}$, with $A$ and $B$ exchanged.
(iii) $P_{A}=I_{A}$ and $P_{B}=I_{B}$. In this case the result is trivially $2^{4|A|+4|B|}$.
After combining all these contributions and including the multiplicity factors coming from the number of Pauli strings in each case, we obtain the final expression in Eq. (4) of the main text.
A.3.3 Eq. (5)
We now present the calculation details of Eq. (5) in the main text. Let us first recall the setup we use for this case. We consider a tripartite system $A$, $B$, and $C$. Since any tripartite stabilizer state can be transformed into a set of single-qubit states, Bell-pair states, and GHZ states via local Clifford unitaries, we consider initial state consisting of $g$ GHZ states shared among $A$, $B$, and $C$; $b_{AB}$, $b_{AC}$, and $b_{BC}$ Bell pairs shared between subsystems $AB$, $AC$, and $BC$, respectively, and $f_{A}$, $f_{B}$, $f_{C}$ single-qubit states in each individual subsystem. Let us start by explicitly writing down a set of stabilizer generators for this state.
For the single-qubit states, we have
$\mathbb{S}_{A}=\langle X_{1},X_{2},X_{3},\ldots,X_{f_{A}}\rangle,$ (25)
and similarly for $\mathbb{S}_{B}$ and $\mathbb{S}_{C}$. As the notation suggests, these are also the stabilizer subgroups on each individual subsystem. For the Bell-pair states shared between subsystem $A$ and $B$, we have
$L_{AB}=\langle X_{f_{A}+1}X_{|A|+f_{B}+1},Z_{f_{A}+1}Z_{|A|+f_{B}+1},\dots,X_{f_{A}+b_{AB}}X_{|A|+f_{B}+b_{AB}},Z_{f_{A}+b_{AB}}Z_{|A|+f_{B}+b_{AB}}\rangle,$ (26)
and similarly for $L_{BC}$ and $L_{AC}$. Finally, for the GHZ state, we define
$H_{X}$ $=$ $\langle X_{|A|-g+1}X_{|A|+|B|-g+1}X_{|A|+|B|+|C|-g},\dots\rangle$
$H_{AB}$ $=$ $\langle Z_{|A|-g+1}Z_{|A|+|B|-g+1},\dots\rangle$
$H_{AC}$ $=$ $\langle Z_{|A|-g+1}Z_{|A|+|B|+|C|-g+1},\dots\rangle$
$H_{BC}$ $=$ $\langle Z_{|A|+|B|-g+1}Z_{|A|+|B|+|C|-g+1},\dots\rangle.$ (27)
The set $H_{X}$, combined with any two of $H_{AB}$, $H_{AC}$, and $H_{BC}$, generates the GHZ states shared among subsystems $A$, $B$, and $C$.
Our approach of moving forward parallels that in the bipartite case. Namely, we decompose the stabilizer group into cosets of $\mathbb{S}_{AB}$ and $\mathbb{S}_{C}$, the stabilizer subgroups of subsystem $AB$ and $C$, respectively. Then we express the initial state in terms of the coset decomposition. The stabilizer group of subsystem ${\mathbb{S}}_{AB}$ is given by
${\mathbb{S}_{AB}}=\langle\mathbb{S}_{A},\ \mathbb{S}_{B},\ L_{AB},\ H_{AB}\rangle.$ (28)
The stabilizer group of total system then can be decomposed as
${\mathbb{S}_{ABC}}=\underbrace{{{\mathbb{S}}_{AB}}\otimes{{\mathbb{S}}_{C}}\ \cup\ l_{AB}^{1}{{\mathbb{S}}_{AB}}\otimes l_{C}^{1}{{\mathbb{S}}_{C}}\ \cup\ ...}_{{{4}^{{{b}_{AC}}+{{b}_{BC}}+g}}\text{ terms}},$ (29)
where $l_{AB}^{k}$ and $l_{C}^{k}$ are logical operators on $AB$ and $C$ as before. In this case, since we are dealing with initial state with a factorized structure, it is easy to directly idenfity the logical operators. For example, $\{l_{AB}^{k}\}$ consists of Pauli operators acting on $b_{AC}+b_{BC}$ Bell pairs, as well as Pauli operators acting on $g$ GHZ states, hence $4^{b_{AC}+b_{BC}+g}$ operators in total. As before, to ensure elements across different cosets have the correct commutation relation as stabilizer group elements, we require (i) if $[l_{AB}^{i},l_{AB}^{j}]=0$, then $[l_{C}^{i},l_{C}^{j}]=0$; or (ii) if $\{l_{AB}^{i},l_{AB}^{j}\}=0$, then $\{l_{C}^{i},l_{C}^{j}\}=0$.
The average $\overline{Y}$ can be written as
$\overline{Y}=\frac{1}{{{2}^{5N}}}\sum\limits_{P\in P_{N}}{\overline{{{\text{Tr}}^{4}}(P{{U}_{AB}\otimes I_{C}}{\mathbb{S}_{ABC}}{U}_{AB}^{\dagger}\otimes I_{C})}}$ (30)
$=\frac{1}{{{2}^{5N}}}\sum_{P\in P_{N}}{\sum\limits_{k}{{\overline{{{\text{Tr}}^{4}}(P_{AB}{{U}_{AB}}l_{AB}^{k}{\mathbb{S}_{AB}}{U}_{AB}^{\dagger})}}\ {{\text{Tr}}^{4}}({{P}_{C}}{{l}_{C}^{k}}{{\mathbb{S}}_{C}})}},$
where $U_{AB}=U_{A}\otimes U_{B}$.
Now recall that $U_{AB}=U_{A}\otimes U_{B}$ in the current setup. Therefore, the term ${{\overline{{{\text{Tr}}^{4}}(P_{AB}{{U}_{AB}}l_{AB}^{k}{\mathbb{S}_{AB}}{U}_{AB}^{\dagger})}}}$ has a similar factorization as in Eq. (20). However, additional complications occur in this case, since, depending on $l_{AB}^{k}$, the set $l_{AB}^{k}\mathbb{S}_{AB}$ is further decomposed in different ways. In what follows, we use $t_{ij}$ to represent the value of $\overline{{{\text{Tr}}^{4}}({P}_{AB}{U}_{AB}l_{AB}^{k}{\mathbb{S}}_{AB}{U}_{AB}^{\dagger})}$ in each case.
(1) $l_{AB}^{k}=I_{AB}$. We can decompose ${\mathbb{S}}_{AB}$ into:
${\mathbb{S}}_{AB}=\underbrace{{\mathbb{S}}_{A}\otimes{\mathbb{S}}_{B}\ \cup\ u_{1}{\mathbb{S}}_{A}\otimes v_{1}{\mathbb{S}}_{B}\ \cup\dots}_{2^{2b_{AB}+g}\text{ terms}}$ (31)
In this case, the quantity $\overline{{{\text{Tr}}^{4}}({P}_{AB}{U}_{AB}l_{AB}^{k}{\mathbb{S}}_{AB}{U}_{AB}^{\dagger})}$ factorizes into $\alpha_{P_{A}}(u_{k_{1}},u_{k_{2}},u_{k_{3}},u_{k_{4}})$ and $\beta_{P_{B}}(v_{k_{1}},v_{k_{2}},v_{k_{3}},v_{k_{4}})$ in a similar way as Eq. (20). We can thus directly apply our previous results Eqs. (13), (21), (23), and (24), noticing that
${\rm dim}(\mathbb{S}_{A})=f_{A},\quad{\rm dim}(\mathbb{S}_{B})=f_{B},\quad\sum_{k_{1},k_{2},k_{3},k_{4}}f_{1}=3\cdot 2^{4b_{AB}+2g}.$ (32)
  • If ${{P}_{A}}\neq{{I}_{A}},{{P}_{B}}\neq{{I}_{B}}$, we have:
    $t_{11}=3\cdot{{2}^{2{{f}_{A}}+2{{f}_{B}}+4{{b}_{AB}}+2g}}\left(\begin{aligned} &1+2\cdot{{2}^{-2{{b}_{AB}}-g}}+\\ &O({{2}^{-{|A|}+{{b}_{AC}}-{{b}_{AB}}}})+O({{2}^{-{|B|}+{{b}_{BC}}-{{b}_{AB}}}})\\ \end{aligned}\right)$ (33)
  • If ${{P}_{A}}={{I}_{A}},{{P}_{B}}\neq{{I}_{B}}$, we have:
    $t_{12}=3\cdot{2}^{4{|A|}+2{{f}_{B}}}\left[1+O({{2}^{b_{AB}+b_{BC}+g-|B|}})\right]$ (34)
  • If ${{P}_{A}}\neq{{I}_{A}},{{P}_{B}}={{I}_{B}}$, we have:
    $t_{13}=3\cdot{2}^{2{{f}_{A}}+4{|B|}}\left[1+O({{2}^{b_{AB}+b_{AC}+g-|A|}})\right]$ (35)
  • If ${{P}_{A}}={{I}_{A}},{{P}_{B}}={{I}_{B}}$, we have:
    $t_{14}=2^{4|A|+4|B|}.$ (36)
(2) $l_{AB}^{k}\otimes l_{C}^{k}\in H_{AC}-I$. There are $N_{2}:=2^{g}-1$ terms, and we can decompose $l_{AB}^{k}{{\mathbb{S}}_{AB}}$ into:
$l_{AB}^{k}{{\mathbb{S}}_{AB}}=\underbrace{{\mathbb{S}}_{A}\otimes v_{1}{\mathbb{S}}_{B}+u_{2}{\mathbb{S}}_{A}\otimes{\mathbb{S}}_{B}+u_{3}{\mathbb{S}}_{A}\otimes v_{3}{\mathbb{S}}_{B}...}_{2^{2b_{AB}+g}\text{ terms}}.$ (37)
To see that the decomposition above is correct, first notice that in this case $l_{AB}^{k}$ are supported on subsystem $A$ only, so there must be a term $u_{2}\mathbb{S}_{A}\otimes\mathbb{S}_{B}$ where $\mathbb{S}_{B}$ is left untouched. Secondly, since $H_{AB}\subset\mathbb{S}_{AB}$, for $l_{AB}^{k}\otimes l_{C}^{k}\in H_{AC}-I$, it must share a common Pauli-$Z$ operator with an element in $H_{AB}$ acting on the same GHZ state spanning $A$, $B$, and $C$. Thus, there will also be a term $\mathbb{S}_{A}\otimes v_{1}\mathbb{S}_{B}$ where $\mathbb{S}_{A}$ is recovered.
We can then immediately obtain the following results depending on the operator content of the Pauli string $P$ on subsystem $A$ and $B$, respectively.
  • If ${{P}_{A}}\neq{{I}_{A}},{{P}_{B}}\neq{{I}_{B}}$, we have:
    $t_{21}=3\cdot{{2}^{2{{f}_{A}}+2{{f}_{B}}+4{{b}_{AB}}+2g}}\left(\begin{aligned} &1+2\cdot{{2}^{-2{{b}_{AB}}-g}}+\\ &O({{2}^{-{|A|}+{{b}_{AC}}-{{b}_{AB}}}})+O({{2}^{-{|B|}+{{b}_{BC}}-{{b}_{AB}}}})\\ \end{aligned}\right)$ (38)
  • If ${{P}_{A}}={{I}_{A}},{{P}_{B}}\neq{{I}_{B}}$, we have:
    $t_{22}=3\cdot{2}^{4{|A|}+2{{f}_{B}}}\left[1+O({{2}^{-b_{AB}-b_{BC}-g-|B|}})\right]$ (39)
  • If ${{P}_{A}}\neq{{I}_{A}},{{P}_{B}}={{I}_{B}}$, we have:
    $t_{23}=3\cdot{2}^{2{{f}_{A}}+4{|B|}}\left[1+O({{2}^{-b_{AB}-b_{AC}-g-|A|}})\right]$ (40)
  • If ${{P}_{A}}={{I}_{A}},{{P}_{B}}={{I}_{B}}$, we have:
    $t_{24}=0,$ (41)
    since $I_{AB}\notin l_{AB}^{k}S_{AB}$.
(3) $l_{AB}^{k}\otimes l_{C}^{k}\in\left\langle L_{BC},\ H_{BC}\right\rangle-H_{BC}$. There are $N_{3}:=2^{2b_{BC}+g}-2^{g}$ terms. The reason we must exclude $H_{BC}$ is that we have already included generators $H_{AC}$ in case (2) discussed above, which, together with $H_{AB}$ will generate $H_{BC}$ already. In other words, elements in $H_{BC}$ are already included in case (2) above, so we should avoid double counting. We can decompose $l_{AB}^{k}{{\mathbb{S}}_{AB}}$ into:
$l_{AB}^{k}{{\mathbb{S}}_{AB}}=\underbrace{{\mathbb{S}}_{A}\otimes v_{1}{\mathbb{S}}_{B}+u_{2}{\mathbb{S}}_{A}\otimes v_{2}{\mathbb{S}}_{B}+u_{3}{\mathbb{S}}_{A}\otimes v_{3}{\mathbb{S}}_{B}...}_{2^{2b_{AB}+g}\text{ terms}},$ (42)
since $l_{AB}^{k}$ only acts on subsystem $B$. Similarly, we obtain the following results.
  • If ${{P}_{A}}\neq{{I}_{A}},{{P}_{B}}\neq{{I}_{B}}$, we have:
    $t_{31}=3\cdot{{2}^{2{{f}_{A}}+2{{f}_{B}}+4{{b}_{AB}}+2g}}\left[1+2\cdot{{2}^{-2{{b}_{AB}}-g}}+O({{2}^{-{|A|}+{{b}_{AC}}-{{b}_{AB}}}})\right]$ (43)
  • If ${{P}_{A}}={{I}_{A}},{{P}_{B}}\neq{{I}_{B}}$, we have:
    $t_{32}=3\cdot{2}^{4{|A|}+2{{f}_{B}}}\left[1+O({{2}^{-b_{AB}-b_{BC}-g-|B|}})\right]$ (44)
  • If ${{P}_{A}}\neq{{I}_{A}},{{P}_{B}}={{I}_{B}}$, we have:
    $t_{33}=0$ (45)
  • If ${{P}_{A}}={{I}_{A}},{{P}_{B}}={{I}_{B}}$, we have:
    $t_{34}=0$ (46)
(4) $l_{AB}^{k}\otimes l_{C}^{k}\in\left\langle L_{AC},\ H_{AC}\right\rangle-H_{AC}$, there are $N_{4}:=2^{2b_{AC}+g}-2^{g}$ terms. As before, since $l_{AB}^{k}\in H_{AC}$ has already been included in case (2) above, we should exclude them to avoid double counting. We can decompose $l_{AB}^{k}{{\mathbb{S}}_{AB}}$ into:
$l_{AB}^{k}{{\mathbb{S}}_{AB}}=\underbrace{{u_{1}\mathbb{S}}_{A}\otimes{\mathbb{S}}_{B}+u_{2}{\mathbb{S}}_{A}\otimes v_{2}{\mathbb{S}}_{B}+u_{3}{\mathbb{S}}_{A}\otimes v_{3}{\mathbb{S}}_{B}...}_{2^{2b_{AB}+g}\text{ terms}}$ (47)
since in this case $l_{AB}^{k}$ leaves subsystem $B$ and hence $\mathbb{S}_{B}$ untouched. Depending on the operator content of the Pauli string $P$ on subsystem $A$ and $B$, we obtain the following results.
  • If ${{P}_{A}}\neq{{I}_{A}},{{P}_{B}}\neq{{I}_{B}}$, we have:
    $t_{41}=3\cdot{{2}^{2{{f}_{A}}+2{{f}_{B}}+4{{b}_{AB}}+2g}}\left[1+2\cdot{{2}^{-2{{b}_{AB}}-g}}+O({{2}^{-{|B|}+{{b}_{BC}}-{{b}_{AB}}}})\right]$ (48)
  • If ${{P}_{A}}={{I}_{A}},{{P}_{B}}\neq{{I}_{B}}$, we have:
    $t_{42}=0$ (49)
  • If ${{P}_{A}}\neq{{I}_{A}},{{P}_{B}}={{I}_{B}}$, we have:
    $t_{43}=3\cdot{2}^{2{{f}_{A}}+4{|B|}}\left[1+O({{2}^{-b_{AB}-b_{AC}-g-|A|}})\right]$ (50)
  • If ${{P}_{A}}={{I}_{A}},{{P}_{B}}={{I}_{B}}$, we have:
    $t_{44}=0$ (51)
(5) Finally, $l_{AB}^{k}\otimes l_{C}^{k}$ does not belong to any of the four cases listed above. There are $N_{5}:=2^{2b_{AC}+2b_{BC}+2g}-2^{2b_{AC}+g}-2^{2b_{BC}+g}+2^{g}$ terms. In this case, $l_{AB}^{k}$ does not leave either of $\mathbb{S}_{A}$ and $\mathbb{S}_{B}$ invariant, and the decomposition of $l_{AB}^{k}{{\mathbb{S}}_{AB}}$ reads:
$l_{AB}^{k}{{\mathbb{S}}_{AB}}=\underbrace{{u_{1}\mathbb{S}}_{A}\otimes v_{1}{\mathbb{S}}_{B}+u_{2}{\mathbb{S}}_{A}\otimes v_{2}{\mathbb{S}}_{B}+u_{3}{\mathbb{S}}_{A}\otimes v_{3}{\mathbb{S}}_{B}...}_{2^{2b_{AB}+g}\text{ terms}}$ (52)
  • If ${{P}_{A}}\neq{{I}_{A}},{{P}_{B}}\neq{{I}_{B}}$, we have:
    $t_{51}=3\cdot{{2}^{2{{f}_{A}}+2{{f}_{B}}+4{{b}_{AB}}+2g}}\left(\begin{aligned} &1+2\cdot{{2}^{-2{{b}_{AB}}-g}}+\\ &+O({{2}^{-{|A|}-{{b}_{AB}}-{{b}_{AC}}-g}})+O({{2}^{-{|B|}-{{b}_{BC}}-{{b}_{AB}}-g}})\\ \end{aligned}\right)$ (53)
  • If ${{P}_{A}}={{I}_{A}},{{P}_{B}}\neq{{I}_{B}}$, we have:
    $t_{52}=0$ (54)
  • If ${{P}_{A}}\neq{{I}_{A}},{{P}_{B}}={{I}_{B}}$, we have:
    $t_{53}=0$ (55)
  • If ${{P}_{A}}={{I}_{A}},{{P}_{B}}={{I}_{B}}$, we have:
    $t_{54}=0$ (56)
The $\overline{Y}$ is then expressed in Eq. (57):
$2^{5|A|+5|B|+5|C|}\cdot 2^{-4|C|}\cdot 2^{-|C|+b_{AC}+b_{BC}+g}\cdot\overline{Y}$ (57)
$=t_{11}+N_{2}t_{21}+N_{3}t_{31}+N_{4}t_{41}+N_{5}t_{51}$
$+(4^{|B|}-1)\left[t_{12}+N_{2}t_{22}+N_{3}t_{32}+N_{4}t_{42}+N_{5}t_{52}\right]$
$+(4^{|A|}-1)\left[t_{13}+N_{2}t_{23}+N_{3}t_{33}+N_{4}t_{43}+N_{5}t_{53}\right]$
$+(4^{|A|}-1)(4^{|B|}-1)\left[t_{14}+N_{2}t_{24}+N_{3}t_{34}+N_{4}t_{44}+N_{5}t_{54}\right]$
In the above summation, the prefactor $2^{-|C|+b_{AC}+b_{BC}+g}$ comes from the size of each coset $|l_{C}^{k}\mathbb{S}_{C}|$. We observe that some terms may include contributions of the form $t_{41}=3\cdot{{2}^{2{{f}_{A}}+2{{f}_{B}}+4{{b}_{AB}}+2g}}[1+2\cdot{{2}^{-2{{b}_{AB}}-g}}+O({{2}^{-{|B|}+{{b}_{BC}}-{{b}_{AB}}}})]$, which raises concerns about whether these subleading terms $O({{2}^{-{|B|}+{{b}_{BC}}-{{b}_{AB}}}})$ can be neglected when the entanglement is large. However, we can verify that the number of $t_{41}$ is $N_{4}$, while the number of $t_{51}$ is $N_{5}$, and we have $N_{4}/N_{5}\approx 2^{-2b_{BC}-g}$. Therefore, the subleading terms contributes $O(2^{-|B|-b_{BC}-b_{AB}-g})$ to the final result.
Appendix B An example of the decomposition of the stabilizer group
In this section, we give a concrete example of the coset decomposition of a stabilizer group that has been widely used in this work. Consider a 5-qubit GHZ state:
$\frac{1}{\sqrt{2}}\left(|00000\rangle+|11111\rangle\right)$ (58)
The elements of the corresponding stabilizer group, denoted as $\mathbb{S}_{AB}$, are listed in Table 1. We partition the system into two subsystems: subsystem $A$ consisting of the first three qubits and subsystem $B$ consisting of the remaining two qubits. The reduced density matrices for these subsystems are given by Eq. (59):
$\rho_{A}=\frac{1}{2^{3}}({III}_{A}+{IZZ}_{A}+{ZIZ}_{A}+{ZZI}_{A}),\quad\rho_{B}=\frac{1}{2^{2}}({II}_{B}+{ZZ}_{B})$ (59)
Defining the stabilizer subgroups as
$\mathbb{S}_{A}:=\{{III}_{A},{IZZ}_{A},{ZIZ}_{A},{ZZI}_{A}\},\quad\mathbb{S}_{B}:=\{{II}_{B},{ZZ}_{B}\},$ (60)
we observe that $\mathbb{S}_{A}\otimes\mathbb{S}_{B}$ forms a subgroup of $\mathbb{S}_{AB}$. Consequently, $\mathbb{S}_{AB}$ can be decomposed into this subgroup and its corresponding cosets, as expressed in Eq. (61):
$\mathbb{S}_{AB}=\mathbb{S}_{A}\otimes\mathbb{S}_{B}\ \cup\ a_{1}\mathbb{S}_{A}\otimes b_{1}\mathbb{S}_{B}\ \cup\ \ a_{2}\mathbb{S}_{A}\otimes b_{2}\mathbb{S}_{B}\ \cup\ a_{3}\mathbb{S}_{A}\otimes b_{3}\mathbb{S}_{B}$ (61)
The states stabilized by $\mathbb{S}_{A}$ form a stabilizer code defined on subsystem $A$, specifically the code space spanned by $\{|000\rangle_{A},|111\rangle_{A}\}$. Similarly, the stabilizer code for subsystem $B$ is spanned by $\{|00\rangle_{B},|11\rangle_{B}\}$. The operators $a_{k}$ and $b_{k}$ correspond to the logical operators of the stabilizer codes on $A$ and $B$, respectively. Notably, these logical operators are combined in a manner that ensures all $a_{k}\otimes b_{k}$ commute with one another.
$\quad\quad\mathbb{S}_{A}\otimes\mathbb{S}_{B}\quad\quad$ $XXX\mathbb{S}_{A}\otimes XX\mathbb{S}_{B}$ $IIZ\mathbb{S}_{A}\otimes ZI\mathbb{S}_{B}$ $-XXY\mathbb{S}_{A}\otimes YX\mathbb{S}_{B}$
IIIII XXXXX IIZZI -XXYYX
IZZII -XYYXX IZIZI -XYXYX
ZZIII -YYXXX ZZZZI YYYYX
ZIZII -YXYXX ZIIZI -YXXYX
IIIZZ -XXXYY IIZIZ -XXYXY
IZZZZ XYYYY IZIIZ -XYXXY
ZZIZZ YYXYY ZZZIZ YYYXY
ZIZZZ YXYYY ZIIIZ -YXXXY
Table 1: Stabilizer group of 5-qubit GHZ state and its decomposition.
Appendix C Factorized unitary acting on three subregions
In this section, we consider the effect of a factorized unitary acting on all three subregions of a tripartite state: $U=U_{A}\otimes U_{B}\otimes U_{C}$. In this case, the final result reads:
$\overline{Y}=4\cdot 2^{-N}\big{[}1+3\cdot{{2}^{-2{{b}_{AB}}-2{{b}_{AC}}-2{g}}}+3\cdot{{2}^{-2{{b}_{AB}}-2{{b}_{BC}}-2{g}}}+3\cdot{{2}^{-2{{b}_{AC}}-2{{b}_{BC}}-2{g}}}$ (62)
$+\frac{3}{2}\cdot{{2}^{-2{{b}_{AB}}-2{{b}_{AC}}-2{{b}_{BC}}-2{g}}}+\frac{9}{2}\cdot{{2}^{-2{{b}_{AB}}-2{{b}_{AC}}-2{{b}_{BC}}-3{g}}}\big{]}$
where once again subleading corrections are of the order $O(2^{-|A|},2^{-|B|},2^{-|C|})$. Similarly to the bipartite result Eq. (4) in the main text, we find that the resulting state also achieves maximal amount of magic, provided that the subsystems are sufficiently entangled. Interestingly, this does require that the $A$, $B$, and $C$ are pairwise entangled. As Eq. (62) suggests, if $A\&B$, $B\&C$ are highly entangled while $A\&C$ are not entangled (i.e. $b_{AB},b_{BC}$ is large, but $b_{AC}$ is zero), the average $\overline{Y}$ still approaches that of a Haar random state.
As a consistency check, we consider a special case of Eq. (62), namely, when only $A\&B$ are entangled, and $C$ is not entangled with $AB$. In this case, we can check that the average $\overline{Y}$ coincides with a product of Eq. (4) in the main text for the bipartite case and a single Haar random state on $C$, which is consistent with our previous results.
We now detail the calculation of Eq. (62). Our strategy parallels that used for the bipartite case. We first decompose the stabilizer group for the system $\mathbb{S}_{ABC}$ as
${\mathbb{S}_{ABC}}=\underbrace{{{\mathbb{S}}_{A}}\otimes{{\mathbb{S}}_{B}}\otimes{{\mathbb{S}}_{C}}\ \cup\ {{a}_{1}}{{\mathbb{S}}_{A}}\otimes{{b}_{1}}{{\mathbb{S}}_{B}}\otimes{{c}_{1}}{{\mathbb{S}}_{C}}\ \cup\ \dots}_{d:={2}^{2b_{AB}+2b_{AC}+2b_{BC}+3g}\ {\rm terms}}$ (63)
where $a_{k}\otimes b_{k}\otimes c_{k}\in\langle L_{AB},L_{BC},L_{AC},H_{X},H_{AB},H_{BC}\rangle$. Notice, however, since our initial state now factorizes into single-qubit, Bell pair and GHZ states, the logical operators $\{a_{k},b_{k},c_{k}\}$ have more structures. For example, the $2^{2b_{AB}+g}$ elements in $\langle L_{AB},H_{AB}\rangle$ are only supported on subsystems $A$ and $B$, hence there are $d_{C}:=2^{2b_{AB}+g}$ terms in the coset decomposition (63) that have the same $c_{k}$’s. Similarly, there are $d_{A}:=2^{2b_{BC}+g}$ cosets with the same $a_{k}$’s, and $d_{B}:=2^{2b_{AC}+g}$ cosets with the same $b_{k}$’s.
After applying $U:=U_{A}\otimes U_{B}\otimes U_{C}$ and averaging over Haar random unitaries, we obtain:
$\overline{Y}=\frac{1}{{2^{5N}}}\sum\limits_{P\in P_{N}}\sum\limits_{k_{1},k_{2},k_{3},k_{4}}\alpha_{P_{A}}(a_{k_{1}},a_{k_{2}},a_{k_{3}},a_{k_{4}})\beta_{P_{B}}(b_{k_{1}},b_{k_{2}},b_{k_{3}},b_{k_{4}})\gamma_{P_{C}}(c_{k_{1}},c_{k_{2}},c_{k_{3}},c_{k_{4}}).$ (64)
Now, depending on the operator contents of the Pauli string $P$ on subsystems $A$, $B$ and $C$ (identity or not), we use Eq. (12) and Eq. (13) to evaluate the above expression. We give one example here. Consider when $P_{A}\neq I_{A}$, $P_{B}\neq I_{B}$, and $P_{C}\neq I_{C}$, using Eq. (12) and Eq. (13), we expand Eq. (64) as follows:
$\overline{{{{\text{Tr}}^{4}}}(PU{\mathbb{S}_{ABC}}{{U}^{\dagger}})}$ (65)
$=2^{2f_{A}+2f_{B}+2f_{C}}\left(3d^{2}+6dd_{A}+6dd_{B}+6dd_{C}+6d\cdot 2^{g}\right)$
$\quad+O(2^{3f_{A}+2f_{B}+2f_{C}-2|A|}d^{2})+O(2^{2f_{A}+3f_{B}+2f_{C}-2|B|}d^{2})$
$\quad+O(2^{2f_{A}+2f_{B}+3f_{C}-2|C|}d^{2})+O(2^{3f_{A}+3f_{B}+3f_{C}-2|A|-2|B|-2|C|}d^{3})$
$=3\cdot 2^{2f_{A}+2f_{B}+2f_{C}}d\left(d+2d_{A}+2d_{B}+2d_{C}+2\cdot 2^{g}+O(2^{-|A|})+O(2^{-|B|})+O(2^{-|C|})\right),$
where $d:=2^{2b_{AB}+2b_{AC}+2b_{BC}+3g}$ is the total number of cosets. The first line can be obtained when only leading order terms, namely, the first line in Eq. (13) is considered. Then we need to calculate the following summations over $\{k_{1},k_{2},k_{3},k_{4}\}$, constrained by products of $\delta$-functions:
$\sum\limits_{k_{1},k_{2},k_{3},k_{4}}\delta_{2}(a_{k_{1}},a_{k_{2}})\delta_{2}(a_{k_{3}},a_{k_{4}})\delta_{2}(b_{k_{1}},b_{k_{2}})\delta_{2}(b_{k_{3}},b_{k_{4}})\delta_{2}(c_{k_{1}},c_{k_{2}})\delta_{2}(c_{k_{3}},c_{k_{4}})=d^{2}$ (66)
$\sum\limits_{k_{1},k_{2},k_{3},k_{4}}\delta_{2}(a_{k_{1}},a_{k_{2}})\delta_{2}(a_{k_{3}},a_{k_{4}})\delta_{2}(b_{k_{1}},b_{k_{2}})\delta_{2}(b_{k_{3}},b_{k_{4}})\delta_{2}(c_{k_{1}},c_{k_{3}})\delta_{2}(c_{k_{2}},c_{k_{4}})=dd_{C}$ (67)
$\sum\limits_{k_{1},k_{2},k_{3},k_{4}}\delta_{2}(a_{k_{1}},a_{k_{2}})\delta_{2}(a_{k_{3}},a_{k_{4}})\delta_{2}(b_{k_{1}},b_{k_{3}})\delta_{2}(b_{k_{2}},b_{k_{4}})\delta_{2}(c_{k_{1}},c_{k_{4}})\delta_{2}(c_{k_{2}},c_{k_{3}})=2^{g}d$ (68)
This is where we need to be careful with the number of cosets having identical $a_{k}$’s, $b_{k}$’s, or $c_{k}$’s. The first summation is straightforward, if we fix an arbitrary $a_{k_{1}}\otimes b_{k_{1}}\otimes c_{k_{1}}$, then we find no more constraints except for $a_{k_{2}}\otimes b_{k_{2}}\otimes c_{k_{2}}=a_{k_{1}}\otimes b_{k_{1}}\otimes c_{k_{1}}$ and $a_{k_{3}}\otimes b_{k_{3}}\otimes c_{k_{3}}=a_{k_{4}}\otimes b_{k_{4}}\otimes c_{k_{4}}$, so the summation over $k_{1}$ and $k_{3}$ can be chosen independently, leading to $d^{2}$.
For the second summation, because of the factors $\delta_{2}(a_{k_{1}},a_{k_{2}}),\delta_{2}(b_{k_{1}},b_{k_{2}})$, we have $a_{k_{1}}\otimes b_{k_{1}}\otimes c_{k_{1}}=a_{k_{2}}\otimes b_{k_{2}}\otimes c_{k_{2}}$, similarly we have $a_{k_{3}}\otimes b_{k_{3}}\otimes c_{k_{3}}=a_{k_{4}}\otimes b_{k_{4}}\otimes c_{k_{4}}$. And considering further the factor $\delta_{2}(c_{k_{1}},c_{k_{3}})$, we find once we fix $a_{k_{1}}\otimes b_{k_{1}}\otimes c_{k_{1}}$, there are only $d_{C}=2^{2b_{AB}+g}$ choices for $a_{k_{3}}\otimes b_{k_{3}}\otimes c_{k_{3}}$, thus the summation is $dd_{C}$.
For the third summation, once we fix $a_{k_{1}}\otimes b_{k_{1}}\otimes c_{k_{1}}$, because of the factor $\delta_{2}(a_{k_{1}},a_{k_{2}})$, we need $a_{k_{2}}\otimes b_{k_{2}}\otimes c_{k_{2}}\in a_{k_{1}}\otimes b_{k_{1}}\otimes c_{k_{1}}\cdot\left\langle L_{BC},H_{BC}\right\rangle$; because of the factor $\delta_{2}(b_{k_{1}},b_{k_{3}})$, we can choose $a_{k_{3}}\otimes b_{k_{3}}\otimes c_{k_{3}}\in a_{k_{1}}\otimes b_{k_{1}}\otimes c_{k_{1}}\cdot\left\langle L_{AC},H_{AC}\right\rangle$; considering further the factor $\delta_{2}(c_{k_{2}},c_{k_{3}})$, we can only choose $a_{k_{2}}\otimes b_{k_{2}}\otimes c_{k_{2}}\in a_{k_{1}}\otimes b_{k_{1}}\otimes c_{k_{1}}\cdot\left\langle H_{BC}\right\rangle$. Furthermore, once $a_{k_{2}}\otimes b_{k_{2}}\otimes c_{k_{2}}$ is fixed, $a_{k_{3}}\otimes b_{k_{3}}\otimes c_{k_{3}}$ is also determined, because $\delta_{2}(c_{k_{2}},c_{k_{3}})$ implies that $a_{k_{3}}\otimes b_{k_{3}}\otimes c_{k_{3}}$ can only be obtained by acting on $a_{k_{1}}\otimes b_{k_{1}}\otimes c_{k_{1}}$ with stabilizer group element of the same GHZ state as in $a_{k_{2}}\otimes b_{k_{2}}\otimes c_{k_{2}}$. Finally, $a_{k_{4}}\otimes b_{k_{4}}\otimes c_{k_{4}}$ is also fully fixed by the $\delta$-functions. Therefore, there are only two independent summations in the choices for ${k_{1}}$ and the GHZ state in $H_{BC}$, leading to the final result $2^{g}d$.
After similar analysis, we obtain:
$\overline{{{\text{Tr}}^{\text{4}}}(PU{\mathbb{S}_{ABC}}{{U}^{\dagger}})}$ (69)
$=$ $\begin{cases}3\cdot{{\text{2}}^{\text{2}{{f}_{A}}+\text{2}{{f}_{B}}+\text{2}{{f}_{C}}}}d({d}+2{{d}_{A}}+2{{d}_{B}}+2{{d}_{C}}+2\cdot{{2}^{g}}),&{{P}_{A}}\neq{{I}_{A}},{{P}_{B}}\neq{{I}_{B}},{{P}_{C}}\neq{{I}_{C}}\\ {{\text{2}}^{4{|A|}+\text{2}{{f}_{B}}+\text{2}{{f}_{C}}}}(3{d}_{A}^{2}+6{{d}_{A}}),&{{P}_{A}}={{I}_{A}},{{P}_{B}}\neq{{I}_{B}},{{P}_{C}}\neq{{I}_{C}}\\ 3\cdot{{\text{2}}^{4{|A|}+4{|B|}+\text{2}{{f}_{C}}}},&{{P}_{A}}={{I}_{A}},{{P}_{B}}={{I}_{B}},{{P}_{C}}\neq{{I}_{C}}\\ {{\text{2}}^{4{|A|}+4{|B|}+4{|C|}}},&{{P}_{A}}={{I}_{A}},{{P}_{B}}={{I}_{B}},{{P}_{C}}={{I}_{C}}\\ \end{cases}.$
The subleading corrections are of the order $O(2^{-|A|},2^{-|B|},2^{-|C|})$ in each line. Cases not listed in Eq. (69) are similar. After combining all terms together, we obtain Eq. (62).
Appendix D Magic injection with shallow-depth brickwork circuits
In this section, we replace the Haar random unitary gate acting on $A$ with a shallow-depth brickwork circuit. First, we still consider the initial state shown in Fig. 1(d) of the main text, followed by one layer of local gates acting on subregion $A$. Since the initial state factorizes into single-qubit states and Bell-pair states, the average $\overline{Y}$ also factorizes into a product of contributions from single-qubit and Bell pair states. Thus, it is easy to see that, $\overline{Y}\propto 2^{-E}$ still holds true in this case.
However, if $U_{A}$ is not Haar random, the specific choice of initial state in Fig. 1(d) is no longer the most generic one. To construct the most general stabilizer initial state with tunable entanglement, notice that since any bipartite stabilizer state is local-unitary equivalent to the state in Fig. 1(d), we can simply recover a generic stabilizer state by acting on Fig. 1(d) with random Clifford unitaries $U^{\rm Cliff}_{A}\otimes U^{\rm Cliff}_{B}$. Since $U^{\rm Cliff}_{B}$ commutes with subsequent magic injection gates supported on $A$ and leaves $Y$ invariant, we can simply apply $U^{\rm Cliff}_{A}$. We then consider brickwork circuits of varying depths and gate ranges, as summarized in Fig. 6. We then numerically compute the average $\overline{Y}$ corresponding to each circuit architecture, and the results are shown in Fig. 7. We find that the qualitative picture of our analytical result is still valid for the brickwork circuit, in that the average $\overline{Y}$ decays exponentially with $E$. However, notice that for extremely shallow depths (single layer and two layers), the curve bends upward and saturates, which indicates magic injection that is below capacity. As the circuit depth increases, the behavior of $\overline{Y}$ approaches that of a Haar random $U_{A}$.
Figure 6: The circuits used to simulate the case where the gate acting on A is not sampled from a global Haar ensemble. The blue lines represent Bell pairs. The blue bricks represent unrelated random Clifford gates, and the yellow bricks represent unrelated random magical gates.
Figure 7: Numerical results for $\overline{Y}$ under brickwork circuits of varying depths and gate ranges, as depicted in Fig. 6.
Appendix E Exact results
In this section, we present an alternative analytical method to derive exact expressions of $\overline{Y}$ in all cases discussed in the main text [^1] We acknowledge Alioscia Hamma for providing insightful notes, which served as the primary inspiration for the analytical method presented in this section.. Recall the linear stabilizer entropy for a pure state $\rho=|\psi\rangle\langle\psi|$ is defined as $Y(\rho):=D^{-1}\sum_{P\in P_{N}}{\rm tr}(P\rho)^{4}$. where $D:={\rm dim}\ \mathcal{H}$. This quantity admits an equivalent expression:
$Y(\rho)=D{\rm tr}(Q\rho^{\otimes 4}),$ (70)
where $Q:=D^{-2}\sum_{P\in P_{N}}P^{\otimes 4}$. For a Hilbert space $\mathcal{H}=\otimes^{n}_{i=1}\mathcal{H}_{i}$, the operator $Q$ factories as $Q=\otimes^{n}_{i=1}Q_{i}$. We can then interpret the map $M(\cdot):=D{\rm tr}(Q(\cdot))$ as a linear operator, which facilitates subsequent calculations.
E.1 Eq. (1)
In this subsection, we present a detailed derivation of the exact form of Eq. (1). The initial state is given by:
$|\psi\rangle=\sum_{i\in\{0,1\}^{\otimes E}}2^{-E/2}{|0\rangle}^{\otimes f_{A}}{|i\rangle}_{A}\otimes{|i\rangle}_{B}.$ (71)
Let $\vec{i}:=(i_{1},i_{2},i_{3},i_{4})$, where each $i_{k}\in\{0,1\}^{\otimes E}$, denote a tuple of four bit strings. Then we have:
${|\psi\rangle\langle\psi|}^{\otimes 4}=2^{-4E}\sum_{\vec{i},\vec{i^{\prime}}}{|0\rangle\langle 0|}^{\otimes 4f_{A}}\otimes{|\vec{i}\rangle\langle{\vec{i^{\prime}}}|}_{A}\otimes{|\vec{i}\rangle\langle\vec{i^{\prime}}|}_{B}.$ (72)
Substituting this into the expression for the quantity in Eq. (1), we have:
$\overline{Y}=\mathbb{E}_{U_{A}}\ Y\left(U_{A}|\psi\rangle\langle\psi|\ U_{A}^{\dagger}\right)=D{\rm tr}\left[Q\mathbb{E}_{U_{A}}\left({U_{A}}^{\otimes 4}\otimes{I_{B}}^{\otimes 4}{|\psi\rangle\langle\psi|}^{\otimes 4}\ {U_{A}^{\dagger}}^{\otimes 4}\otimes{I_{B}}^{\otimes 4}\right)\right].$ (73)
Notice the Haar average of an operator $O\in\mathcal{B}(\mathcal{H}^{\otimes k})$ reads:
$\mathbb{E}_{U}U^{\otimes k}O{U^{\dagger}}^{\otimes k}=\sum_{\sigma,\pi\in S_{k}}V_{\sigma,\pi}{\rm tr}(T_{\sigma}O)T_{\pi},$ (74)
with $V_{\sigma,\pi}$ being the Weingarten function and $T_{\sigma},T_{\pi}$ being the aforementioned permutation operators. Applying this to Eq. (73), we have:
$\overline{Y}=\sum_{\sigma,\pi}\sum_{\vec{i},\vec{i^{\prime}}}2^{-4E}DV^{A}_{\sigma,\pi}{\rm tr}\left(Q_{A}T^{A}_{\pi}\right){\rm tr}\left(T^{A-f_{A}}_{\sigma}{|\vec{i}\rangle\langle{\vec{i^{\prime}}}|}_{A}\right){\rm tr}\left(Q_{B}{|\vec{i}\rangle\langle\vec{i^{\prime}}|}_{B}\right).$ (75)
We have used the fact that $Q=Q_{A}\otimes Q_{B}$ and $T^{A}_{\sigma}=T^{f_{A}}_{\sigma}\otimes T^{A-f_{A}}_{\sigma}$. Furthermore, because of the tensor product structure, we have:
$F_{1}(\sigma,E)$ ::=:= $\sum_{\vec{i},\vec{i^{\prime}}\in{\{0,1\}}^{\otimes 4E}}{\rm tr}\left(T^{A-f_{A}}_{\sigma}{|\vec{i}\rangle\langle{\vec{i^{\prime}}}|}_{A}\right){\rm tr}\left(Q_{B}{|\vec{i}\rangle\langle\vec{i^{\prime}}|}_{B}\right)$
$=$ $\sum_{\vec{i}\in{\{0,1\}}^{\otimes 4E}}{\rm tr}\left(Q_{B}{|\vec{i}\rangle\langle\sigma(\vec{i})|}_{B}\right)=\left[\sum_{\vec{j}\in{\{0,1\}}^{\otimes 4}}{\rm tr}\left(Q_{1}{|\vec{j}\rangle\langle\sigma(\vec{j})|}_{1}\right)\right]^{E},$
where $\sigma(\vec{i}):=(i_{\sigma(1)},i_{\sigma(2)},i_{\sigma(3)},i_{\sigma(4)})$. Then the Eq. (75) can be written as the formula shown in Eq. (77).
$\overline{Y}=\sum_{\sigma,\pi}2^{-4E}DV^{A}_{\sigma,\pi}{\rm tr}\left(Q_{A}T^{A}_{\pi}\right)F_{1}(\sigma,E),$ (77)
where all the terms represent rational fractions, and the entire summation can be completed using symbolic computation. Let $D_{A}:=2^{|A|},D_{E}:=2^{E}$. Finally, we have the exact expression Eq. (78).
$\overline{Y}=\frac{4(D_{A}^{2}D_{E}^{2}-3D_{A}D_{E}-6D_{E}^{2}+6)}{D_{A}(D_{A}^{2}-9)D_{E}^{3}}.$ (78)
E.2 Eq. (4)
In this subsection, we present the analytical method to derive the exact expression of Eq. (LABEL:eq:UaUb). The initial state is shown in Eq. (79).
$|\psi\rangle=\sum_{i\in\{0,1\}^{\otimes E}}2^{-E/2}{|0\rangle}^{\otimes f_{A}}{|i\rangle}_{A}\otimes{|0\rangle}^{\otimes f_{B}}{|i\rangle}_{B},$ (79)
and the tensor product of its density matrix reads:
${|\psi\rangle\langle\psi|}^{\otimes 4}=2^{-4E}\sum_{\vec{i},\vec{i^{\prime}}}{|0\rangle\langle 0|}^{\otimes 4f_{A}}\otimes{|\vec{i}\rangle\langle{\vec{i^{\prime}}}|}_{A}\otimes{|0\rangle\langle 0|}^{\otimes 4f_{B}}\otimes{|\vec{i}\rangle\langle\vec{i^{\prime}}|}_{B}.$ (80)
The quantity in Eq. (LABEL:eq:UaUb) then can be written as:
$\overline{Y}=\mathbb{E}_{U_{A}}\mathbb{E}_{U_{B}}\ Y\left(U_{A}U_{B}|\psi\rangle\langle\psi|\ U_{A}^{\dagger}U_{B}^{\dagger}\right)=D{\rm tr}\left[Q\mathbb{E}_{U_{A}}\mathbb{E}_{U_{B}}\left({U_{A}}^{\otimes 4}\otimes{U_{B}}^{\otimes 4}{|\psi\rangle\langle\psi|}^{\otimes 4}\ {U_{A}^{\dagger}}^{\otimes 4}\otimes{U_{B}^{\dagger}}^{\otimes 4}\right)\right].$ (81)
Applying Eq. (74), we have:
$\overline{Y}=\sum_{\sigma_{A},\pi_{A},\sigma_{B},\pi_{B}}\sum_{\vec{i},\vec{i^{\prime}}}2^{-4E}DV^{A}_{\sigma_{A},\pi_{A}}V^{B}_{\sigma_{B},\pi_{B}}{\rm tr}\left(Q_{A}T^{A}_{\pi_{A}}\right){\rm tr}\left(Q_{B}T^{B}_{\pi_{B}}\right){\rm tr}\left(T^{A-f_{A}}_{\sigma_{A}}{|\vec{i}\rangle\langle{\vec{i^{\prime}}}|}_{A}\right){\rm tr}\left(T^{B-f_{B}}_{\sigma_{B}}{|\vec{i}\rangle\langle{\vec{i^{\prime}}}|}_{B}\right).$ (82)
The summation over $\vec{i},\vec{i^{\prime}}$ can be done similarly in Eq. (E.1):
$F_{2}(\sigma_{A},\sigma_{B},E)$ ::=:= $\sum_{\vec{i},\vec{i^{\prime}}\in{\{0,1\}}^{\otimes 4E}}{\rm tr}\left(T^{A-f_{A}}_{\sigma_{A}}{|\vec{i}\rangle\langle{\vec{i^{\prime}}}|}_{A}\right){\rm tr}\left(T^{B-f_{B}}_{\sigma_{B}}{|\vec{i}\rangle\langle{\vec{i^{\prime}}}|}_{B}\right)$
$=$ $\sum_{\vec{i}\in{\{0,1\}}^{\otimes 4E}}\delta_{\sigma_{A}(\vec{i}),\sigma_{B}(\vec{i})}=\left(\sum_{\vec{j}\in{\{0,1\}}^{\otimes 4}}\delta_{\sigma_{A}(\vec{j}),\sigma_{B}(\vec{j})}\right)^{E}.$
Then we derive a summation formula Eq. (84), whose terms contains rational fractions, so the entire summation can be completed using symbolic computation.
$\overline{Y}=\sum_{\sigma_{A},\pi_{A},\sigma_{B},\pi_{B}}2^{-4E}DV^{A}_{\sigma_{A},\pi_{A}}V^{B}_{\sigma_{B},\pi_{B}}{\rm tr}\left(Q_{A}T^{A}_{\pi_{A}}\right){\rm tr}\left(Q_{B}T^{B}_{\pi_{B}}\right)F_{2}(\sigma_{A},\sigma_{B},E).$ (84)
Let $D_{A}:=2^{|A|},D_{B}:=2^{|B|},D_{E}:=2^{E}$, we have:
$\overline{Y}=\frac{4}{D_{A}D_{B}(D_{A}^{2}-9)(D_{B}^{2}-9)D_{E}^{3}}\big{(}D_{A}^{2}D_{B}^{2}D_{E}(D_{E}^{2}+3)-6D_{A}D_{B}(D_{A}+D_{B})(D_{E}^{2}+1)$ (85)
$-6(D_{A}^{2}+D_{B}^{2}-9)D_{E}(D_{E}^{2}-1)+3D_{A}D_{B}D_{E}(D_{E}^{2}+11)\big{)}.$
E.3 Eq. (5)
In this subsection, we present the analytical method to derive the exact expression of Eq. (5). The initial state is shown in Eq. (86).
$|\psi\rangle=2^{-(b_{AB}+b_{AC}+b_{BC}+g)/2}\sum_{i_{AB},i_{AC},i_{BC},i_{g}}{|0\rangle}^{\otimes f_{A}}{|i_{AB}\rangle}{|i_{AC}\rangle}{|i_{g}\rangle}_{A}\otimes{|0\rangle}^{\otimes f_{B}}{|i_{AB}\rangle}{|i_{BC}\rangle}{|i_{g}\rangle}_{B}\otimes{|i_{AC}\rangle}{|i_{BC}\rangle}{|i_{g}\rangle}_{C}.$ (86)
The tensor product of its density matrix reads:
${|\psi\rangle\langle\psi|}^{\otimes 4}=2^{-4(b_{AB}+b_{AC}+b_{BC}+g)}\sum_{\vec{i}_{AB},\vec{i}^{\prime}_{AB},\vec{i}_{AC},\vec{i}^{\prime}_{AC},\vec{i}_{BC},\vec{i}^{\prime}_{BC},\vec{i}_{g},\vec{i}^{\prime}_{g}}{{|0\rangle\langle 0|}^{\otimes 4f_{A}}{|\vec{i}_{AB}\rangle\langle\vec{i}^{\prime}_{AB}|\otimes}{|\vec{i}_{AC}\rangle\langle\vec{i}^{\prime}_{AC}|}\otimes|\vec{i}_{g}\rangle\langle\vec{i}^{\prime}_{g}|}_{A}$ (87)
$\otimes{{|0\rangle\langle 0|}^{\otimes 4f_{B}}{|\vec{i}_{AB}\rangle\langle\vec{i}^{\prime}_{AB}|\otimes}{|\vec{i}_{BC}\rangle\langle\vec{i}^{\prime}_{BC}|}\otimes|\vec{i}_{g}\rangle\langle\vec{i}^{\prime}_{g}|}_{B}\otimes{{|\vec{i}_{AC}\rangle\langle\vec{i}^{\prime}_{AC}|\otimes}{|\vec{i}_{BC}\rangle\langle\vec{i}^{\prime}_{BC}|}\otimes|\vec{i}_{g}\rangle\langle\vec{i}^{\prime}_{g}|}_{C}.$
The quantity in Eq. (5) then can be written as:
$\overline{Y}$ $=$ $\mathbb{E}_{U_{A}}\mathbb{E}_{U_{B}}\ Y\left(U_{A}U_{B}|\psi\rangle\langle\psi|\ U_{A}^{\dagger}U_{B}^{\dagger}\right)$
$=$ $D{\rm tr}\left[Q\mathbb{E}_{U_{A}}\mathbb{E}_{U_{B}}\left({U_{A}}^{\otimes 4}\otimes{U_{B}}^{\otimes 4}\otimes{I_{B}}^{\otimes 4}{|\psi\rangle\langle\psi|}^{\otimes 4}\ {U_{A}^{\dagger}}^{\otimes 4}\otimes{U_{B}^{\dagger}}^{\otimes 4}\otimes{I_{C}}^{\otimes 4}\right)\right].$
Applying Eq. (74), we have:
$\overline{Y}=2^{-4(b_{AB}+b_{AC}+b_{BC}+g)}D\sum_{\sigma_{A},\pi_{A},\sigma_{B},\pi_{B}}\sum_{\vec{i}_{AB},\vec{i}^{\prime}_{AB},\vec{i}_{AC},\vec{i}^{\prime}_{AC},\vec{i}_{BC},\vec{i}^{\prime}_{BC},\vec{i}_{g},\vec{i}^{\prime}_{g}}V^{A}_{\sigma_{A},\pi_{A}}V^{B}_{\sigma_{B},\pi_{B}}\times$ (89)
${\rm tr}\left(T^{A-f_{A}}_{\sigma_{A}}{{|\vec{i}_{AB}\rangle\langle\vec{i}^{\prime}_{AB}|\otimes}{|\vec{i}_{AC}\rangle\langle\vec{i}^{\prime}_{AC}|}\otimes|\vec{i}_{g}\rangle\langle\vec{i}^{\prime}_{g}|}_{A}\right){\rm tr}\left(T^{B-f_{B}}_{\sigma_{B}}{{|\vec{i}_{AB}\rangle\langle\vec{i}^{\prime}_{AB}|\otimes}{|\vec{i}_{BC}\rangle\langle\vec{i}^{\prime}_{BC}|}\otimes|\vec{i}_{g}\rangle\langle\vec{i}^{\prime}_{g}|}_{B}\right)\times$
${\rm tr}\left(Q_{A}T^{A}_{\pi_{A}}\right){\rm tr}\left(Q_{B}T^{B}_{\pi_{B}}\right){\rm tr}\left(Q_{C}{{|\vec{i}_{AC}\rangle\langle\vec{i}^{\prime}_{AC}|\otimes}{|\vec{i}_{BC}\rangle\langle\vec{i}^{\prime}_{BC}|}\otimes|\vec{i}_{g}\rangle\langle\vec{i}^{\prime}_{g}|}_{C}\right).$
The summation over $\vec{i}_{AB},\vec{i}^{\prime}_{AB},\vec{i}_{AC},\vec{i}^{\prime}_{AC},\vec{i}_{BC},\vec{i}^{\prime}_{BC},\vec{i}_{g},\vec{i}^{\prime}_{g}$ can be done in the similar way as those Eq. (E.1) and Eq. (E.2).
$F_{3}(\sigma_{A},\sigma_{B},b_{AB},b_{AC},b_{BC},g)$
::=:= $\sum_{\vec{i}_{AB},\vec{i}^{\prime}_{AB},\vec{i}_{AC},\vec{i}^{\prime}_{AC},\vec{i}_{BC},\vec{i}^{\prime}_{BC},\vec{i}_{g},\vec{i}^{\prime}_{g}}{\rm tr}\left(Q_{C}{{|\vec{i}_{AC}\rangle\langle\vec{i}^{\prime}_{AC}|\otimes}{|\vec{i}_{BC}\rangle\langle\vec{i}^{\prime}_{BC}|}\otimes|\vec{i}_{g}\rangle\langle\vec{i}^{\prime}_{g}|}_{A}\right)\times$
$\times$ ${\rm tr}\left(T^{A-f_{A}}_{\sigma_{A}}{{|\vec{i}_{AB}\rangle\langle\vec{i}^{\prime}_{AB}|\otimes}{|\vec{i}_{AC}\rangle\langle\vec{i}^{\prime}_{AC}|}\otimes|\vec{i}_{g}\rangle\langle\vec{i}^{\prime}_{g}|}_{A}\right){\rm tr}\left(T^{B-f_{B}}_{\sigma_{B}}{{|\vec{i}_{AB}\rangle\langle\vec{i}^{\prime}_{AB}|\otimes}{|\vec{i}_{BC}\rangle\langle\vec{i}^{\prime}_{BC}|}\otimes|\vec{i}_{g}\rangle\langle\vec{i}^{\prime}_{g}|}_{B}\right)$
$=$ $\sum_{\vec{i}_{AB},\vec{i}_{AC},\vec{i}_{BC},\vec{i}_{g}}{\rm tr}\left(Q_{AC}|\vec{i}_{AC}\rangle\langle\sigma_{A}(\vec{i}_{AC})|\right){\rm tr}\left(Q_{BC}{|\vec{i}_{BC}\rangle\langle\sigma_{B}(\vec{i}_{BC})|}\right){\rm tr}\left(Q_{g}|\vec{i}_{g}\rangle\langle\sigma_{A}(\vec{i}_{g})|\right)\delta_{\sigma_{A}(\vec{i}_{AB}),\sigma_{B}(\vec{i}_{AB})}\delta_{\sigma_{A}(\vec{i}_{g}),\sigma_{B}(\vec{i}_{g})}$
$=$ $\left(\sum_{\vec{j}}\delta_{\sigma_{A}(\vec{j}),\sigma_{B}(\vec{j})}\right)^{b_{AB}}\left[\sum_{\vec{j}}{\rm tr}\left(Q_{1}|\vec{j}\rangle\langle\sigma_{A}(\vec{j})|\right)\right]^{b_{AC}}\left[\sum_{\vec{j}}{\rm tr}\left(Q_{1}|\vec{j}\rangle\langle\sigma_{B}(\vec{j})|\right)\right]^{b_{BC}}\left[\sum_{\vec{j}}{\rm tr}\left(Q_{1}|\vec{j}\rangle\langle\sigma_{B}(\vec{j})|\right)\delta_{\sigma_{A}(\vec{j}),\sigma_{B}(\vec{j})}\right]^{g}.$
Then we derive a summation formula Eq. (91), whose terms contains rational fractions, so the entire summation can be completed using symbolic computation.
$\overline{Y}=2^{-4(b_{AB}+b_{AC}+b_{BC}+g)}D\sum_{\sigma_{A},\pi_{A},\sigma_{B},\pi_{B}}V^{A}_{\sigma_{A},\pi_{A}}V^{B}_{\sigma_{B},\pi_{B}}{\rm tr}\left(Q_{A}T^{A}_{\pi_{A}}\right){\rm tr}\left(Q_{B}T^{B}_{\pi_{B}}\right)F_{3}(\sigma_{A},\sigma_{B},b_{AB},b_{AC},b_{BC},g).$ (91)
Let $D_{A}:=2^{|A|},D_{B}:=2^{|B|},D_{C}:=2^{|C|},D_{AB}:=2^{b_{AB}},D_{AC}:=2^{b_{AC}},D_{BC}:=2^{b_{BC}},D_{g}:=2^{g}$, we have:
$\overline{Y}$ $=$ $\frac{4}{D_{A}D_{B}(D_{A}^{2}-9)(D_{B}^{2}-9)D_{AB}^{3}D_{AC}^{3}D_{BC}^{3}D_{g}^{3}}\big{(}D_{A}^{2}D_{B}^{2}D_{AB}^{3}D_{AC}^{2}D_{BC}^{2}D_{g}^{2}+3D_{A}^{2}D_{B}^{2}D_{AB}D_{AC}^{2}D_{BC}^{2}D_{g}$
-- $6D_{A}^{2}D_{AB}^{3}D_{AC}^{2}D_{BC}^{2}D_{g}^{2}-6D_{A}^{2}D_{AB}^{2}D_{AC}^{2}D_{B}D_{BC}D_{g}-18D_{A}^{2}D_{AB}D_{AC}^{2}D_{BC}^{2}D_{g}+24D_{A}^{2}D_{AB}D_{AC}^{2}$
-- $6D_{A}^{2}D_{AC}^{2}D_{B}D_{BC}+3D_{A}D_{AB}^{3}D_{AC}D_{B}D_{BC}D_{g}-6D_{A}D_{AB}^{2}D_{AC}D_{B}^{2}D_{BC}^{2}D_{g}+36D_{A}D_{AB}^{2}D_{AC}D_{BC}^{2}D_{g}$
-- $36D_{A}D_{AB}^{2}D_{AC}+3D_{A}D_{AB}D_{AC}D_{B}D_{BC}D_{g}+30D_{A}D_{AB}D_{AC}D_{B}D_{BC}-6D_{A}D_{AC}D_{B}^{2}D_{BC}^{2}$
$+$ $36D_{A}D_{AC}D_{BC}^{2}-36D_{A}D_{AC}-6D_{AB}^{3}D_{AC}^{2}D_{B}^{2}D_{BC}^{2}D_{g}^{2}+36D_{AB}^{3}D_{AC}^{2}D_{BC}^{2}D_{g}^{2}+18D_{AB}^{3}$
$+$ $36D_{AB}^{2}D_{AC}^{2}D_{B}D_{BC}D_{g}-36D_{AB}^{2}D_{B}D_{BC}-18D_{AB}D_{AC}^{2}D_{B}^{2}D_{BC}^{2}D_{g}+108D_{AB}D_{AC}^{2}D_{BC}^{2}D_{g}$
-- $144D_{AB}D_{AC}^{2}+24D_{AB}D_{B}^{2}D_{BC}^{2}-144D_{AB}D_{BC}^{2}+126D_{AB}+36D_{AC}^{2}D_{B}D_{BC}-36D_{B}D_{BC}\big{)}.$
Appendix F Simulation method
In this section, we briefly discuss the numerical simulation method we used. We use two different methods for numerically evaluting the average $\overline{Y}$. The first one is a brute-force evaluation according to the definition of $Y$. However, this requires a direct enumeration of all $4^{N}$ Pauli strings, which can be time-consuming for larger system sizes. We thus adopt a second method, which is inspired by our analytical calculations to avoid summing over exponentially many Pauli strings. According to Eq. (17), there are only four different values for ${{\alpha}_{{{P}_{A}}}}({{a}_{k}},{{a}_{k}},{{a}_{k}},{{a}_{k}})$, depending on whether $P_{A}=I_{A}$ and $a_{k}=I_{A}$. Therefore, we only need to evaluate the corresponding $\alpha$’s to efficiently compute the average $\overline{Y}$. Similarly, there are four possible contributions from different $P$ in Eq. (20), so we can simply simulate these values to obtain an unbiased estimate for $\overline{Y}$.

### v3

@@
-Stabilizer Entanglement as a Magic Highway
+Stabilizer Entanglement Enhances Magic Injection
 Abstract
 Non-stabilizerness is a key resource for fault-tolerant quantum computation, yet its interplay with entanglement in dynamical settings remains underexplored.
-We study a well-controlled, analytically tractable setup that isolates entanglement generation from magic injection.
-We analytically and numerically demonstrate that stabilizer entanglement functions as a highway that facilitates the spreading of locally injected magic throughout the entire system.
-Specifically, for an initial stabilizer state with bipartite entanglement $E$, the total magic growth, quantified by the linear stabilizer entropy $Y$, follows $\overline{Y}\propto 2^{-|A|-E}$ under a Haar random unitary on a local subregion $A$.
-Moreover, when applying a tensor product of local Haar random unitaries, the resulting state’s global magic approaches that of a genuine Haar random state if the initial stabilizer state is sufficiently entangled by a system-size-independent amount.
-Similar results are also obtained for tripartite stabilizer entanglement.
-We further extend our analysis to non-stabilizer entanglement and magic injection via a shallow-depth brickwork circuit, and find that the qualitative picture of our conclusion remains unchanged.
+We address this by analyzing a well-controlled, analytically tractable setup, where we show that entanglement acts as a conduit that teleports magic across the system, thereby enhancing magic injection.
+Using exact calculations, we prove that when a Haar-random unitary $U_{A}$ is applied to a subsystem $A$ of an entangled stabilizer state, the total injected magic increases with the entanglement between $A$ and its complement.
+More generally, for any unitary $U_{A}$, we show that this enhancement is maximized when $A$ is maximally entangled with its complement, in which case the total injected magic is exactly given by the unitary stabilizer Rényi entropy we introduce.
+This quantity provides both a directly computable measure of unitary magic and a lower bound on the minimum number of $T$ gates required to synthesize $U_{A}$.
+We further extend our analysis to tripartite stabilizer entanglement, non-stabilizer entanglement, and magic injection via shallow-depth brickwork circuits, finding that the qualitative picture remains unchanged.
 Introduction.- Understanding what distinguishes quantum systems from classical ones has been a challenging and multi-faceted endeavor.
-Much of the impact of quantumness in physics has been focusing on the idea of quantum entanglement, which played a central role in characterizing topological order [1, 2, 3], dynamically induced phases [4], quantum chaos, and even the emergence of spacetime [5].
+Much of the study of quantumness in physics has focused on the idea of quantum entanglement, which played a central role in characterizing topological order [1, 2, 3], dynamically induced phases [4], quantum chaos, and even the emergence of spacetime [5].
 However, entanglement being a quantum correlation, is just one facet of quantumness.
 Quantum advantage, for example, is concerned with the hardness of simulating quantum systems on a classical computer.
-This facet is distinct from entanglement as many highly entangled systems prepared using Clifford operations violate Bell’s inequalities but can be efficiently simulated classically [6, 7].
+This facet is not captured by entanglement alone as many highly entangled systems prepared using Clifford operations violate Bell’s inequalities but can be efficiently simulated classically [6, 7].
 The notion of classical hardness therefore constitutes a second layer of quantumness and is intimately connected to non-stabilizerness or magic [8, 9, 10, 11, 12, 13].
 Like entanglement, magic is an important but distinct quantum resource for realizing fault-tolerant universal quantum computation [13, 14].
-Therefore, it is crucial to understand any physical consequences played by this orthogonal dimension to entanglement in quantum systems.
-The importance of non-stabilizerness has been largely overlooked in quantum many-body physics until recently, where a flurry of activities have been dedicated to study its total extent in various systems, which has implication for resource estimation in state preparation and quantum simulation [15, 16, 17, 18, 19, 20], in characterizing phase transitions [21, 22], in quantum dynamics [23, 24, 25, 26, 27], in improved quantum simulation with tensor networks [28, 29], and in quantum gravity [30, 31].
+The importance of non-stabilizerness has been largely overlooked in quantum many-body physics until recently, where a flurry of activities have been dedicated to study its total extent in various systems, which has implication for resource estimation in state preparation and quantum simulation [15, 16, 17, 18, 19, 20, 21, 22], in characterizing phase transitions [23, 24], in quantum dynamics [25, 26, 27, 28, 29, 30], in improved quantum simulation with tensor networks [31, 32, 33, 34, 35], and in quantum gravity [36, 37].
 Arguably, the most fascinating perspective comes from the interplay between magic and entanglement [38, 37, 39, 40, 41, 42, 29] as both are needed for quantum advantage.
-Recent work revealed that, unlike the spread of entanglement which takes time linear in system size in spatially local Haar random circuits, magic appears to saturate in $O(\log n)$ time [37] (whereas it spreads much more slowly in Hamiltonian systems [38]).
-In the long time limit, it is also understood that magic in these highly entangled system is similar to the scrambling/encoding of information in quantum error correcting codes where it quickly delocalizes and can only be found when one can access a large part of the system [39, 26, 40].
-This asymptotic behavior is consistent with the observation that subsystem magic quickly rises and decays to almost zero as a state thermalizes [24].
-Despite these preliminary findings, the fundamental mechanisms underlying magic dynamics, particularly the interplay between magic injection and entanglement, remain poorly understood.
-Specifically, the role of entanglement in facilitating the growth and spread of magic, and vice versa, is yet to be determined in a quantitative manner.
-In generic quantum circuits and Hamiltonian dynamics, this interplay is inherently complex, as each time step of evolution simultaneously injects magic and increases entanglement [41, 37].
-To gain deeper insights into magic dynamics, a well-controlled setup that isolates different contributing factors is essential.
-Here we study precisely one such setting and show that entanglement, or at least stabilizer entanglement, acts as a conduit for magic, allowing it to spread much faster to parts of the system that are entangled.
-This would constitute an important first step in understanding how magic dynamics is distinct and yet interlinked with various types of entanglement dynamics.
-In this work, we analytically and numerically demonstrate that, given an initial stabilizer state with bipartite entanglement $E$, the total growth of magic, as quantified by the linear stabilizer entropy $Y$ under magic injection via a Haar-random unitary acting on a local subregion $A$, is enhanced by the presence of entanglement: $\overline{Y}\propto 2^{-|A|-E}$.
-Furthermore, when a unitary of the form $U=U_{A}\otimes U_{B}$, consisting of independently Haar-random local unitaries on subsystems $A$ and $B$, is applied, the global magic of the resulting state becomes nearly indistinguishable from that of a fully Haar-random state, provided the initial state is sufficiently entangled.
-Interestingly, the amount of entanglement necessary for such indistinguishability from Haar random magic is independent of system size.
-Similar results are also obtained for tripartite stabilizer entanglement.
-We further extend our analysis to non-stabilizer entanglement and magic injection via a shallow-depth brickwork circuit, and find that the qualitative picture of our conclusion remains unchanged in these cases.
-Quantum magic and its measure.- Quantum magic quantifies the amount of “non-stabilizer” or “non-Clifford” resources in a quantum state.
-A computable measure for quantum magic, which we will primarily focus on in this work, is given by the linear stabilizer entropy.
-Denote by $P_{N}$ the set of $N$-qubit Pauli strings.
-The linear stabilizer entropy for a pure state $\rho=|\psi\rangle\langle\psi|$ is defined as $Y(\rho):=2^{-N}\sum_{P\in P_{N}}{\rm tr}(P\rho)^{4}$.
-To gain some intuition about this quantity, first consider a stabilizer pure state.
-There are precisely $2^{N}$ Pauli strings corresponding to the stabilizer group elements of $\rho$ for which ${\rm tr}(P\rho)^{4}=1$, and ${\rm tr}(P\rho)^{4}=0$ for all other Pauli strings.
-Thus $Y(\rho)=1$ for stabilizer states.
-For a generic non-stabilizer pure state, the Pauli spectrum satisfies the normalization condition: $12^{-N}\sum_{P\in P_{N}}{\rm tr}(P\rho)^{2}=$ 1.
-Therefore, $Y(\rho)$ defined above can be thought of as the second moment of the Pauli spectrum of $\rho$, which characterizes how delocalized the expectation values of $P$ are over all possible Pauli strings.
-The linear stabilizer entropy satisfies the following properties: (i) faithfulness: let $\text{STAB}_{n}$ denote the set of all $n$-qubit pure stabilizer states, then $Y(\rho)=1$ iff $\rho\in\text{STAB}_{n}$, otherwise $Y(\rho)<1$; (ii) invariant under Clifford unitaries $U_{C}$: $Y(U_{C}\rho U^{\dagger}_{C})=Y(\rho)$; (iii) multiplicative: $Y(\rho_{A}\otimes\rho_{B})=Y(\rho_{A})Y(\rho_{B})$; (iv) the closely related quantity $M^{\rm lin}:=1-Y$ is a strong pure-state monotone under free operations [42].
-Hence, $Y(\rho)$ serves as a good measure for quantum magic from the point of view of magic-state resource theory.
-Moreover, the logarithm of $Y(\rho)$ is related to the second stabilizer Rényi entropy $M_{2}(\rho)=-{\rm log}Y(\rho)$, another widely used magic measure [43].
-Figure 1: Schematics of the setup used for analytical and numerical calculations.
-The initial states are bipartite or tripartite stabilizer pure states.
+Recent work revealed that magic in highly entangled systems is similar to the scrambling/encoding of information in quantum error correcting codes where it quickly delocalizes and can only be found when one can access a large part of the system [43, 28, 44].
+This asymptotic behavior is consistent with the observation that subsystem magic quickly rises and decays to almost zero as the system thermalizes [26].
+Despite recent progress, the fundamental mechanisms governing magic dynamics, particularly the interplay between magic injection and entanglement, remain poorly understood.
+In generic quantum circuits and Hamiltonian dynamics, entanglement and magic dynamics are interwined, as each time step of evolution simultaneously injects magic and increases entanglement [45, 46].
+To disentangle their roles, it is essential to study a controlled setting where one can isolate their contributions.
+In this work, by constructing such a setup, we aim to address the following question: how does entanglement affect the injection of magic into a quantum system?
+We show that preexisting stabilizer entanglement (i.e.
+entanglement of a nonmagical, stabilizer state) can enhance magic injection via gate teleportation.
+More explicitly, using exact calculations, we prove that when a Haar-random unitary $U_{A}$ is applied to a subsystem $A$ of an entangled stabilizer state, the total injected magic increases with the entanglement between $A$ and its complement.
+More generally, for any unitary $U_{A}$, we show that this enhancement is maximized when $A$ is maximally entangled with its complement, in which case the total injected magic is exactly given by the unitary stabilizer Rényi entropy we introduce.
+This quantity provides both a directly computable measure of unitary magic and a lower bound on the minimum number of $T$ gates (defined as $T=e^{i\frac{\pi}{8}Z}$) required to synthesize $U_{A}$, also known as $T$-count.
+We further extend our analysis to tripartite stabilizer entanglement, non-stabilizer entanglement, and magic injection via shallow-depth brickwork circuits, finding that the qualitative picture remains unchanged.
+Quantum magic and its measure.- Clifford operations consist of unitaries that map Pauli operators to Pauli operators, the preparation of Pauli eigenstates, and the measurement in the Pauli basis among other things.
+Their simplicity around Pauli operators allows them to be efficiently simulable on a classical computer.
+The states that can be prepared with Clifford operations are called stabilizer states.
+A stabilizer pure state $|\psi\rangle$ is one where $s|\psi\rangle=|\psi\rangle$ for all $s\in S$ where $S$ is a size-$2^{N}$ Abelian subgroup of the Pauli group $P_{N}$ over $N$ qubits.
+It can be shown that $\rho=|\psi\rangle\langle\psi|=\frac{1}{2^{N}}\sum_{s\in S}s$ — a stabilizer pure state is an equal superposition of the $2^{N}$ stabilizer group elements.
+Quantum magic quantifies the amount of “non-Clifford” or “non-stabilizer” resources needed in a quantum state or process.
+In this work, we focus on a computable measure of magic in quantum states called linear stabilizer entropy [47] defined as
+$Y^{\rm lin}(\rho):=1-\frac{1}{2^{N}}\sum_{P\in P_{N}}{\rm tr}(P\rho)^{4}.$ (1)
+[…]
+The linear stabilizer entropy satisfies the following properties: (i) faithfulness: let $\text{STAB}_{N}$ denote the set of all $N$-qubit pure stabilizer states, then $Y^{\rm lin}(\rho)=0$ iff $\rho\in\text{STAB}_{N}$, otherwise $Y^{\rm lin}(\rho)>0$; (ii) invariant under Clifford unitaries $U_{C}$: $Y^{\rm lin}(U_{C}\rho U^{\dagger}_{C})=Y^{\rm lin}(\rho)$; (iii) monotonicity: $Y^{\rm lin}$ is a strong pure-state monotone under free operations [48].
+Hence, $Y^{\rm lin}(\rho)$ serves as a good measure for quantum magic from the point of view of magic-state resource theory, which increases and approaches unity as magic increases.
+Notice that $Y^{\rm lin}(\rho)$ is also closely related to the second stabilizer Rényi entropy (SRE) $M_{2}(\rho)=-{\rm log}(1-Y^{\rm lin}(\rho))$, another widely used magic measure [47].
+Figure 1: Schematics of the setup considered in this work.
+The initial states are bipartite stabilizer pure states.
 (a) A Haar random unitary injects magic by acting on a subregion $A$.
 (b) A factorized unitary $U=U_{A}\otimes U_{B}$, where $U_{A}$ and $U_{B}$ are independently Haar-random, acts on the initial stabilizer state.
-(c) A factorized unitary $U=U_{A}\otimes U_{B}$ injects magic by acting on subregion $AB$ of a tripartite system.
-(d) Initial state (2) for numerical simulations of the setup in (a).
-The state consists of $f_{A}$ and $f_{B}$ qubits in |+ket|$\rangle$| + ⟩ state in subsystem $A$ and $B$, respectively, along with $E$ Bell pairs shared between $A$ and $B$.
-(e) Initial state for obtaining Eq.
-(5).
-The state initial state consists of $g$ GHZ states shared among $A$, $B$, and $C$; $b_{AB}$, $b_{AC}$, and $b_{BC}$ Bell pairs shared between subsystems $AB$, $AC$, and $BC$, respectively, and $f_{A}$, $f_{B}$, $f_{C}$ single-qubit states in each individual subsystem.
-Stabilizer entanglement facilitates magic injection.-To illustrate how stabilizer entanglement can influence magic injection, we examine the effect of a single-qubit $T$ gate on two distinct stabilizer states: an unentangled product state, $|\psi_{1}\rangle=|+\rangle^{\otimes N}=\left((|0\rangle+|1\rangle)/\sqrt{2}\right)^{\otimes N}$, and a GHZ state with long-range entanglement, $|\psi_{2}\rangle=(|0\rangle^{\otimes N}+|1\rangle^{\otimes N})/\sqrt{2}$.
-In both cases, the initial states contain no magic, and the $T$ gate injects magic locally.
-However, the manner in which this injected magic manifests differs significantly between the two cases.
-In $T|\psi_{1}\rangle$, the injected magic remains locally accessible and can be directly probed at the specific qubit where the $T$ gate is applied.
-In contrast, for $T|\psi_{2}\rangle$, the state appears non-magical to any local observer with access only to a subregion of the system.
-This simple example illustrates that the propagation of locally injected magic throughout the system is intrinsically tied to the entanglement structure of the initial state.
-To further quantify how stabilizer entanglement facilitates magic spreading, we consider the setup illustrated in Fig.
-1(a), which enables an analytical calculation.
-Specifically, we begin with an initial stabilizer pure state $\rho=|\psi\rangle\langle\psi|$, partitioned into subregion $A$ and its complement $B$.
+(c) Schematic of gate teleportation.
+Because of the Bell pairs between $A$ and $B$, $U_{A}$ acting on $A$ is equivalent to $U^{\prime}_{A}$} and $U^{\prime}_{B}$} acting on $A$ and $B$.
+Stabilizer entanglement enhances magic injection.- To see how entanglement impacts magic injection, consider the setup illustrated in Fig.
+1(a).
+We take an initial stabilizer pure state $\rho$, partitioned into subregion $A$ and its complement $B$.
 A Haar random unitary $U_{A}$ is then applied to subregion $A$, and we are interested in computing the linear stabilizer entropy of the resulting state $\sigma=U_{A}|\psi\rangle\langle\psi|U_{A}^{\dagger}$, averaged over Haar random unitaries.
-We first make a few remarks on this particular choice of setup before showing the result.
-This construction is a natural generalization of the toy example discussed above, replacing the single-qubit unitary with a Haar random unitary on a finite subregion.
-While the Haar-random unitary may appear nonlocal within subregion $A$, it is local relative to the entire system if $|A|\ll|B|$.
-The use of Haar random unitaries primarily serves to render the analytical calculations more tractable.
-We will show below that key qualitative features of our results persist even when considering more structured unitaries within subregion $A$.
-Secondly, this setup clearly separates entanglement generation with respect to the bipartition and magic injection.
-The unitary acting on $A$ injects magic to the system while preserving the amount of entanglement between $A$ and $B$, including the full entanglement spectrum.
-As many, if not most, operationally convenient circuits apply Clifford gates and non-Clifford phase gates in separate stages, e.g.
-Clifford+T decomposition [13] and the PFC ensemble that prepare pseudorandom unitaries [44], our setup is to probe the internal dynamics that take place between the most elementary layers in those circuits.
-This is complementary to the random unitary circuit model used in Ref.
-[37] to study magic saturation.
-In that model, unitary gates simultaneously generate entanglement and inject magic, leading to universal late-time magic dynamics prior to saturation, regardless of the initial state.
-We analytically compute the linear stabilizer entropy $Y(U_{A}|\psi\rangle\langle\psi|U_{A}^{\dagger})$, averaged over the ensemble of Haar random unitaries $U_{A}$.
-In the limit $|A|\gg 1$, we obtain, to leading order,
-$\overline{Y}=\mathbb{E}_{U_{A}}\ Y\left(U_{A}|\psi\rangle\langle\psi|\ U_{A}^{\dagger}\right)=4\cdot{{2}^{-|A|-E}}\left[1+O({2}^{-{|A|-E}})\right],$ (1)
-where $E=-\rm{tr}(\rho_{A}\text{log}_{2}\rho_{A})$ is the amount of entanglement between $A$ and $B$, which is an integer for a stabilizer state $|\psi\rangle$.
-An exact calculation of $\overline{Y}$ for arbitrary $|A|$ and $E$ is possible, and is shown in the Supplemental Material (SM) [45].
-Eq.
-(1) establishes a quantitative relation between the efficiency of local magic injection and the amount of stabilizer entanglement contained in the target state.
-For a fixed subregion $A$, the average $Y(\rho)$ decays exponentially with the amount of entanglement between $A$ and its complement.
-Therefore, global magic injection on $AB$ is enhanced with entanglement, which acts as a throttle that controls how much magic can flow to region $B$ from $A$.
-A detailed derivation of Eq.
-(1) is provided in the SM [45], yet its qualitative features, particularly the factor $2^{-|A|-E}$ can be understood straightforwardly.
-Since the unitary acts only on subsystem $A$, stabilizers supported entirely on $B$ remains unchanged under $U_{A}$ and contribute to $Y(\rho)$.
-A simple counting shows that the number of surviving stabilizers is precisely $2^{|B|-E}$, leading to the factor $2^{-|A|-E}$ in Eq.
-(1).
-The prefactor 4 in front of $2^{-|A|-E}$ is, in some sense, non-universal.
-To see this, let us define the generalized linear stabilizer entropy $Y_{\alpha}(\rho):=2^{-N}\sum_{P\in P_{N}}{\rm tr}(P\rho)^{2\alpha}$.
-Eq.
-(1) thus corresponds to the specific case $\alpha=2$.
-Similarly, $Y_{\alpha}(\rho)$ is related to the $\alpha$-th stabilizer Rényi entropy: $M_{\alpha}(\rho)=\frac{1}{1-\alpha}{\rm log}Y_{\alpha}(\rho)$.
-In the limit $\alpha\rightarrow\infty$, only the surviving stabilizers supported on $B$ contribute, yielding $\overline{Y_{\infty}}=2^{-|A|-E}$.
-For arbitrary $\alpha$, we have $\overline{Y_{\alpha}}=[1+c_{\alpha}(|A|,E)]2^{-|A|-E}$, with the one coming from the surviving stabilizers, and $c_{\alpha}(|A|,E)$ coming from contributions of all other Pauli strings, which in general depends on both $|A|$ and $E$ [46].
-Specifically, Eq.
-(1) implies that for $\alpha=2$, we have $c_{2}=3$.
-Figure 2: (a) Numerical simulations of the setup depicted in Fig.
-1(a).
-We consider initial states of the form Eq.
-(2) with varying $|A|$ and $E$.
-We compute the average value of $\overline{Y}$ (upper triangles), as well as $2^{-\overline{M_{2}}}$ (lower triangles), where $\overline{M_{2}}$ is the average value of the second stabilizer Rényi entropy.
-The analytical result Eq.
-(1) is shown as dashed lines.
-(b) Numerical simulations of the setup depicted in Fig.
-1(b).
-Choice of the initial state is the same as in (a).
-The analytic result Eq.
-(LABEL:eq:UaUb) is shown as dashed lines.
-(c) Exact results (squares) and leading order contributions Eq.
-(5) of the setup depicted in Fig.
-1(c).
-(d) Same as in (a) but for non-stabilizer initial states with $k$ imperfect Bell pairs $|\phi_{\theta}\rangle=\cos\theta|00\rangle+\sin\theta|11\rangle$ shared between $A$ and $B$.
-We compute the magic already existing in the initial state $\overline{Y}_{\text{ini}}$ (dashed lines) and the change after the action of $U_{A}$: $\overline{Y}_{\text{final}}/\overline{Y}_{\text{ini}}$ (solid lines).
-(e) Same as in (c) but initial states are now constructed with four different types of entanglement spectra (see main text).
-Error bars are smaller than the size of the data points and hence not visible.
-(f) Same as in (d)&(e), but now plotting $\overline{Y}_{\rm final}$ in each case considered.
-To test Eq.
-(1), particularly the regime of validity of neglecting subleading corrections, we perform numerical simulations of the setup depicted in Fig.
-1(a).
-Recall that any stabilizer state $|\psi\rangle$ can be brought to the form of Eq.
-(2) by local Clifford unitaries [47]: $|\Phi\rangle=U^{\rm Cliff}_{A}\otimes U^{\rm Cliff}_{B}|\psi\rangle$ such that
-$|\Phi\rangle=|+\rangle^{\otimes f_{A}}\ |{\rm Bell}\rangle_{AB}^{\otimes E}\ |+\rangle^{\otimes f_{B}},$ (2)
-which is illustrated in Fig.
-1(d).
-The state consists of $f_{A}$ and $f_{B}$ qubits in |+ket|$\rangle$| + ⟩ state in subsystem $A$ and $B$, respectively, along with $E$ Bell pairs shared between $A$ and $B$, so that the entanglement entropy between $A$ and $B$ is $E$.
-Since $\overline{Y}$ is obtained by averaging over the Haar random ensemble, $U^{\rm Cliff}_{A}$ can be simply absorbed into the average.
-Moreover, $Y(\rho)$ is invariant under Clifford unitary $U_{B}^{\rm Cliff}$.
+This setup preserves the entanglement spectrum between $A$ and $B$ and models the process where magic is locally injected by a non-Clifford unitary gate to a stabilizer state when $|A|\ll|B|$.
+We will show in SM [49] that key qualitative features of our results also persist even when considering more structured unitaries within subregion $A$.
+We then compute Haar-averaged the linear stabilizer entropy $Y^{\rm lin}(U_{A}|\psi\rangle\langle\psi|U_{A}^{\dagger})$ exactly. To leading order,
+$\overline{Y^{\rm lin}}$ $=$ $\mathbb{E}_{U_{A}}\ Y^{\rm lin}\left(U_{A}|\psi\rangle\langle\psi|\ U_{A}^{\dagger}\right)$ (2)
+$=$ $1-4\cdot{{2}^{-|A|-E}}\left[1+O({2}^{-{|A|-E}})\right],$
+where $|A|\gg 1$ and $E=-\rm{tr}(\rho_{A}\text{log}\rho_{A})$ is the amount of entanglement between $A$ and $B$, which is an integer for a stabilizer state.
+An exact closed-form expression for $\overline{Y^{\rm lin}}$ at arbitrary $|A|$ and $E$ is derived in the Supplemental Material (SM) [49].
+It turns out that the leading-order analytical formula already provides an excellent approximation for moderate system sizes and entanglement: numerical simulations show deviations become negligible once $|A|+E\gtrsim 4$ [Fig.
+2(a)].
+Consistently, we find that the Haar-averaged second SRE $\overline{M_{2}}$ is nearly indistinguishable from $\log(1-\overline{Y^{\rm lin}})$, confirming that the average in Eq.
+(2) is typical with vanishing fluctuations [49].
 Thus we have
-$\mathbb{E}_{U_{A}}Y\left(U_{A}|\psi\rangle\langle\psi|U_{A}^{\dagger}\right)=\mathbb{E}_{U_{A}}Y\left(U_{A}|\Phi\rangle\langle\Phi|U_{A}^{\dagger}\right)$ (3)
-for an arbitrary stabilizer state $|\psi\rangle$.
-We numerically average over $Y(U_{A}|\Phi\rangle\langle\Phi|U_{A}^{\dagger})$ with Haar random $U_{A}$ for various values of $|A|$ and $E$.
-The results are shown in Fig.
-2(a), where we find excellent agreement with the analytical prediction of Eq.
-(1).
-In particular, the subleading corrections become negligible already at moderate subsystem sizes and entanglement $|A|+E\gtrsim 4$.
-We also simulate the Haar averaged second stabilizer Rényi entropy $\overline{M_{2}}$, which turns out to be extremely close to ${\rm log}\ {\overline{Y}}$.
-This implies that the average in Eq.
-(1) is also typical, as sample-to-sample fluctuations are extremely small.
-The same qualitative behavior holds if $U_{A}$ is given by a brickwork circuit instead of Haar random unitary, but the total injected magic is limited by the depth (see SM [45]).
-One can further explore the most general approach to magic injection while preserving the entanglement spectrum.
-Starting from a stabilizer state, we now apply a unitary that is a tensor product of local unitaries acting on subsytems $A$ and $B$: $U=U_{A}\otimes U_{B}$, as illustrated in Fig.
+$\overline{M_{2}}=|A|+E+2+O(2^{-|A|-E}).$ (3)
+Notice that $\overline{M_{2}}$ is extensive, as expected.
+Eqs.
+(2) and (3) establish a quantitative relation between the efficiency of local magic injection and the amount of stabilizer entanglement contained in the target state.
+For a fixed subregion $A$, $\overline{Y^{\rm lin}}(\rho)$ approaches $1$ exponentially with the amount of entanglement $E$ between $A$ and its complement, and correspondingly the average second SRE increases linearly with $E$, indicating that global magic injection on $AB$ is enhanced by preexisting entanglement.
+Similar behavior arises when $U_{A}$ is implemented by a brickwork circuit rather than a Haar-random unitary, although in that case the total injected magic is limited by the circuit depth (see SM [49]).
+Intuitively, the enhancement of magic injection by stabilizer entanglement as revealed by Eqs.
+(2) and (3) can be understood as a consequence of gate ‘teleportation’ mediated by Bell pairs spanning subsystems $A$ and $B$ [see Fig.
+1(c)].
+As a simple example, consider a $2N$-qubit maximally entangled state between $A$ and $B$: $|{\rm Bell}\rangle_{AB}^{\otimes N}$.
+Then any unitary acting on half of the Bell pairs in $A$ will teleport to its partner in $B$ since $U\otimes I\ |\mathrm{Bell}\rangle_{AB}^{\otimes N}=I\otimes U^{T}\ |\mathrm{Bell}\rangle_{AB}^{\otimes N}$.
+It has been shown that any bipartite stabilizer state $|\psi\rangle_{AB}$ can be converted into a collection of single-qubit states (say $|+\rangle$) and Bell pairs through local Clifford unitaries [50]:
+$U_{A}^{\rm Cliff}\otimes U^{\rm Cliff}_{B}|\psi\rangle_{AB}=|+\rangle^{\otimes f_{A}}|\mathrm{Bell}\rangle^{\otimes E}_{AB}|+\rangle^{\otimes f_{B}}.$ (4)
+Hence for a stabilizer state that has entanglement, one can inject magic on $B$ through $A$ where magic flows to $B$ via the “entanglement conduit” formed by the $E$ Bell pairs.
+This simple picture also offers an intuitive explanation for the observed dynamics of local magic as a system thermalizes [26]—while time evolution increases local magic initially when the spins are unentangled, this local magic quickly spreads through the conduit as entanglement also grows in the system.
+Figure 2: Numerical simulations of $\overline{Y^{\rm lin}}$ for the setup depicted in (a) Fig.
+1(a), and (b) Fig.
+1(b), respectively.
+We consider stabilizer initial states with varying system sizes and bipartite entanglement $E$.
+The analytical results Eqs.
+(2) and (LABEL:eq:UaUb) are shown as dashed lines.
+A further manifestation of entanglement-enhanced magic injection is revealed by applying a unitary that is a tensor product of local unitaries acting on subsytems $A$ and $B$: $U=U_{A}\otimes U_{B}$, as illustrated in Fig.
 1(b).
@@
 In the limit where $|A|\gg 1$ and $|B|\gg 1$, we find [49]
-$\overline{Y}$ $=$ $\mathbb{E}_{U_{A}}\mathbb{E}_{U_{B}}Y\left(U_{A}U_{B}|\psi\rangle\langle\psi|U_{A}^{\dagger}U_{B}^{\dagger}\right)$
-$=$ $4\cdot{2}^{-N}\left[1+3\cdot 2^{-2E}+O({{2}^{-|A|-E}})+O({{2}^{-|B|-E}})\right],$
+$\overline{Y^{\rm lin}}$ $=$ $\mathbb{E}_{U_{A}}\mathbb{E}_{U_{B}}Y^{\rm lin}\left(U_{A}U_{B}|\psi\rangle\langle\psi|U_{A}^{\dagger}U_{B}^{\dagger}\right)$
+$=$ $1-4\cdot{2}^{-N}\left[1+3\cdot 2^{-2E}+O({{2}^{-|A|-E}},{{2}^{-|B|-E}})\right],$
 where $N=|A|+|B|$ is the total number of qubits.
-This result reveals an intriguing mechanism by which stabilizer entanglement facilitates magic saturation.
-Notice that the average value of $\overline{Y}$ for Haar random states is given by $4\cdot 2^{-N}$.
+First of all, Eq.
+(LABEL:eq:UaUb) is consistent with Eq.
+(2) and our teleportation picture upon taking $|A|=|B|=E$.
+Furthermore, notice that the average value of $\overline{Y^{\rm lin}}$ for Haar random states is given by $1-4\cdot 2^{-N}$.
 Eq.
-(LABEL:eq:UaUb) shows that the gap between $\overline{Y}$ for a stabilizer state acted upon by $U_{A}\otimes U_{B}$ and that for a Haar random state decays exponentially with the amount of entanglement.
-In other words, despite the applied unitary being a tensor product of local unitaries on subsystems $A$ and $B$, from the perspective of magic, its effect on magic injection closely resembles that of a genuinely global Haar-random unitary, provided that the initial state is sufficiently entangled.
+(LABEL:eq:UaUb) shows that the gap between $\overline{Y^{\rm lin}}$ for a stabilizer state acted upon by $U_{A}\otimes U_{B}$ and that for a Haar random state decays exponentially with the amount of entanglement.
+Therefore, despite the applied unitary being a tensor product of local unitaries on subsystems $A$ and $B$, from the perspective of magic, its effect on magic injection closely resembles that of a genuinely global Haar-random unitary on $AB$, provided that the initial state is sufficiently entangled.
 Interestingly, the amount of entanglement necessary for such indistinguishability from Haar random magic is independent of system size.
-In Fig.
-2(b), we numerically simulate this set up using the initial state (2) and averaging over Haar random unitaries $U_{A}$ and $U_{B}$.
-Again, we find excellent agreement with the prediction of Eq.
-(LABEL:eq:UaUb), with the subleading corrections negligible for moderate subsystem sizes and entanglement.
-Moreover, the average $\overline{Y}$ saturates to that of a Haar random state already at $E\approx 3$, which is independent of system size.
-Tripartite entanglement.- The above results for bipartite systems can be naturally extended to tripartite systems.
-Consider a stabilizer state shared among three parties $A$, $B$ and $C$.
-Since any tripartite stabilizer state can be transformed via local unitaries acting on $A$, $B$, and $C$ into a collection of GHZ state, Bell pair states, and single-qubit states [48], we can, without loss of generality, consider initial states as depicted in Fig.
-1(e).
-Specifically, the initial state consists of $g$ GHZ states shared among $A$, $B$, and $C$; $b_{AB}$, $b_{AC}$, and $b_{BC}$ Bell pairs shared between subsystems $AB$, $AC$, and $BC$, respectively, and $f_{A}$, $f_{B}$, $f_{C}$ single-qubit states in each individual subsystem.
-Since magic injection via unitary operations on any single subsystem in this setting mirrors the bipartite case discussed earlier, we first consider unitaries of the form $U=U_{A}\otimes U_{B}$ [Fig.
-1(c)].
-In the limit where $|A|\gg 1$ and $|B|\gg 1$, we find that the average $\overline{Y}$ after applying $U$ is given by [45]
-$\overline{Y}$ $=$ $\mathbb{E}_{U_{A}}\mathbb{E}_{U_{B}}Y\left(U_{A}U_{B}|\psi\rangle\langle\psi|U_{A}^{\dagger}U_{B}^{\dagger}\right)$ (5)
-$=$ $4\cdot{2}^{-|A|-|B|-g-{{b}_{AC}}-{{b}_{BC}}}\left[1+3\cdot{2}^{-2{{b}_{AB}}-g}\right],$
-where subleading corrections are of order $O(2^{-|A|},2^{-|B|})$.
-Eq.
-(5) can be understood as follows.
-The average value $\overline{Y}$ is once again exponentially suppressed in both the total size of the subregion where the unitary is applied: $|A|+|B|$, and the amount of entanglement between subregion $AB$ and its complement: $g+b_{AC}+b_{BC}$.
-Hence the physical interpretation of the prefactor in Eq.
-(5) is identical to that of Eq.
-(1).
-The discrepancy between these two cases lies in the additional factor $3\cdot{2}^{-2{{b}_{AB}}-g}$, which becomes negligible once $A$ and $B$ are sufficiently entangled.
-This is precisely what we have seen in Eq.
-(LABEL:eq:UaUb): the effect of a factorized unitary $U_{A}\otimes U_{B}$ becomes essentially indistinguishable from that of a global unitary from the standpoint of magic, once $A$ and $B$ are sufficiently entangled.
-Thus, Eq.
-(5) encapsulates the combined effects of Eqs.
-(1) and (LABEL:eq:UaUb) in the tripartite setting.
-In the SM [45], we further consider a factorized unitary acting on all three subregions: $U=U_{A}\otimes U_{B}\otimes U_{C}$, where we confirm that in this case the effect of a factorized Haar random unitary on magic injection again becomes indistinguishable from that of a global Haar random unitary, provided that the subregions are sufficiently entangled with one another.
-Non-stabilizer entanglement.- So far, we have focused on the propagation of magic on stabilizer states.
+Fig.
+2(b) shows numerical simulations of Haar-averaged $U_{A}$ and $U_{B}$, which again closely match Eq.
+(LABEL:eq:UaUb) with subleading corrections already negligible at moderate $|A|$ and $E$.
+Remarkably, $\overline{Y^{\rm lin}}$ saturates to the Haar-random value once $E\approx 3$, independent of system size.
+In SM, we generalize Eqs. (2) and (LABEL:eq:UaUb) to tripartite states, where similar results are obtained [49].
+T count and unitary SRE.- The preceding results establish that entanglement enhances magic injection, but they also raise a natural question: what sets the ultimate limit of this enhancement?
+From a resource-theoretic perspective, all non-Clifford resources originate from the applied unitary $U_{A}$, so the total injected magic must be upper bounded by the non-Clifford content of $U_{A}$ itself.
+We begin by introducing the unitary stabilizer Rényi entropy of a given unitary, which provides a direct bridge between our previous results and the non-Clifford resources contained in $U$.
+Definition.
+Let $U$ be an $N$-qubit unitary, then the $\alpha$-th unitary stabilizer Rényi entropy of $U$ is defined as
+[…]
+where $P_{i}\in P_{N}$.
+To gain some intuition on the above quantity, notice that $UP_{i}U^{\dagger}=\sum c_{ij}P_{j}$ for a single Pauli string $P_{i}$, where $c_{ij}=2^{-N}{\rm tr}(P_{j}UP_{i}U^{\dagger})$, and $\sum_{j}|c_{ij}|^{2}=1$.
+Thus Eq.
+(6) is essentially the logarithm of $2^{-2N}\sum_{ij}|c_{ij}|^{2\alpha}$, where the additional factor of $2^{-2N}$ is due to the summation over all $P_{i}$.
+The definition clearly parallels the state SRE, but now for the unitary directly, which basically characterizes the similarity between $U$ and a Clifford unitary for which $c_{ij}=\pm 1$ for one specific $j$.
+We prove that this measure establishes a lower bound on the $T$-count, i.e. the minimal number of $T$ gates needed to synthesize a unitary $U$. More precisely,
+Theorem 1.
+For any $\alpha\geq 0$,
+$t(U)\geq v(U)\geq H_{\alpha}(U),$ (7)
+where the unitary stabilizer nullity is defined as $v(U):=2N-{\rm log}|s(U)|$, where $s(U):=\{P_{i}\in P_{N}|UP_{i}U^{\dagger}=\pm P_{j}\}$ [51].
+If the total amount of magic of the unitary is given by the unitary SRE above, what initial stabilizer state $|\psi\rangle$, if it exists, can be used to extract all the magic hidden in $U_{A}$?
+The answer is simply a state that is maximally entangled between $A$ and an ancillary system $B$ of the same size.
+Theorem 2.
+The unitary SRE of $U$ is equal to the state SRE of $U$’s Choi state $|U\rangle:=U_{A}\otimes I_{B}|\rm Bell\rangle_{AB}^{\otimes N}$:
+$H_{\alpha}(U)=M_{\alpha}(|U\rangle).$ (8)
+Notice that the right hand side of Eq.
+(8) is precisely the kind of setup we consider in this work, namely, a unitary acts on a stabilizer initial state consisting of $N$ Bell pairs shared between subsystem $A$ and its complement.
+Explicitly, the linear stabilizer entropy of acting $U$ on subsystem $A$ when the initial stabilizer state has $|A|=E\leq|B|$ is $Y^{\rm lin}=1-\exp(-H_{2}(U))$.
+The above theorem also provides an intuitive interpretation of the unitary SRE we introduce — if we unfold a unitary operator into its dual state using the channel-state duality, then the magic of a unitary operator is precisely equal to that of its dual (Choi) state.
+Detailed proofs of Theorem 1 and 2 are given in the SM [49].
+Figure 3: Numerical simulations of magic injection to initial states with non-stabilizer entanglement.
+(a)&(b): imperfect Bell states; (c)&(d): states with engineered entanglement spectra (see text for a precise description).
+In (a)&(c), we plot the amount of preexisting magic in the initial states, while in (b)&(d) we plot the amount of injected magic: $\Delta\overline{M_{2}}=\overline{M_{2}^{\rm final}}-\overline{M_{2}^{\rm init}}$.
+Non-stabilizer entanglement.- So far, we have focused on initial states with only stabilizer entanglement.
 However, realistic dynamics almost always contains magic in the initial state.
-Does non-stabilizerness of this kind impose any road blocks on the magic highway?
-Our findings strongly suggest that, to leading order, the answer is no.
-Again consider an entangled initial state on $AB$ where $U_{A}$ is applied to inject magic.
-Since any initial magic local to $A$ can be absorbed via the averaging of $U_{A}$, let us start with a state where non-stabilizerness only lives in the quantum correlation.
+Does entanglement still enhance magic injection when the initial state is non-stabilizer?
+Our results indicate that, at leading order, the answer is yes.
+Consider an entangled state on $AB$ with magic injected by $U_{A}$.
+Because any local non-stabilizerness in $A$ is washed out by averaging over $U_{A}$, it suffices to focus on states where the non-stabilizer content resides entirely in the correlations between $A$ and $B$.
 Such a state is known to have non-local magic, which is non-trivial as long as the entanglement spectrum is non-flat [37].
@@
 For example, initial states with non-local magic can be parametrized in the form $|\psi\rangle=|0\rangle^{\otimes f_{A}}\sum_{i=1}^{2^{k}}\lambda_{i}|i\rangle_{A}\otimes|i\rangle_{B}$, where $|i\rangle$ denotes a computational basis state of $k$ qubits.
-We choose four arbitrary entanglement spectra: (i) $\lambda_{i}\propto e^{-i/2^{k}}$; (ii) $\lambda_{i}\propto i$; (iii) $\lambda_{i}\propto i^{2}$; (iv) $\lambda_{i}\propto i^{3}$ and examine the magic spread numerically.
-Magic propagation is again displaying an $E$ dependence similar to having a stabilizer state or imperfect Bell states as the initial state [Fig.
-2.(f)].
-Nonetheless, more care is needed going beyond leading order.
+We choose four arbitrary entanglement spectra: (i) $\lambda_{i}\propto e^{-i/2^{k}}$; (ii) $\lambda_{i}\propto i$; (iii) $\lambda_{i}\propto i^{2}$; (iv) $\lambda_{i}\propto i^{3}$ and examine the magic growth numerically.
+However, more care is needed going beyond leading order.
 It is clear that there is non-stabilizerness dependence even in the simplest example with imperfect Bell states at the order of $O(\sqrt{k})$.
 Furthermore, because initial states can contain magic, the remaining magic capacity for further injection is reduced, i.e., the change of total magic per magic state injection is smaller compared to stabilizer entanglement.
-We numerically calculate both the magic already existing in the initial states $\overline{Y}_{\text{ini}}$ and the change after the action of $U_{A}$: $\overline{Y}_{\text{final}}/\overline{Y}_{\text{ini}}$.
+We numerically calculate both the preexisting magic in the initial states and the injected magic by $U_{A}$, as quantified by the increase in the second SRE: $\Delta\overline{M_{2}}=\overline{M_{2}^{\rm final}}-\overline{M_{2}^{\rm init}}$.
 As shown in Fig.
-2(d,e), $\overline{Y}_{\rm ini}$ itself in general decays with the increase of entanglement.
-Nonetheless, we find that the ratio $\overline{Y}_{\text{final}}/\overline{Y}_{\text{ini}}$ also decays exponentially with the amount of entanglement.
-For the same $E$, the change in global magic becomes smaller as the initial state becomes more magical.
-Discussion.- In this work, we demonstrate that stabilizer entanglement functions as a highway that facilitates the spread of locally injected magic throughout the entire system.
-In particular, we consider an analytically tractable setup where a local Haar random unitary acts on a subregion of a non-magical but potentially highly entangled state, thereby separating the process of magic injection from entanglement generation.
-We find that the average magic, as quantified by the linear stabilizer entropy, is enhanced by stabilizer entanglement.
+3(a,c), $\overline{Y^{\rm lin}}$ of both classes of initial states grow with entanglement, indicating that on average the initial state contains more magic when more entangled.
+The injected magic $\Delta{\overline{M_{2}}}$ also increases with the amount of entanglement [Fig.
+3(b,d)].
+This suggests that even non-stabilizer entanglement enhances the effect of magic injection, albeit at a much smaller rate.
+Indeed, as seen from Fig.
+3, the amount of injected magic $\Delta\overline{M_{2}}$ becomes smaller as the initial state becomes more magical at the same $E$.
+Discussion.- In this work, we demonstrate that entanglement enhances the effect of locally injected magic.
+We showed this in two ways — by locally injecting magic using a Haar random unitary, and by directly computing the upper limit of maximally extractable magic using unitary SRE.
+The former shows a precise increase of magic as a function of entanglement while the latter proves that the upper limit is attained when the initial state is maximally entangled.
 We further generalize our results to tripartite entanglement and non-stabilizer entanglement.
-A few interesting open questions remain.
-Our results for imperfect Bell-pair states suggest that the average amount of injected magic decreases as the amount of magic already existing in the target state increases.
-This implies that there may be a notion of intrinsic magic capacity of a given state that depends on both its entanglement and magic.
-Addressing this question would require a more systematic extension of our results to non-stabilizer states, which we leave for future work.
-Secondly, once we replace the Haar random $U_{A}$ with a more structured brickwork circuit, the question of what is the minimal circuit depth required for the effect of magic injection to coincide with that of a Haar random $U_{A}$ is worth exploring.
-The question is perhaps intimately related to the complexity growth of local random unitary circuits with circuit depth.
-Acknowledgement.- We thank Cheng Wang, Yingfei Gu, and Yuzhen Zhang for helpful discussions, and Xhek Turkeshi for useful comment on the manuscript.
+As any stabilizer state is equivalent to the tensor product of Bell states, GHZ states and unentangled ancilla through local Clifford rotations, our results apply generally to any bipartite and tripartite stabilizer state.
+On the resource-theoretic front, we define a new computable measure $H_{\alpha}(U)$ for estimating $t(U)$, which leads to a natural interpretation of entanglement-enhanced magic injection as finding an initial stabilizer state that most efficiently extract the magic contained in a given unitary.
+Furthermore, it does not rely on optimization and can be computed far more efficiently via tensor network methods, unlike unitary or state nullity.
+This makes $H_{\alpha}$ a more computable unitary magic measure and a powerful estimate of unitary magic.
+Several open questions remain.
+Our analysis of imperfect Bell-pair states suggests that the average amount of injected magic decreases as the target state itself becomes more magical.
+This points to a possible notion of an intrinsic magic capacity for a quantum state, determined jointly by its entanglement and existing magic.
+A systematic extension of our results to generic non-stabilizer states would be needed to make this precise.
+Another direction concerns replacing the Haar-random $U_{A}$ with a structured brickwork circuit: what minimal circuit depth is required for the injected magic to match that of a Haar-random unitary?
+This question is likely connected to the growth of complexity in local random unitary circuits as a function of depth.
+Acknowledgement.- We thank Zi-Wen Liu, Jong-Yeon Lee, Cheng Wang, Yingfei Gu, and Yuzhen Zhang for helpful discussions, and Xhek Turkeshi for useful comments on the manuscript.
 We especially thank Daniele Iannotti, Gianluca Esposito, Lorenzo Campos Venuti, and Alioscia Hamma for sharing their note on obtaining the exact expression of average stabilizer purity at a fixed amount of entanglement.
-This work is supported by Grant No.
-12375027 from the National Natural Science Foundation of China (Z.-C.
-Y.).
+Z.-C.Y.
+is supported by Grant No.
+12375027 from the National Natural Science Foundation of China.
 Numerical simulations were performed on the High-performance Computing Platform of Peking University.
-Supplemental Material for “Stabilizer entanglement as a magic highway”
-Appendix A Analytical method
+Supplemental Material for “Stabilizer Entanglement Enhances Magic Injection”
+Appendix A Additional numerical results on the second stabilizer Rényi entropy $\overline{M_{2}}$
+In this section, we provide additional numerical results on the Haar-average value of the second stabilizer Rényi entropy $\overline{M_{2}}$, for the setup depicted in Fig.
+1(a) of the main text.
+We find that the the average $\overline{M_{2}}$ is extremely close to ${\rm log}(1-\overline{Y^{\rm lin}})$, see Fig.
+4.
+This indicates that the average $\overline{Y^{\rm lin}}$ is also typical, as sample-to-sample fluctuations are small.
+Figure 4: Numerical simulations of $\overline{Y^{\rm lin}}$ and $\overline{M_{2}}$ for the setup depicted in Fig.1(a) of the main text.
+$1-\overline{Y^{\rm lin}}$ is shown as upper triangles and $2^{-\overline{M_{2}}}$ is shown as lower triangles.
+Appendix B Tripartite entanglement
+In this section, we extend the main results presented in the main text to tripartite entanglement.
+Consider a stabilizer state shared among three parts $A$, $B$ and $C$.
+Since any tripartite stabilizer state can be transformed via local unitaries acting on $A$, $B$, and $C$ into a collection of GHZ states, Bell states, and single-qubit states [50], we can, without loss of generality, consider initial states as depicted in Fig.
+5(a).
+Specifically, the initial state consists of $g$ GHZ states shared among $A$, $B$, and $C$; $b_{AB}$, $b_{AC}$, and $b_{BC}$ Bell states shared between subsystems $AB$, ACAC, and $BC$, respectively, and $f_{A}$, $f_{B}$, $f_{C}$ single-qubit states in each individual subsystem.
+Since magic injection via unitary operations on any single subsystem in this setting mirrors the bipartite case discussed earlier, we first consider unitaries of the form $U=U_{A}\otimes U_{B}$ [Fig.
+5(b)].
+In the limit where $|A|\gg 1$ and $|B|\gg 1$, we find that the average $\overline{Y^{\rm lin}}$ after applying $U$ is given by
+$\overline{Y^{\rm lin}}$ $=$ $\mathbb{E}_{U_{A}}\mathbb{E}_{U_{B}}Y^{\rm lin}\left(U_{A}U_{B}|\psi\rangle\langle\psi|U_{A}^{\dagger}U_{B}^{\dagger}\right)$ (9)
+$=$ $1-4\cdot{2}^{-|A|-|B|-g-{{b}_{AC}}-{{b}_{BC}}}\left[1+3\cdot{2}^{-2{{b}_{AB}}-g}\right],$
+where subleading corrections are of order $O(2^{-|A|},2^{-|B|})$.
+A detailed derivation of Eq.
+(9) is given in Sec.
+C.3.
+Eq.
+(9) can be understood as follows.
+The average value $1-\overline{Y^{\rm lin}}$ is once again exponentially suppressed in both the total size of the subregion where the unitary is applied: $|A|+|B|$, and the amount of entanglement between subregion $AB$ and its complement: $g+b_{AC}+b_{BC}$.
+Hence the physical interpretation of the prefactor in Eq.
+(9) is identical to that of Eq.
+(2) in the main text.
+The discrepancy between these two cases lies in the additional factor $3\cdot{2}^{-2{{b}_{AB}}-g}$, which becomes negligible once $A$ and $B$ are sufficiently entangled.
+This is precisely what we have seen in Eq.
+(LABEL:eq:UaUb) of the main text: the effect of a factorized unitary $U_{A}\otimes U_{B}$ becomes essentially indistinguishable from that of a global unitary from the standpoint of magic, once $A$ and $B$ are sufficiently entangled.
+Thus, Eq.
+(9) encapsulates the combined effects of Eqs.
+(2) and  (LABEL:eq:UaUb) of the main text in the tripartite setting.
+In Sec.
+E, we further consider a factorized unitary acting on all three subregions: $U=U_{A}\otimes U_{B}\otimes U_{C}$, where we confirm that in this case the effect of a factorized Haar random unitary on magic injection again becomes indistinguishable from that of a global Haar random unitary, provided that the subregions are sufficiently entangled with one another.
+Figure 5: (a) Schematic of the initial stabilizer state used for tripartite systems.
+Notice that any tripartite stabilizer state can be brought into this form via local Clifford unitaries acting on subsystems $A$, $B$ and $C$.
+(b) Schematic of the setup for $U_{A}\otimes U_{B}$ acting on tripartite systems.
+(c) Numerical results for $U_{A}\otimes U_{B}$ acting on tripartite systems.
+The dashed lines represent the quantity shown in Eq.
+(9) and the squares represent the exact result Eq.
+(114).
+Appendix C Analytical method
 In this section, we present the analytical method used to derive the main results quoted in the main text.
-In Section A.1, we provide a brief introduction to a diagrammatic approach for computing polynomials of matrix elements of unitary matrices averaged over the Haar ensemble.
-Section A.2 introduces a key quantity used throughout our analysis and outlines approximations made to simplify both the calculation and the final results.
-Finally, in Section A.3, we detail the step-by-step derivation of the results presented in the main text.
+In Section C.1, we provide a brief introduction to a diagrammatic approach for computing polynomials of matrix elements of unitary matrices averaged over the Haar ensemble.
+Section C.2 introduces a key quantity used throughout our analysis and outlines approximations made to simplify both the calculation and the final results.
+Finally, in Section C.3, we detail the step-by-step derivation of the results presented in the main text.
 Exact results of all setups considered in the main text (i.e.
-without any approximation such as $|A|\gg 1$) can also be obtained using an alternative analytical method, as provided in Appendix E.
+without any approximation such as $|A|\gg 1$) can also be obtained using an alternative analytical method, as provided in Appendix H.
 This alternative approach, while more compatible with symbolic computations, offers less physical intuition.
-A.1 A diagrammatic approach
+C.1 A diagrammatic approach
 In this subsection, we outline a diagrammatic method for systematically computing expectation values of polynomials in the matrix elements of unitary matrices under the Haar ensemble.
@@
 The diagrams consist of the building blocks shown in Fig.
-3.
+6.
 The matrix elements $U_{ab}$ or $U^{*}_{\alpha\beta}$ are represented by thick dotted lines.
@@
 As an example, the functions $f(U)=\text{Tr}(AUBU^{\dagger})$ and $g(U)=\text{Tr}(AUBUCU^{\dagger}DU^{\dagger})$ are represented in Fig.
-4.
+7.
 Figure 6: Diagrammatic representations for the unitary matrix $U$ and $U^{*}$, the fixed matrix $A$ and the Kronecker delta.
@@
 This pairing is represented by adding additional thin lines connecting the endpoints of $U$ and that of $U^{*}$ in the diagram, with black dots connected to black dots, and white dots connected to white dots, as depicted in Fig.
-4.
+7.
 To associate each diagram with a specific value, we follow the following rules:
@@
 By connecting the dots with thin lines, we obtain the diagrams shown in Fig.
-4.
+7.
 For $f$, there is only one diagram, which consists of a single $U$-cycle of length 1 (with weight $V_{1}$) and two $T$-cycles, which generate $\text{Tr}A$ and $\text{Tr}B$.
@@
 $\overline{g(U)}=V_{1,1}\left[\text{Tr}(ABCD)+\text{Tr}A\ \text{Tr}(BD)\ \text{Tr}C\right]+V_{2}\left[\text{Tr}A\ \text{Tr}(BDC)+\text{Tr}(ADB)\ \text{Tr}C\right].$ (13)
-A.2 A useful quantity and its approximations
-In this subsection, we define a quantity that will be used intensively in evaluating $\overline{Y}$.
+C.2 A useful quantity and its approximations
+In this subsection, we define a quantity that will be used intensively in evaluating $\overline{Y^{\rm lin}}$.
 We will calculate the value of this quantity in its most general form, to leading order in $2^{|A|+E}$, where $E$ is the entanglement entropy.
@@
 In the above definition, we also abuse notation by using $\mathbb{S}_{A}$ to represent an equal weight summation of all elements in $\mathbb{S}_{A}$: $\sum_{g\in\mathbb{S}_{A}}g$.
-The connection between the above quantity and $\overline{Y}$ will be explained in detail in the next subsection.
+The connection between the above quantity and $\overline{Y^{\rm lin}}$ will be explained in detail in the next subsection.
 Basically it naturally arises when computing ${\rm Tr}(PU_{A}\rho U_{A}^{\dagger})^{4}$, with $\rho$ represented as a projector onto the stabilizer subspace.
-Eq. (10) can be represented using the diagrammatic method introduced in Sec. A.1 as follows:
+Eq. (14) can be represented using the diagrammatic method introduced in Sec. C.1 as follows:
 Figure 8: (a) Leading order and the corresponding diagrams for $\alpha_{P_{A}}(u_{1},u_{2},u_{3},u_{4})$.
-There are $576{4!}^{2}=$ 576 diagrams in total.
+There are ${4!}^{2}=576$ diagrams in total.
 If $P_{A}\neq I_{A}$, then $\text{Tr}(P_{A})=\text{Tr}(P_{A}^{3})=0$, and there are 216 diagrams with non-zero contributions.
@@
 $V_{1,1,1,1}=\frac{N^{4}-8N^{2}+6}{N^{2}(N^{2}-1)(N^{2}-4)(N^{2}-9)}=2^{-4|A|}+6\cdot 2^{-6|A|}+O(2^{-8|A|}),$ (15)
-$V_{2,1,1}=\frac{-N^{3}+4N}{N^{2}(N^{2}-1)(N^{2}-4)(N^{2}-9)}=-2^{-5|A|}+18\cdot 2^{-7|A|}+O(2^{-9|A|}),$
-$V_{2,2}=\frac{N^{2}+6}{N^{2}(N^{2}-1)(N^{2}-4)(N^{2}-9)}=2^{-6|A|}+20\cdot 2^{-8|A|}+O(2^{-10|A|}),$
+[…]
+[…]
 $V_{3,1}=\frac{2N^{2}-6}{N^{2}(N^{2}-1)(N^{2}-4)(N^{2}-9)}=2\cdot 2^{-6|A|}+8\cdot 2^{-8|A|}+O(2^{-10|A|}),$
@@
 In the above equation, the first line comes from terms like $V_{1,1,1,1}{\rm Tr}^{2}(P_{A}^{2}){\rm Tr}(u_{1}u_{2}\mathbb{S}_{A}^{2}){\rm Tr}(u_{3}u_{4}\mathbb{S}_{A}^{2})$, and the 2nd line comes from $V_{2,1,1}{\rm Tr}^{2}(P_{A}^{2}){\rm Tr}(u_{1}u_{2}\mathbb{S}_{A}^{2}){\rm Tr}(u_{3}\mathbb{S}_{A}){\rm Tr}(u_{4}\mathbb{S}_{A})$.
-A.3 Derivation of the key results presented in the main text
-In this subsection, we provide a detailed derivation of Eqs. (1), (4) and (5) in the main text.
-A.3.1 Eq. (1)
+C.3 Derivation of the key results presented in the main text
+In this subsection, we provide a detailed derivation of Eqs.
+(2), (LABEL:eq:UaUb) in the main text, as well as Eq.
+(9) in Sec.
+B.
+Since what we essentially need to compute is $1-\overline{Y^{\rm lin}}$, we will focus on this part below.
+C.3.1 Eq. (2)
 We start with Eq.
 (2).
-The key to our derivation is to represent the initial state as a projector onto the stabilizer subspace, and a decomposition of the stabilizer group of the full sytem in terms of the cosets of stabilizer subgroups of subsystems $A$ and $B$.
+The key to our derivation is to represent the initial state as a projector onto the stabilizer subspace, and a decomposition of the stabilizer group of the full system in terms of the cosets of stabilizer subgroups of subsystems $A$ and $B$.
 Let $\mathbb{S}_{AB}$ denote the stabilizer group of the initial state.
@@
 where $a_{i}\notin\mathbb{S}_{A}$ and $b_{i}\notin\mathbb{S}_{B}$ are the logical operators of subsystems $A$ and $B$, respectively.
-To ensure the proper commutation relation, the logical operators must be paired up into $(a_{i},b_{i})$ such that: (i) if $[a_{i},a_{j}]=0$, then $[b_{i},b_{j}]=0$; or (ii) if $\{a_{i},a_{j}\}=0$, then $\{b_{i},b_{j}\}=0$.
+To ensure the proper commutation relation, the logical operators must be paired up into $(a_{i},b_{i})$ such that: (i) if $a_{i},a_{j}]=0$, then $b_{i},b_{j}]=0$; or (ii) if $\{a_{i},a_{j}\}=0$, then $\{b_{i},b_{j}\}=0$.
 While it is not hard to convince oneself that the above decomposition is correct for pure states with ${\rm dim}(\mathbb{S}_{AB})=N$, let us check that the dimensions work out.
-First of all, since the entanglement between $A$ and $B$ is $E$, there are precisely $E$ logical qubits encoded in both subsystems, with $4^{E}$ logical operators in each subsytem, and hence $4^{E}$ terms in the decomposition.
+First of all, since the entanglement between $A$ and $B$ is $E$, there are precisely $E$ logical qubits encoded in both subsystems, with $4^{E}$ logical operators in each subsystem, and hence $4^{E}$ terms in the decomposition.
 Secondly, we can count the total number of group elements in $\mathbb{S}_{AB}$ from the right-hand side of the decomposition.
 We have $4^{E}\times 2^{{\rm dim}(\mathbb{S}_{A})}\times 2^{{\rm dim}(\mathbb{S}_{B})}=2^{|A|+|B|}=2^{N}$, as it should.
-We will give a concrete example of such a decomposition in Appendix B.
+We will give a concrete example of such a decomposition in Appendix D.
 We can now express the initial state using the above coset decomposition of $\mathbb{S}_{AB}$:
 ${{\rho}_{0}}=\frac{1}{{2}^{N}}{\mathbb{S}_{AB}}=\frac{1}{2^{N}}\left(\underbrace{\mathbb{S}_{A}\otimes\mathbb{S}_{B}\ +\ a_{1}\mathbb{S}_{A}\otimes b_{1}\mathbb{S}_{B}\ +\ a_{2}\mathbb{S}_{A}\otimes b_{2}\mathbb{S}_{B}\ +\ldots}_{4^{E}\ \rm{terms}}\right)$ (20)
-Armed with the above representation of $\rho_{0}$, we can cast the average value $\overline{Y}$ as:
-$\overline{Y}=\frac{1}{{{2}^{5N}}}\sum\limits_{P\in P_{N}}{\overline{{{\text{Tr}}^{4}}(P{{U}_{A}}{\mathbb{S}_{AB}}{{U}_{A}}^{\dagger})}}$ (17)
+Armed with the above representation of $\rho_{0}$, we can cast the average value $\overline{Y^{\rm lin}}$ as:
+$1-\overline{Y^{\rm lin}}=\frac{1}{{{2}^{5N}}}\sum\limits_{P\in P_{N}}{\overline{{{\text{Tr}}^{4}}(P{{U}_{A}}{\mathbb{S}_{AB}}{{U}_{A}}^{\dagger})}}$ (21)
 $=\frac{1}{{{2}^{5N}}}\sum_{P\in P_{N}}{\sum\limits_{k=0}^{{{2}^{2E}-1}}{{{\alpha}_{{{P}_{A}}}}({{a}_{k}},{{a}_{k}},{{a}_{k}},{{a}_{k}})\ {{\text{Tr}}^{4}}({{P}_{B}}{{b}_{k}}{{\mathbb{S}}_{B}})}}.$
@@
 (16) and (17).
-Consider first a fixed Pauli string $P$ and $k$, which corresponds to one term of the summation in $\overline{Y}$. We have the following result:
+Consider first a fixed Pauli string $P$ and $k$, which corresponds to one term of the summation in $\overline{Y^{\rm lin}}$. We have the following result:
 ${{{\alpha}_{{{P}_{A}}}}({{a}_{k}},{{a}_{k}},{{a}_{k}},{{a}_{k}})\ \text{Tr}^{4}({{P}_{B}}{{b}_{k}}{{\mathbb{S}}_{B}})}$ (22)
-$=$ $\delta(P_{B},b_{k}\mathbb{S}_{B})\times\begin{cases}3\cdot{{2}^{2{|A|}+4{|B|}-2E}}\left[1-2\cdot{{2}^{E-{|A|}}}+O({{2}^{-E-{|A|}}})\right],&\quad{{P}_{A}}\neq{{I}_{A}},\quad k=0\\ 3\cdot{{2}^{2{|A|}+4{|B|}-2E}}\left[1+O({{2}^{-E-{|A|}}})\right],&\quad{{P}_{A}}\neq{{I}_{A}},\quad k\neq 0\\ {{2}^{4{|A|}+4{|B|}}},&\quad{{P}_{A}}={{I}_{A}},\quad k=0\\ 0,&\quad{{P}_{A}}={{I}_{A}},\quad k\neq 0\\ \end{cases}$
+[…]
 In the above equation, $k=0$ corresponds to the first term in the coset decomposition, with $a_{0}=b_{0}=I$; $\delta(P_{B},b_{k}\mathbb{S}_{B})=1$ iff $P_{B}\in b_{k}\mathbb{S}_{B}$, and zero otherwise.
@@
 (2) quoted in the main text after dividing by $2^{5N}$.
-A.3.2 Eq. (4)
+The qualitative features of Eq.
+(2), particularly the factor $2^{-|A|-E}$, can also be understood—since the unitary acts only on subsystem $A$, stabilizers supported entirely on $B$ remain unchanged under $U_{A}$ and contribute to $1-Y^{\rm lin}(\rho)$.
+A simple counting shows that the number of surviving stabilizers is precisely $2^{|B|-E}$, leading to the factor $2^{-|A|-E}$ in Eq.
+(2).
+The prefactor 4 in front of $2^{-|A|-E}$ is, in some sense, non-universal.
+To see this, let us define the generalized linear stabilizer entropy $Y_{\alpha}^{\rm lin}(\rho):=1-2^{-N}\sum_{P\in P_{N}}{\rm tr}(P\rho)^{2\alpha}$.
+Eq.
+(2) thus corresponds to the specific case $\alpha=2$.
+Similarly, $1-Y_{\alpha}^{\rm lin}(\rho)$ is related to the $\alpha$-th stabilizer Renyi entropy: $M_{\alpha}(\rho)=\frac{1}{1-\alpha}{\rm log}(1-Y_{\alpha}^{\rm lin}(\rho))$.
+In the limit $\alpha\rightarrow\infty$, only the surviving stabilizers supported on $B$ contribute, yielding $\overline{Y_{\infty}^{\rm lin}}=1-2^{-|A|-E}$.
+For arbitrary $\alpha$, we have $1-\overline{Y_{\alpha}^{\rm lin}}=[1+c_{\alpha}(|A|,E)]2^{-|A|-E}$, with the one coming from the surviving stabilizers, and $c_{\alpha}(|A|,E)$ coming from contributions of all other Pauli strings, which in general depends on both $|A|$ and $E$  [^1] We thank Xhek Turkeshi for pointing this out to us..
+Specifically, Eq.
+(2) implies that for $\alpha=2$, we have $c_{2}=3$.
+C.3.2 Eq. (LABEL:eq:UaUb)
 We provide detailed calculations for Eq.
-(4) in the main text.
-Since the unitary acting on $\rho_{0}$ takes a factorized form $U=U_{A}\otimes U_{B}$, the average $\overline{Y}$ similarly factorizes into
-$\overline{Y}=\frac{1}{{{2}^{5N}}}\sum\limits_{P\in P_{N}}{\sum\limits_{{{k}_{1}},{{k}_{2}},{{k}_{3}},{{k}_{4}}}{{{\alpha}_{{{P}_{A}}}}({{a}_{{{k}_{1}}}},{{a}_{{{k}_{2}}}},{{a}_{{{k}_{3}}}},{{a}_{{{k}_{4}}}}){{\beta}_{{{P}_{B}}}}({{b}_{{{k}_{1}}}},{{b}_{{{k}_{2}}}},{{b}_{{{k}_{3}}}},{{b}_{{{k}_{4}}}})}},$ (20)
+(LABEL:eq:UaUb) in the main text.
+Since the unitary acting on $\rho_{0}$ takes a factorized form $U=U_{A}\otimes U_{B}$, the average $\overline{Y^{\rm lin}}$ similarly factorizes into
+$1-\overline{Y^{\rm lin}}=\frac{1}{{{2}^{5N}}}\sum\limits_{P\in P_{N}}{\sum\limits_{{{k}_{1}},{{k}_{2}},{{k}_{3}},{{k}_{4}}}{{{\alpha}_{{{P}_{A}}}}({{a}_{{{k}_{1}}}},{{a}_{{{k}_{2}}}},{{a}_{{{k}_{3}}}},{{a}_{{{k}_{4}}}}){{\beta}_{{{P}_{B}}}}({{b}_{{{k}_{1}}}},{{b}_{{{k}_{2}}}},{{b}_{{{k}_{3}}}},{{b}_{{{k}_{4}}}})}},$ (24)
 where $\beta_{P_{B}}\left(b_{k_{1}},b_{k_{2}},b_{k_{3}},b_{k_{4}}\right)$ is defined analogously to Eq. (14).
@@
 We now analyze the order of magnitude of each term.
-Since $\sum\limits_{{{k}_{1}},{{k}_{2}},{{k}_{3}},{{k}_{4}}}f_{i}\cdot(...)\leq\sum\limits_{{{k}_{1}},{{k}_{2}},{{k}_{3}},{{k}_{4}}}f_{i}$, where ()(...)( … ) represents any product of $\delta_{1},\delta_{2}$, or $\delta_{4}$ (as additional $\delta$-functions further reduce the number of unconstrained summations over $k_{1},\ldots,k_{4}$), we have
+Since $\sum\limits_{{{k}_{1}},{{k}_{2}},{{k}_{3}},{{k}_{4}}}f_{i}\cdot(...)\leq\sum\limits_{{{k}_{1}},{{k}_{2}},{{k}_{3}},{{k}_{4}}}f_{i}$, where ()(...) represents any product of $\delta_{1},\delta_{2}$, or $\delta_{4}$ (as additional $\delta$-functions further reduce the number of unconstrained summations over $k_{1},\ldots,k_{4}$), we have
 $\sum_{k_{1},k_{2},k_{3},k_{4}}f_{i}^{2}=\sum_{k_{1},k_{2},k_{3},k_{4}}f_{i}+O(2^{-E}),\quad\sum_{k_{1},k_{2},k_{3},k_{4}}f_{i}f_{j}\sim O(2^{-E})\times{\rm min}(\sum f_{i},\sum f_{j}).$ (26)
@@
 (iii) $P_{A}=I_{A}$ and $P_{B}=I_{B}$. In this case the result is trivially $2^{4|A|+4|B|}$.
-After combining all these contributions and including the multiplicity factors coming from the number of Pauli strings in each case, we obtain the final expression in Eq. (4) of the main text.
-A.3.3 Eq. (5)
+After combining all these contributions and including the multiplicity factors coming from the number of Pauli strings in each case, we obtain the final expression in Eq.
+(LABEL:eq:UaUb) of the main text.
+C.3.3 Eq. (9)
 We now present the calculation details of Eq.
-(5) in the main text.
+(9) in Sec.
+B.
 Let us first recall the setup we use for this case.
 We consider a tripartite system $A$, $B$, and $C$.
-Since any tripartite stabilizer state can be transformed into a set of single-qubit states, Bell-pair states, and GHZ states via local Clifford unitaries, we consider initial state consisting of $g$ GHZ states shared among $A$, $B$, and $C$; $b_{AB}$, $b_{AC}$, and $b_{BC}$ Bell pairs shared between subsystems $AB$, $AC$, and $BC$, respectively, and $f_{A}$, $f_{B}$, $f_{C}$ single-qubit states in each individual subsystem.
+Since any tripartite stabilizer state can be transformed into a set of single-qubit states, Bell states, and GHZ states via local Clifford unitaries, we consider initial state consisting of $g$ GHZ states shared among $A$, $B$, and $C$; $b_{AB}$, $b_{AC}$, and $b_{BC}$ Bell pairs shared between subsystems $AB$, ACAC, and $BC$, respectively, and $f_{A}$, $f_{B}$, $f_{C}$ single-qubit states in each individual subsystem, shown in Fig.
+5(a).
 Let us start by explicitly writing down a set of stabilizer generators for this state.
@@
 where $l_{AB}^{k}$ and $l_{C}^{k}$ are logical operators on $AB$ and $C$ as before.
-In this case, since we are dealing with initial state with a factorized structure, it is easy to directly idenfity the logical operators.
+In this case, since we are dealing with initial state with a factorized structure, it is easy to directly identify the logical operators.
 For example, $\{l_{AB}^{k}\}$ consists of Pauli operators acting on $b_{AC}+b_{BC}$ Bell pairs, as well as Pauli operators acting on $g$ GHZ states, hence $4^{b_{AC}+b_{BC}+g}$ operators in total.
-As before, to ensure elements across different cosets have the correct commutation relation as stabilizer group elements, we require (i) if $[l_{AB}^{i},l_{AB}^{j}]=0$, then $[l_{C}^{i},l_{C}^{j}]=0$; or (ii) if $\{l_{AB}^{i},l_{AB}^{j}\}=0$, then $\{l_{C}^{i},l_{C}^{j}\}=0$.
-The average $\overline{Y}$ can be written as
-$\overline{Y}=\frac{1}{{{2}^{5N}}}\sum\limits_{P\in P_{N}}{\overline{{{\text{Tr}}^{4}}(P{{U}_{AB}\otimes I_{C}}{\mathbb{S}_{ABC}}{U}_{AB}^{\dagger}\otimes I_{C})}}$ (30)
+As before, to ensure elements across different cosets have the correct commutation relation as stabilizer group elements, we require (i) if $l_{AB}^{i},l_{AB}^{j}]=0$, then $l_{C}^{i},l_{C}^{j}]=0$; or (ii) if $\{l_{AB}^{i},l_{AB}^{j}\}=0$, then $\{l_{C}^{i},l_{C}^{j}\}=0$.
+The average $\overline{Y^{\rm lin}}$ can be written as
+$1-\overline{Y^{\rm lin}}=\frac{1}{{{2}^{5N}}}\sum\limits_{P\in P_{N}}{\overline{{{\text{Tr}}^{4}}(P{{U}_{AB}\otimes I_{C}}{\mathbb{S}_{ABC}}{U}_{AB}^{\dagger}\otimes I_{C})}}$ (34)
 $=\frac{1}{{{2}^{5N}}}\sum_{P\in P_{N}}{\sum\limits_{k}{{\overline{{{\text{Tr}}^{4}}(P_{AB}{{U}_{AB}}l_{AB}^{k}{\mathbb{S}_{AB}}{U}_{AB}^{\dagger})}}\ {{\text{Tr}}^{4}}({{P}_{C}}{{l}_{C}^{k}}{{\mathbb{S}}_{C}})}},$
@@
     $t_{54}=0$ (60)
-The $\overline{Y}$ is then expressed in Eq. (57):
-$2^{5|A|+5|B|+5|C|}\cdot 2^{-4|C|}\cdot 2^{-|C|+b_{AC}+b_{BC}+g}\cdot\overline{Y}$ (57)
+The $\overline{Y^{\rm lin}}$ is then expressed in Eq. (61):
+$2^{5|A|+5|B|+5|C|}\cdot 2^{-4|C|}\cdot 2^{-|C|+b_{AC}+b_{BC}+g}\cdot(1-\overline{Y^{\rm lin}})$ (61)
 $=t_{11}+N_{2}t_{21}+N_{3}t_{31}+N_{4}t_{41}+N_{5}t_{51}$
@@
 Therefore, the subleading terms contributes $O(2^{-|B|-b_{BC}-b_{AB}-g})$ to the final result.
-Appendix B An example of the decomposition of the stabilizer group
+Appendix D An example of the decomposition of the stabilizer group
 In this section, we give a concrete example of the coset decomposition of a stabilizer group that has been widely used in this work. Consider a 5-qubit GHZ state:
@@
 Notably, these logical operators are combined in a manner that ensures all $a_{k}\otimes b_{k}$ commute with one another.
-$\quad\quad\mathbb{S}_{A}\otimes\mathbb{S}_{B}\quad\quad$ $XXX\mathbb{S}_{A}\otimes XX\mathbb{S}_{B}$ $IIZ\mathbb{S}_{A}\otimes ZI\mathbb{S}_{B}$ $-XXY\mathbb{S}_{A}\otimes YX\mathbb{S}_{B}$
+$\quad\quad\mathbb{S}_{A}\otimes\mathbb{S}_{B}\quad\quad$ $XXX\mathbb{S}_{A}\otimes XX\mathbb{S}_{B}$ $IIZ\mathbb{S}_{A}\otimes ZI\mathbb{S}_{B}$ $XXY\mathbb{S}_{A}\otimes YX\mathbb{S}_{B}$
 IIIII XXXXX IIZZI -XXYYX
@@
 Table 1: Stabilizer group of 5-qubit GHZ state and its decomposition.
-Appendix C Factorized unitary acting on three subregions
+Appendix E Factorized unitary acting on three subregions
 In this section, we consider the effect of a factorized unitary acting on all three subregions of a tripartite state: $U=U_{A}\otimes U_{B}\otimes U_{C}$. In this case, the final result reads:
-$\overline{Y}=4\cdot 2^{-N}\big{[}1+3\cdot{{2}^{-2{{b}_{AB}}-2{{b}_{AC}}-2{g}}}+3\cdot{{2}^{-2{{b}_{AB}}-2{{b}_{BC}}-2{g}}}+3\cdot{{2}^{-2{{b}_{AC}}-2{{b}_{BC}}-2{g}}}$ (62)
+$\overline{Y^{\rm lin}}=1-4\cdot 2^{-N}\big[1+3\cdot{{2}^{-2{{b}_{AB}}-2{{b}_{AC}}-2{g}}}+3\cdot{{2}^{-2{{b}_{AB}}-2{{b}_{BC}}-2{g}}}+3\cdot{{2}^{-2{{b}_{AC}}-2{{b}_{BC}}-2{g}}}$ (66)
 $+\frac{3}{2}\cdot{{2}^{-2{{b}_{AB}}-2{{b}_{AC}}-2{{b}_{BC}}-2{g}}}+\frac{9}{2}\cdot{{2}^{-2{{b}_{AB}}-2{{b}_{AC}}-2{{b}_{BC}}-3{g}}}\big]$
@@
 Similarly to the bipartite result Eq.
-(4) in the main text, we find that the resulting state also achieves maximal amount of magic, provided that the subsystems are sufficiently entangled.
-Interestingly, this does require that the $A$, $B$, and $C$ are pairwise entangled.
+(LABEL:eq:UaUb), we find that the resulting state also achieves maximal amount of magic, provided that the subsystems are sufficiently entangled.
+Interestingly, this doesn’t require that the $A$, $B$, and $C$ are pairwise entangled.
 As Eq.
-(62) suggests, if $A\&B$, $B\&C$ are highly entangled while $A\&C$ are not entangled (i.e.
-$b_{AB},b_{BC}$ is large, but $b_{AC}$ is zero), the average $\overline{Y}$ still approaches that of a Haar random state.
+(66) suggests, if A&BA\&B, B&CB\&C are highly entangled while A&CA\&C are not entangled (i.e.
+$b_{AB},b_{BC}$ is large, but $b_{AC}$ is zero), the average $\overline{Y^{\rm lin}}$ still approaches that of a Haar random state.
 As a consistency check, we consider a special case of Eq.
-(62), namely, when only $A\&B$ are entangled, and $C$ is not entangled with $AB$.
-In this case, we can check that the average $\overline{Y}$ coincides with a product of Eq.
-(4) in the main text for the bipartite case and a single Haar random state on $C$, which is consistent with our previous results.
+(66), namely, when only A&BA\&B are entangled, and $C$ is not entangled with $AB$.
+In this case, we can check that the average $1-\overline{Y^{\rm lin}}$ coincides with a product of Eq.
+(LABEL:eq:UaUb) in the main text for the bipartite case and a single Haar random state on $C$, which is consistent with our previous results.
 We now detail the calculation of Eq. (66). Our strategy parallels that used for the bipartite case. We first decompose the stabilizer group for the system $\mathbb{S}_{ABC}$ as
@@
 After applying $U:=U_{A}\otimes U_{B}\otimes U_{C}$ and averaging over Haar random unitaries, we obtain:
-$\overline{Y}=\frac{1}{{2^{5N}}}\sum\limits_{P\in P_{N}}\sum\limits_{k_{1},k_{2},k_{3},k_{4}}\alpha_{P_{A}}(a_{k_{1}},a_{k_{2}},a_{k_{3}},a_{k_{4}})\beta_{P_{B}}(b_{k_{1}},b_{k_{2}},b_{k_{3}},b_{k_{4}})\gamma_{P_{C}}(c_{k_{1}},c_{k_{2}},c_{k_{3}},c_{k_{4}}).$ (64)
+$1-\overline{Y^{\rm lin}}=\frac{1}{{2^{5N}}}\sum\limits_{P\in P_{N}}\sum\limits_{k_{1},k_{2},k_{3},k_{4}}\alpha_{P_{A}}(a_{k_{1}},a_{k_{2}},a_{k_{3}},a_{k_{4}})\beta_{P_{B}}(b_{k_{1}},b_{k_{2}},b_{k_{3}},b_{k_{4}})\gamma_{P_{C}}(c_{k_{1}},c_{k_{2}},c_{k_{3}},c_{k_{4}}).$ (68)
 Now, depending on the operator contents of the Pauli string $P$ on subsystems $A$, $B$ and $C$ (identity or not), we use Eq.
@@
 (17) to evaluate the above expression.
-We give one example here.
+We give an example here.
 Consider when $P_{A}\neq I_{A}$, $P_{B}\neq I_{B}$, and $P_{C}\neq I_{C}$, using Eq.
@@
 The subleading corrections are of the order $O(2^{-|A|},2^{-|B|},2^{-|C|})$ in each line. Cases not listed in Eq. (73) are similar. After combining all terms together, we obtain Eq. (66).
-Appendix D Magic injection with shallow-depth brickwork circuits
+Figure 9: The circuits used to simulate the case where the gate acting on A is not sampled from a global Haar ensemble.
+The blue lines represent Bell pairs.
+The blue bricks represent unrelated random Clifford gates, and the yellow bricks represent unrelated random magical gates.
+Figure 10: Numerical results for $\overline{Y^{\rm lin}}$ under brickwork circuits of varying depths and gate ranges, as depicted in Fig. 9.
+Appendix F Magic injection with shallow-depth brickwork circuits
 In this section, we replace the Haar random unitary gate acting on $A$ with a shallow-depth brickwork circuit.
@@
 1(d) of the main text, followed by one layer of local gates acting on subregion $A$.
-Since the initial state factorizes into single-qubit states and Bell-pair states, the average $\overline{Y}$ also factorizes into a product of contributions from single-qubit and Bell pair states.
-Thus, it is easy to see that, $\overline{Y}\propto 2^{-E}$ still holds true in this case.
+Since the initial state factorizes into single-qubit states and Bell-pair states, the average $1-\overline{Y^{\rm lin}}$ also factorizes into a product of contributions from single-qubit and Bell states.
+Thus, it is easy to see that, $1-\overline{Y^{\rm lin}}\propto 2^{-E}$ still holds true in this case.
 However, if $U_{A}$ is not Haar random, the specific choice of initial state in Fig.
@@
 We then consider brickwork circuits of varying depths and gate ranges, as summarized in Fig.
-6.
-We then numerically compute the average $\overline{Y}$ corresponding to each circuit architecture, and the results are shown in Fig.
-7.
-We find that the qualitative picture of our analytical result is still valid for the brickwork circuit, in that the average $\overline{Y}$ decays exponentially with $E$.
+9.
+We then numerically compute the average $\overline{Y^{\rm lin}}$ corresponding to each circuit architecture, and the results are shown in Fig.
+10.
+We find that the qualitative picture of our analytical result is still valid for the brickwork circuit, in that the average $1-\overline{Y^{\rm lin}}$ decays exponentially with $E$.
 However, notice that for extremely shallow depths (single layer and two layers), the curve bends upward and saturates, which indicates magic injection that is below capacity.
-As the circuit depth increases, the behavior of $\overline{Y}$ approaches that of a Haar random $U_{A}$.
-Figure 6: The circuits used to simulate the case where the gate acting on A is not sampled from a global Haar ensemble.
-The blue lines represent Bell pairs.
-The blue bricks represent unrelated random Clifford gates, and the yellow bricks represent unrelated random magical gates.
-Figure 7: Numerical results for $\overline{Y}$ under brickwork circuits of varying depths and gate ranges, as depicted in Fig. 6.
-Appendix E Exact results
-In this section, we present an alternative analytical method to derive exact expressions of $\overline{Y}$ in all cases discussed in the main text [^1] We acknowledge Alioscia Hamma for providing insightful notes, which served as the primary inspiration for the analytical method presented in this section..
-Recall the linear stabilizer entropy for a pure state $\rho=|\psi\rangle\langle\psi|$ is defined as $Y(\rho):=D^{-1}\sum_{P\in P_{N}}{\rm tr}(P\rho)^{4}$.
+As the circuit depth increases, the behavior of $\overline{Y^{\rm lin}}$ approaches that of a Haar random $U_{A}$.
+Appendix G Unitary stabilizer Rényi entropy
+Here we develop the formalism for unitary magic and T count based on ideas similar to the stabilizer Rényi entropy and unitary nullity.
+Definition G.1.
+Let $U$ be an $N$-qubit unitary, then the $\alpha$-th unitary stabilizer Rényi entropy of $U$ is defined as
+[…]
+where $P_{i}\in P_{N}$.
+For a shorthand notation, let $Q_{\alpha}:=\sum_{P_{i}\in P_{N}}{P_{i}}^{\otimes 2\alpha}$, then we can write it as:
+$H_{\alpha}(U)=\frac{1}{1-\alpha}\text{log}[{\text{tr}}(Q_{\alpha}U^{\otimes 2\alpha}Q_{\alpha}{U^{\dagger\otimes 2\alpha}})]-2N\frac{1+\alpha}{1-\alpha}.$ (75)
+As explained in the main text, the quantity $H_{\alpha}$ can be viewed as the $\alpha$-th Rényi entropy of the probability distribution associated with the weight $|c_{ij}|^{2}$ of $UP_{i}U^{\dagger}=\sum_{j}c_{ij}P_{j}$ over all $4^{N}$ Pauli strings.
+G.1 The relationship between unitary SRE and state SRE
+Consider a $2N$-qubit maximally entangled state between the original system and an auxiliary system, where a unitary $U$ acts on the $N$-qubit subsystem.
+The unitary stabilizer Rényi entropy $H_{\alpha}(U)$ is directly related to the state stabilizer Rényi entropy $M_{\alpha}$, defined as:
+$M_{\alpha}(\rho):=\frac{1}{1-\alpha}\text{log}[\text{tr}({\rho}^{\otimes 2\alpha}Q_{\alpha})]-N\frac{1}{1-\alpha}.$ (76)
+Theorem 3.
+Suppose $U$ is an $N$-qubit unitary operator.
+Let $|U\rangle_{AB}=U_{A}\otimes I_{B}|{{\rm Bell}\rangle}^{\otimes N}_{AB})$ be the Choi state of $U$ where $|A|=|B|=2^{N}$ and $|\mathrm{Bell}\rangle=(|00\rangle+|11\rangle)/\sqrt{2}$ is maximally entangled between $A$ and $B$ Then
+$M_{\alpha}(|U\rangle)=H_{\alpha}(U).$
+Proof.
+Consider a $2N$-qubit maximally entangled state between the original system $A$ and an auxiliary system $B$ of equal size: $\rho_{0}=|{\rm Bell}\rangle\langle{\rm Bell}|^{\otimes N}$.
+Recall that a $2$-qubit Bell state can be written as:
+$|{\rm Bell}\rangle\langle{\rm Bell}|=\frac{1}{4}(II+XX+ZZ-YY).$ (77)
+[…]
+$\rho_{0}=\frac{1}{2^{2N}}\sum_{P_{i}\in P_{N}}\underbrace{P_{i}}_{A}\underbrace{P_{i}}_{B}.$ (78)
+Now we consider the stabilizer Rényi entropy of state $|U\rangle=U_{A}\otimes I_{B}\ |{\rm Bell}\rangle_{AB}^{\otimes N}$:
+$M_{\alpha}(|U\rangle)=\frac{1}{1-\alpha}\text{log}[\text{tr}({{U_{A}}^{\otimes 2\alpha}\otimes{I_{B}}^{\otimes 2\alpha}{\rho_{0}}^{\otimes 2\alpha}{U_{A}}^{\dagger\otimes 2\alpha}\otimes{I_{B}}^{\otimes 2\alpha}}Q^{AB}_{\alpha})]-2N\frac{1}{1-\alpha},$ (79)
+the last term is not $N\frac{1}{1-\alpha}$ because $|U\rangle$ is a $2N$-qubit state. The key problem is to calculate the quantity inside the logarithm:
+$\text{tr}({{U_{A}}^{\otimes 2\alpha}\otimes{I_{B}}^{\otimes 2\alpha}{\rho_{0}}^{\otimes 2\alpha}{U_{A}}^{\dagger\otimes 2\alpha}\otimes{I_{B}}^{\otimes 2\alpha}}Q^{AB}_{\alpha})$ (80)
+$=$ $\sum_{P_{i_{1}},\dots,P_{i_{2\alpha}}\in P_{N}}2^{-4\alpha N}\text{tr}(Q^{AB}_{\alpha}{U_{A}}^{\otimes 2\alpha}\otimes{I_{B}}^{\otimes 2\alpha}\underbrace{P_{i_{1}}\cdots P_{i_{2\alpha}}}_{A}\otimes\underbrace{P_{i_{1}}\cdots P_{i_{2\alpha}}}_{B}{U_{A}}^{\dagger\otimes 2\alpha}\otimes{I_{B}}^{\otimes 2\alpha}).$
+From the definition of $Q_{\alpha}$, $Q^{AB}_{\alpha}$ can be written as a tensor product $Q^{AB}_{\alpha}=Q^{A}_{\alpha}\otimes Q^{B}_{\alpha}$, then Eq. (80) can be written as:
+$\text{tr}({{U_{A}}^{\otimes 2\alpha}\otimes{I_{B}}^{\otimes 2\alpha}{\rho_{0}}^{\otimes 2\alpha}{U_{A}}^{\dagger\otimes 2\alpha}\otimes{I_{B}}^{\otimes 2\alpha}}Q^{AB}_{\alpha})$ (81)
+$=$ $\sum_{P_{i_{1}},\dots,P_{i_{2\alpha}}\in P_{N}}2^{-4\alpha N}\text{tr}(Q_{\alpha}{U_{A}}^{\otimes 2\alpha}\underbrace{P_{i_{1}}\cdots P_{i_{2\alpha}}}_{A}{U_{A}}^{\dagger\otimes 2\alpha})\text{tr}(Q_{\alpha}{I_{B}}^{\otimes 2\alpha}\underbrace{P_{i_{1}}\cdots P_{i_{2\alpha}}}_{B}{I_{B}}^{\otimes 2\alpha}),$
+because $Q_{\alpha}:=\sum_{P_{i}\in P_{N}}{P_{i}}^{\otimes{2\alpha}}$, it will select those $P_{i_{1}}\cdots P_{i_{2\alpha}}$ with $P_{i_{1}}=P_{i_{2}}=\cdots=P_{i_{2\alpha}}$.
+Then Eq.
+(81) further simplifies:
+$\text{tr}({{U_{A}}^{\otimes 2\alpha}\otimes{I_{B}}^{\otimes 2\alpha}{\rho_{0}}^{\otimes 2\alpha}{U_{A}}^{\dagger\otimes 2\alpha}\otimes{I_{B}}^{\otimes 2\alpha}}Q^{AB}_{\alpha})$ (82)
+$=$ $\sum_{P_{i}\in P_{N}}2^{-4\alpha N}\text{tr}(Q_{\alpha}{U_{A}}^{\otimes 2\alpha}\underbrace{P_{i}\cdots P_{i}}_{2\alpha}{U_{A}}^{\dagger\otimes 2\alpha})\text{tr}(Q_{\alpha}{I_{B}}^{\otimes 2\alpha}\underbrace{P_{i}\cdots P_{i}}_{2\alpha}{I_{B}}^{\otimes 2\alpha})$
+$=$ $2^{-2\alpha N}\text{tr}(Q_{\alpha}{U_{A}}^{\otimes 2\alpha}Q_{\alpha}{U_{A}}^{\dagger\otimes 2\alpha}).$
+Then $M_{\alpha}(|U\rangle)$ can be written as:
+$M_{\alpha}(|U\rangle)=$ $\frac{1}{1-\alpha}\text{log}[\text{tr}({{U_{A}}^{\otimes 2\alpha}\otimes{I_{B}}^{\otimes 2\alpha}{\rho_{0}}^{\otimes 2\alpha}{U_{A}}^{\dagger\otimes 2\alpha}\otimes{I_{B}}^{\otimes 2\alpha}}Q^{AB}_{\alpha})]-2N\frac{1}{1-\alpha}$ (83)
+$=$ $\frac{1}{1-\alpha}\text{log}[2^{-2\alpha N}\text{tr}(Q_{\alpha}{U_{A}}^{\otimes 2\alpha}Q_{\alpha}{U_{A}}^{\dagger\otimes 2\alpha)}]-2N\frac{1}{1-\alpha}$
+$=$ $\frac{1}{1-\alpha}\text{log}[\text{tr}(Q_{\alpha}{U_{A}}^{\otimes 2\alpha}Q_{\alpha}{U_{A}}^{\dagger\otimes 2\alpha)}]-2N\frac{1+\alpha}{1-\alpha},$
+this is exactly the definition of $H_{\alpha}(U)$ in Eq. (75), so we get $H_{\alpha}(U)=M_{\alpha}(|U\rangle)$. ∎
+We have thus established that the unitary stabilizer Rényi entropy $H_{\alpha}(U)$ is equal to the state Rényi stabilizer entropy of a maximally entangled state between the original system and an auxiliary system of equal size, acted upon by $U$.
+Notice that this result does not involve averaging over a particular unitary ensemble, and holds for each and arbitrary $U$.
+G.2 Lower bound on the unitary nullity and $T$ count
+In this section, we establish the full inequality of Eq.
+(7) of the main text.
+Let $p_{ij}(U):={\text{tr}}^{2}(P_{i}UP_{j}U^{\dagger})/2^{4N}$.
+Because $\sum_{i,j}p_{ij}=1$, the coefficients $p_{ij}$ form a normalized probability distribution.
+Then the $\alpha$-th unitary stabilizer Rényi entropy we have defined can be written in the following way:
+$H_{\alpha}(U):=\frac{1}{1-\alpha}\text{log}\sum_{i,j}{p_{ij}(U)}^{\alpha}-2N.$ (84)
+To connect the unitary SRE to unitary nullity, we first note the following lemma.
+Theorem 4.
+Let $t(U)$ be the T-count of $U$, i.e., the number of T gates needed to synthesize $U$ in a Clifford+T decomposition and let $\nu(U)$ be the unitary nullity. Then $H_{0}(U)\leq\nu(U)\leq t(U)$.
+Proof.
+Define $s(U):=\{P_{i}|\ UP_{i}U^{\dagger}\in P_{N}\}$ and $g(U):=Us(U)U^{\dagger}$.
+Both $s(U)$ and $g(U)$ are subgroups of $P_{N}$.
+Recall the definition of unitary nullity [51] is $v(U):=2N-\text{log}|s(U)|$.
+We will show that, for a fixed $P_{i}$, the number of non-zero $p_{ij}$ is upper bounded by $\frac{2^{2N}}{|s(U)|}$.
+The group $g(U)$ acts on the operator space via conjugation, $R(X):=RXR,\ R\in g(U)$.
+For any $R\in g(U)$, there exists $Q\in s(U)$ with $R=UQU^{\dagger}$.
+Thus for any Pauli string $P_{i}$ and any $R\in g(U)$, we have:
+$R(UP_{i}U^{\dagger})=RUP_{i}U^{\dagger}R=U(QP_{i}Q)U^{\dagger},$ (85)
+for some $Q\in s(U)$. Since Pauli strings either commute or anticommute, that is, $QP_{i}=\pm P_{i}Q$, we must have
+$R(UP_{i}U^{\dagger})=\pm UP_{i}U^{\dagger}.$ (86)
+Thus, regarding $UP_{i}U^{\dagger}$ as a vector in operator space and $R(\cdot)$ as a linear transformation, we see that $UP_{i}U^{\dagger}$ is a common eigenvector of all the $R\in g(U)$, with eigenvalues $\pm 1$.
+Expanding $UP_{i}U^{\dagger}$ in Pauli basis,
+$UP_{i}U^{\dagger}=\sum_{S\in P_{N}}c_{S}S,$ (87)
+and applying $R(\cdot)$ gives
+$R(UP_{i}U^{\dagger})=\sum_{S}c_{S}(RSR)=\pm\sum_{S}c_{S}S.$ (88)
+Since $RSR=\pm S$, Eq.
+(86) imposes the constraint that all Pauli strings in the above expansion with $c_{S}\neq 0$ must share the same commutation relation with $R$.
+Modulo a global phase, the space of $P_{N}$ is isomorphic to the binary symplectic space $\mathbb{F}_{2}^{2N}$.
+Each Pauli string $S=X_{1}^{s_{1}}Z_{1}^{s_{2}}X_{2}^{s_{3}}Z_{2}^{s_{4}}...X_{N}^{s_{2N-1}}Z_{N}^{s_{2N}}$ corresponds to a vector $s=(s_{1},s_{2},...,s_{2N})$, and each $R\in g(U)$ can be equivalently represented as a vector $r\in\mathbb{F}_{2}^{2N}$.
+The commutation relation between $R$ and $S$ is encoded by the symplectic form:
+$\varepsilon(R,S)=(-1)^{\langle r,s\rangle},\quad\langle r,s\rangle:=\sum^{N}_{i=1}r_{2i-1}s_{2i}+r_{2i}s_{2i-1}\quad({\rm mod}\ 2)$ (89)
+where the $\langle r,s\rangle$ is the symplectic inner product of $r,s$.
+Therefore, the constraint that all nonzero $c_{S}$ share the same commutation relation with $R\in g(U)$ becomes a system of linear equations:
+$\langle r,s\rangle=b(r)\in\{0,1\}\quad\forall\ r\in G,$ (90)
+where $G\subseteq\mathbb{F}^{2N}_{2}$ is the vector space corresponding to $g(U)$, with $\text{dim}G=\text{log}|g(U)|=\text{log}|s(U)|$.
+Thus the solution space lies within the affine translation of the symplectic orthogonal complement $G^{\perp}$, whose size is $|G^{\perp}|=2^{2N-\text{dim}G}=\frac{2^{2N}}{|s(U)|}$.
+Hence, the number of nonzero coefficients $c_{S}$ is at most $\frac{2^{2N}}{|s(U)|}$.
+Summing over all Pauli strings $P_{i}$, we obtain:
+$H_{0}(U)=\text{log}(\#\{p_{ij}\neq 0\})-2N\leq\text{log}(\frac{2^{4N}}{|s(U)|})-2N=v(U).$ (91)
+Since $t(U)\geq v(U)$ [51], it follows that $t(U)\geq v(U)\geq H_{0}(U)$. ∎
+Corollary G.1.
+For any unitary $U$, its $T$-count is lower bounded such that $t(U)\geq H_{\alpha}(U)$.
+Proof.
+Recall that Rényi entropy is non-increasing in $\alpha$ such that $H_{\infty}(U)\leq H_{\alpha}(U)\leq H_{0}(U)$ where $H_{0}(U)=\text{log}(\#\{p_{ij}\neq 0\})-2N$. ∎
+Although $H_{\alpha}(U)$ is not as tight a lower bound compare to unitary nullity, the calculation of $H_{\alpha}(U)$ does not involve any optimization.
+Furthermore, if $U_{A}$ has a finer structure, such as a finite depth circuit, then the stabilizer Renyi entropy of its Choi state can be computed efficiently using tensor network methods.
+Hence we obtain a computable measure of the unitary magic and a computable lower bound of $t(U)$.
+Appendix H Exact results
+In this section, we present an alternative analytical method to derive exact expressions of $\overline{Y^{\rm lin}}$ in all cases discussed in the main text [^2] We acknowledge Alioscia Hamma for providing insightful notes, which served as the primary inspiration for the analytical method presented in this section..
+Recall the linear stabilizer entropy for a pure state $\rho=|\psi\rangle\langle\psi|$ is defined as $Y^{\rm lin}(\rho):=1-\frac{1}{D}\sum_{P\in P_{N}}{\rm tr}(P\rho)^{4}$.
 where $D:={\rm dim}\ \mathcal{H}$.
 This quantity admits an equivalent expression:
-$Y(\rho)=D{\rm tr}(Q\rho^{\otimes 4}),$ (70)
-where $Q:=D^{-2}\sum_{P\in P_{N}}P^{\otimes 4}$.
+$Y^{\rm lin}(\rho)=1-\frac{1}{D}{\rm tr}(Q\rho^{\otimes 4}),$ (92)
+where $Q:=\sum_{P_{i}\in P_{N}}{P_{i}}^{\otimes 4}$.
 For a Hilbert space $\mathcal{H}=\otimes^{n}_{i=1}\mathcal{H}_{i}$, the operator $Q$ factories as $Q=\otimes^{n}_{i=1}Q_{i}$.
-We can then interpret the map $M(\cdot):=D{\rm tr}(Q(\cdot))$ as a linear operator, which facilitates subsequent calculations.
-E.1 Eq. (1)
+We can then interpret the map ${\rm tr}(Q(\cdot))$ as a linear operator, which facilitates subsequent calculations.
+H.1 Eq. (2)
 In this subsection, we present a detailed derivation of the exact form of Eq. (2). The initial state is given by:
@@
 Substituting this into the expression for the quantity in Eq. (2), we have:
-$\overline{Y}=\mathbb{E}_{U_{A}}\ Y\left(U_{A}|\psi\rangle\langle\psi|\ U_{A}^{\dagger}\right)=D{\rm tr}\left[Q\mathbb{E}_{U_{A}}\left({U_{A}}^{\otimes 4}\otimes{I_{B}}^{\otimes 4}{|\psi\rangle\langle\psi|}^{\otimes 4}\ {U_{A}^{\dagger}}^{\otimes 4}\otimes{I_{B}}^{\otimes 4}\right)\right].$ (73)
+$\overline{Y^{\rm lin}}=\mathbb{E}_{U_{A}}\ Y^{\rm lin}\left(U_{A}|\psi\rangle\langle\psi|\ U_{A}^{\dagger}\right)=1-\frac{1}{D}{\rm tr}\left[Q\mathbb{E}_{U_{A}}\left({U_{A}}^{\otimes 4}\otimes{I_{B}}^{\otimes 4}{|\psi\rangle\langle\psi|}^{\otimes 4}\ {U_{A}^{\dagger}}^{\otimes 4}\otimes{I_{B}}^{\otimes 4}\right)\right].$ (95)
 Notice the Haar average of an operator $O\in\mathcal{B}(\mathcal{H}^{\otimes k})$ reads:
@@
 with $V_{\sigma,\pi}$ being the Weingarten function and $T_{\sigma},T_{\pi}$ being the aforementioned permutation operators. Applying this to Eq. (95), we have:
-$\overline{Y}=\sum_{\sigma,\pi}\sum_{\vec{i},\vec{i^{\prime}}}2^{-4E}DV^{A}_{\sigma,\pi}{\rm tr}\left(Q_{A}T^{A}_{\pi}\right){\rm tr}\left(T^{A-f_{A}}_{\sigma}{|\vec{i}\rangle\langle{\vec{i^{\prime}}}|}_{A}\right){\rm tr}\left(Q_{B}{|\vec{i}\rangle\langle\vec{i^{\prime}}|}_{B}\right).$ (75)
+$1-\overline{Y^{\rm lin}}=\sum_{\sigma,\pi}\sum_{\vec{i},\vec{i^{\prime}}}\frac{1}{D}2^{-4E}V^{A}_{\sigma,\pi}{\rm tr}\left(Q_{A}T^{A}_{\pi}\right){\rm tr}\left(T^{A-f_{A}}_{\sigma}{|\vec{i}\rangle\langle{\vec{i^{\prime}}}|}_{A}\right){\rm tr}\left(Q_{B}{|\vec{i}\rangle\langle\vec{i^{\prime}}|}_{B}\right).$ (97)
 We have used the fact that $Q=Q_{A}\otimes Q_{B}$ and $T^{A}_{\sigma}=T^{f_{A}}_{\sigma}\otimes T^{A-f_{A}}_{\sigma}$. Furthermore, because of the tensor product structure, we have:
-$F_{1}(\sigma,E)$ ::=:= $\sum_{\vec{i},\vec{i^{\prime}}\in{\{0,1\}}^{\otimes 4E}}{\rm tr}\left(T^{A-f_{A}}_{\sigma}{|\vec{i}\rangle\langle{\vec{i^{\prime}}}|}_{A}\right){\rm tr}\left(Q_{B}{|\vec{i}\rangle\langle\vec{i^{\prime}}|}_{B}\right)$
+$F_{1}(\sigma,E)$ $:=$ $\sum_{\vec{i},\vec{i^{\prime}}\in{\{0,1\}}^{\otimes 4E}}{\rm tr}\left(T^{A-f_{A}}_{\sigma}{|\vec{i}\rangle\langle{\vec{i^{\prime}}}|}_{A}\right){\rm tr}\left(Q_{B}{|\vec{i}\rangle\langle\vec{i^{\prime}}|}_{B}\right)$
 $=$ $\sum_{\vec{i}\in{\{0,1\}}^{\otimes 4E}}{\rm tr}\left(Q_{B}{|\vec{i}\rangle\langle\sigma(\vec{i})|}_{B}\right)=\left[\sum_{\vec{j}\in{\{0,1\}}^{\otimes 4}}{\rm tr}\left(Q_{1}{|\vec{j}\rangle\langle\sigma(\vec{j})|}_{1}\right)\right]^{E},$
 where $\sigma(\vec{i}):=(i_{\sigma(1)},i_{\sigma(2)},i_{\sigma(3)},i_{\sigma(4)})$. Then the Eq. (97) can be written as the formula shown in Eq. (99).
-$\overline{Y}=\sum_{\sigma,\pi}2^{-4E}DV^{A}_{\sigma,\pi}{\rm tr}\left(Q_{A}T^{A}_{\pi}\right)F_{1}(\sigma,E),$ (77)
+$1-\overline{Y^{\rm lin}}=\sum_{\sigma,\pi}2^{-4E}DV^{A}_{\sigma,\pi}{\rm tr}\left(Q_{A}T^{A}_{\pi}\right)F_{1}(\sigma,E),$ (99)
 where all the terms represent rational fractions, and the entire summation can be completed using symbolic computation.
@@
 (100).
-$\overline{Y}=\frac{4(D_{A}^{2}D_{E}^{2}-3D_{A}D_{E}-6D_{E}^{2}+6)}{D_{A}(D_{A}^{2}-9)D_{E}^{3}}.$ (78)
-E.2 Eq. (4)
+$\overline{Y^{\rm lin}}=1-\frac{4(D_{A}^{2}D_{E}^{2}-3D_{A}D_{E}-6D_{E}^{2}+6)}{D_{A}(D_{A}^{2}-9)D_{E}^{3}}.$ (100)
+H.2 Eq. (LABEL:eq:UaUb)
 In this subsection, we present the analytical method to derive the exact expression of Eq. (LABEL:eq:UaUb). The initial state is shown in Eq. (101).
@@
 The quantity in Eq. (LABEL:eq:UaUb) then can be written as:
-$\overline{Y}=\mathbb{E}_{U_{A}}\mathbb{E}_{U_{B}}\ Y\left(U_{A}U_{B}|\psi\rangle\langle\psi|\ U_{A}^{\dagger}U_{B}^{\dagger}\right)=D{\rm tr}\left[Q\mathbb{E}_{U_{A}}\mathbb{E}_{U_{B}}\left({U_{A}}^{\otimes 4}\otimes{U_{B}}^{\otimes 4}{|\psi\rangle\langle\psi|}^{\otimes 4}\ {U_{A}^{\dagger}}^{\otimes 4}\otimes{U_{B}^{\dagger}}^{\otimes 4}\right)\right].$ (81)
+$\overline{Y^{\rm lin}}=\mathbb{E}_{U_{A}}\mathbb{E}_{U_{B}}\ Y^{\rm lin}\left(U_{A}U_{B}|\psi\rangle\langle\psi|\ U_{A}^{\dagger}U_{B}^{\dagger}\right)=1-\frac{1}{D}{\rm tr}\left[Q\mathbb{E}_{U_{A}}\mathbb{E}_{U_{B}}\left({U_{A}}^{\otimes 4}\otimes{U_{B}}^{\otimes 4}{|\psi\rangle\langle\psi|}^{\otimes 4}\ {U_{A}^{\dagger}}^{\otimes 4}\otimes{U_{B}^{\dagger}}^{\otimes 4}\right)\right].$ (103)
 Applying Eq. (96), we have:
-$\overline{Y}=\sum_{\sigma_{A},\pi_{A},\sigma_{B},\pi_{B}}\sum_{\vec{i},\vec{i^{\prime}}}2^{-4E}DV^{A}_{\sigma_{A},\pi_{A}}V^{B}_{\sigma_{B},\pi_{B}}{\rm tr}\left(Q_{A}T^{A}_{\pi_{A}}\right){\rm tr}\left(Q_{B}T^{B}_{\pi_{B}}\right){\rm tr}\left(T^{A-f_{A}}_{\sigma_{A}}{|\vec{i}\rangle\langle{\vec{i^{\prime}}}|}_{A}\right){\rm tr}\left(T^{B-f_{B}}_{\sigma_{B}}{|\vec{i}\rangle\langle{\vec{i^{\prime}}}|}_{B}\right).$ (82)
-The summation over $\vec{i},\vec{i^{\prime}}$ can be done similarly in Eq. (E.1):
-$F_{2}(\sigma_{A},\sigma_{B},E)$ ::=:= $\sum_{\vec{i},\vec{i^{\prime}}\in{\{0,1\}}^{\otimes 4E}}{\rm tr}\left(T^{A-f_{A}}_{\sigma_{A}}{|\vec{i}\rangle\langle{\vec{i^{\prime}}}|}_{A}\right){\rm tr}\left(T^{B-f_{B}}_{\sigma_{B}}{|\vec{i}\rangle\langle{\vec{i^{\prime}}}|}_{B}\right)$
+$1-\overline{Y^{\rm lin}}=\sum_{\sigma_{A},\pi_{A},\sigma_{B},\pi_{B}}\sum_{\vec{i},\vec{i^{\prime}}}2^{-4E}\frac{1}{D}V^{A}_{\sigma_{A},\pi_{A}}V^{B}_{\sigma_{B},\pi_{B}}{\rm tr}\left(Q_{A}T^{A}_{\pi_{A}}\right){\rm tr}\left(Q_{B}T^{B}_{\pi_{B}}\right){\rm tr}\left(T^{A-f_{A}}_{\sigma_{A}}{|\vec{i}\rangle\langle{\vec{i^{\prime}}}|}_{A}\right){\rm tr}\left(T^{B-f_{B}}_{\sigma_{B}}{|\vec{i}\rangle\langle{\vec{i^{\prime}}}|}_{B}\right).$ (104)
+The summation over $\vec{i},\vec{i^{\prime}}$ can be done similarly in Eq. (H.1):
+$F_{2}(\sigma_{A},\sigma_{B},E)$ $:=$ $\sum_{\vec{i},\vec{i^{\prime}}\in{\{0,1\}}^{\otimes 4E}}{\rm tr}\left(T^{A-f_{A}}_{\sigma_{A}}{|\vec{i}\rangle\langle{\vec{i^{\prime}}}|}_{A}\right){\rm tr}\left(T^{B-f_{B}}_{\sigma_{B}}{|\vec{i}\rangle\langle{\vec{i^{\prime}}}|}_{B}\right)$
 $=$ $\sum_{\vec{i}\in{\{0,1\}}^{\otimes 4E}}\delta_{\sigma_{A}(\vec{i}),\sigma_{B}(\vec{i})}=\left(\sum_{\vec{j}\in{\{0,1\}}^{\otimes 4}}\delta_{\sigma_{A}(\vec{j}),\sigma_{B}(\vec{j})}\right)^{E}.$
 Then we derive a summation formula Eq. (106), whose terms contains rational fractions, so the entire summation can be completed using symbolic computation.
-$\overline{Y}=\sum_{\sigma_{A},\pi_{A},\sigma_{B},\pi_{B}}2^{-4E}DV^{A}_{\sigma_{A},\pi_{A}}V^{B}_{\sigma_{B},\pi_{B}}{\rm tr}\left(Q_{A}T^{A}_{\pi_{A}}\right){\rm tr}\left(Q_{B}T^{B}_{\pi_{B}}\right)F_{2}(\sigma_{A},\sigma_{B},E).$ (84)
+$1-\overline{Y^{\rm lin}}=\sum_{\sigma_{A},\pi_{A},\sigma_{B},\pi_{B}}2^{-4E}\frac{1}{D}V^{A}_{\sigma_{A},\pi_{A}}V^{B}_{\sigma_{B},\pi_{B}}{\rm tr}\left(Q_{A}T^{A}_{\pi_{A}}\right){\rm tr}\left(Q_{B}T^{B}_{\pi_{B}}\right)F_{2}(\sigma_{A},\sigma_{B},E).$ (106)
 Let $D_{A}:=2^{|A|},D_{B}:=2^{|B|},D_{E}:=2^{E}$, we have:
-$\overline{Y}=\frac{4}{D_{A}D_{B}(D_{A}^{2}-9)(D_{B}^{2}-9)D_{E}^{3}}\big{(}D_{A}^{2}D_{B}^{2}D_{E}(D_{E}^{2}+3)-6D_{A}D_{B}(D_{A}+D_{B})(D_{E}^{2}+1)$ (85)
+$1-\overline{Y^{\rm lin}}=\frac{4}{D_{A}D_{B}(D_{A}^{2}-9)(D_{B}^{2}-9)D_{E}^{3}}\big(D_{A}^{2}D_{B}^{2}D_{E}(D_{E}^{2}+3)-6D_{A}D_{B}(D_{A}+D_{B})(D_{E}^{2}+1)$ (107)
 $-6(D_{A}^{2}+D_{B}^{2}-9)D_{E}(D_{E}^{2}-1)+3D_{A}D_{B}D_{E}(D_{E}^{2}+11)\big).$
-E.3 Eq. (5)
+H.3 Eq. (9)
 In this subsection, we present the analytical method to derive the exact expression of Eq. (9). The initial state is shown in Eq. (108).
@@
 The quantity in Eq. (9) then can be written as:
-$\overline{Y}$ $=$ $\mathbb{E}_{U_{A}}\mathbb{E}_{U_{B}}\ Y\left(U_{A}U_{B}|\psi\rangle\langle\psi|\ U_{A}^{\dagger}U_{B}^{\dagger}\right)$
-$=$ $D{\rm tr}\left[Q\mathbb{E}_{U_{A}}\mathbb{E}_{U_{B}}\left({U_{A}}^{\otimes 4}\otimes{U_{B}}^{\otimes 4}\otimes{I_{B}}^{\otimes 4}{|\psi\rangle\langle\psi|}^{\otimes 4}\ {U_{A}^{\dagger}}^{\otimes 4}\otimes{U_{B}^{\dagger}}^{\otimes 4}\otimes{I_{C}}^{\otimes 4}\right)\right].$
+$\overline{Y^{\rm lin}}$ $=$ $\mathbb{E}_{U_{A}}\mathbb{E}_{U_{B}}\ Y^{\rm lin}\left(U_{A}U_{B}|\psi\rangle\langle\psi|\ U_{A}^{\dagger}U_{B}^{\dagger}\right)$
+$=$ $1-\frac{1}{D}{\rm tr}\left[Q\mathbb{E}_{U_{A}}\mathbb{E}_{U_{B}}\left({U_{A}}^{\otimes 4}\otimes{U_{B}}^{\otimes 4}\otimes{I_{B}}^{\otimes 4}{|\psi\rangle\langle\psi|}^{\otimes 4}\ {U_{A}^{\dagger}}^{\otimes 4}\otimes{U_{B}^{\dagger}}^{\otimes 4}\otimes{I_{C}}^{\otimes 4}\right)\right].$
 Applying Eq. (96), we have:
-$\overline{Y}=2^{-4(b_{AB}+b_{AC}+b_{BC}+g)}D\sum_{\sigma_{A},\pi_{A},\sigma_{B},\pi_{B}}\sum_{\vec{i}_{AB},\vec{i}^{\prime}_{AB},\vec{i}_{AC},\vec{i}^{\prime}_{AC},\vec{i}_{BC},\vec{i}^{\prime}_{BC},\vec{i}_{g},\vec{i}^{\prime}_{g}}V^{A}_{\sigma_{A},\pi_{A}}V^{B}_{\sigma_{B},\pi_{B}}\times$ (89)
+$1-\overline{Y^{\rm lin}}=2^{-4(b_{AB}+b_{AC}+b_{BC}+g)}\frac{1}{D}\sum_{\sigma_{A},\pi_{A},\sigma_{B},\pi_{B}}\sum_{\vec{i}_{AB},\vec{i}^{\prime}_{AB},\vec{i}_{AC},\vec{i}^{\prime}_{AC},\vec{i}_{BC},\vec{i}^{\prime}_{BC},\vec{i}_{g},\vec{i}^{\prime}_{g}}V^{A}_{\sigma_{A},\pi_{A}}V^{B}_{\sigma_{B},\pi_{B}}\times$ (111)
 ${\rm tr}\left(T^{A-f_{A}}_{\sigma_{A}}{{|\vec{i}_{AB}\rangle\langle\vec{i}^{\prime}_{AB}|\otimes}{|\vec{i}_{AC}\rangle\langle\vec{i}^{\prime}_{AC}|}\otimes|\vec{i}_{g}\rangle\langle\vec{i}^{\prime}_{g}|}_{A}\right){\rm tr}\left(T^{B-f_{B}}_{\sigma_{B}}{{|\vec{i}_{AB}\rangle\langle\vec{i}^{\prime}_{AB}|\otimes}{|\vec{i}_{BC}\rangle\langle\vec{i}^{\prime}_{BC}|}\otimes|\vec{i}_{g}\rangle\langle\vec{i}^{\prime}_{g}|}_{B}\right)\times$
@@
 The summation over $\vec{i}_{AB},\vec{i}^{\prime}_{AB},\vec{i}_{AC},\vec{i}^{\prime}_{AC},\vec{i}_{BC},\vec{i}^{\prime}_{BC},\vec{i}_{g},\vec{i}^{\prime}_{g}$ can be done in the similar way as those Eq.
-(E.1) and Eq.
-(E.2).
+(H.1) and Eq.
+(H.2).
 $F_{3}(\sigma_{A},\sigma_{B},b_{AB},b_{AC},b_{BC},g)$
-::=:= $\sum_{\vec{i}_{AB},\vec{i}^{\prime}_{AB},\vec{i}_{AC},\vec{i}^{\prime}_{AC},\vec{i}_{BC},\vec{i}^{\prime}_{BC},\vec{i}_{g},\vec{i}^{\prime}_{g}}{\rm tr}\left(Q_{C}{{|\vec{i}_{AC}\rangle\langle\vec{i}^{\prime}_{AC}|\otimes}{|\vec{i}_{BC}\rangle\langle\vec{i}^{\prime}_{BC}|}\otimes|\vec{i}_{g}\rangle\langle\vec{i}^{\prime}_{g}|}_{A}\right)\times$
+$:=$ $\sum_{\vec{i}_{AB},\vec{i}^{\prime}_{AB},\vec{i}_{AC},\vec{i}^{\prime}_{AC},\vec{i}_{BC},\vec{i}^{\prime}_{BC},\vec{i}_{g},\vec{i}^{\prime}_{g}}{\rm tr}\left(Q_{C}{{|\vec{i}_{AC}\rangle\langle\vec{i}^{\prime}_{AC}|\otimes}{|\vec{i}_{BC}\rangle\langle\vec{i}^{\prime}_{BC}|}\otimes|\vec{i}_{g}\rangle\langle\vec{i}^{\prime}_{g}|}_{A}\right)\times$
 $\times$ ${\rm tr}\left(T^{A-f_{A}}_{\sigma_{A}}{{|\vec{i}_{AB}\rangle\langle\vec{i}^{\prime}_{AB}|\otimes}{|\vec{i}_{AC}\rangle\langle\vec{i}^{\prime}_{AC}|}\otimes|\vec{i}_{g}\rangle\langle\vec{i}^{\prime}_{g}|}_{A}\right){\rm tr}\left(T^{B-f_{B}}_{\sigma_{B}}{{|\vec{i}_{AB}\rangle\langle\vec{i}^{\prime}_{AB}|\otimes}{|\vec{i}_{BC}\rangle\langle\vec{i}^{\prime}_{BC}|}\otimes|\vec{i}_{g}\rangle\langle\vec{i}^{\prime}_{g}|}_{B}\right)$
@@
 Then we derive a summation formula Eq. (113), whose terms contains rational fractions, so the entire summation can be completed using symbolic computation.
-$\overline{Y}=2^{-4(b_{AB}+b_{AC}+b_{BC}+g)}D\sum_{\sigma_{A},\pi_{A},\sigma_{B},\pi_{B}}V^{A}_{\sigma_{A},\pi_{A}}V^{B}_{\sigma_{B},\pi_{B}}{\rm tr}\left(Q_{A}T^{A}_{\pi_{A}}\right){\rm tr}\left(Q_{B}T^{B}_{\pi_{B}}\right)F_{3}(\sigma_{A},\sigma_{B},b_{AB},b_{AC},b_{BC},g).$ (91)
+$1-\overline{Y^{\rm lin}}=2^{-4(b_{AB}+b_{AC}+b_{BC}+g)}\frac{1}{D}\sum_{\sigma_{A},\pi_{A},\sigma_{B},\pi_{B}}V^{A}_{\sigma_{A},\pi_{A}}V^{B}_{\sigma_{B},\pi_{B}}{\rm tr}\left(Q_{A}T^{A}_{\pi_{A}}\right){\rm tr}\left(Q_{B}T^{B}_{\pi_{B}}\right)F_{3}(\sigma_{A},\sigma_{B},b_{AB},b_{AC},b_{BC},g).$ (113)
 Let $D_{A}:=2^{|A|},D_{B}:=2^{|B|},D_{C}:=2^{|C|},D_{AB}:=2^{b_{AB}},D_{AC}:=2^{b_{AC}},D_{BC}:=2^{b_{BC}},D_{g}:=2^{g}$, we have:
-$\overline{Y}$ $=$ $\frac{4}{D_{A}D_{B}(D_{A}^{2}-9)(D_{B}^{2}-9)D_{AB}^{3}D_{AC}^{3}D_{BC}^{3}D_{g}^{3}}\big{(}D_{A}^{2}D_{B}^{2}D_{AB}^{3}D_{AC}^{2}D_{BC}^{2}D_{g}^{2}+3D_{A}^{2}D_{B}^{2}D_{AB}D_{AC}^{2}D_{BC}^{2}D_{g}$
+$1-\overline{Y^{\rm lin}}$ $=$ $\frac{4}{D_{A}D_{B}(D_{A}^{2}-9)(D_{B}^{2}-9)D_{AB}^{3}D_{AC}^{3}D_{BC}^{3}D_{g}^{3}}\big(D_{A}^{2}D_{B}^{2}D_{AB}^{3}D_{AC}^{2}D_{BC}^{2}D_{g}^{2}+3D_{A}^{2}D_{B}^{2}D_{AB}D_{AC}^{2}D_{BC}^{2}D_{g}$ (114)
 - $6D_{A}^{2}D_{AB}^{3}D_{AC}^{2}D_{BC}^{2}D_{g}^{2}-6D_{A}^{2}D_{AB}^{2}D_{AC}^{2}D_{B}D_{BC}D_{g}-18D_{A}^{2}D_{AB}D_{AC}^{2}D_{BC}^{2}D_{g}+24D_{A}^{2}D_{AB}D_{AC}^{2}$
@@
 - $144D_{AB}D_{AC}^{2}+24D_{AB}D_{B}^{2}D_{BC}^{2}-144D_{AB}D_{BC}^{2}+126D_{AB}+36D_{AC}^{2}D_{B}D_{BC}-36D_{B}D_{BC}\big).$
-Appendix F Simulation method
+Appendix I Simulation method
 In this section, we briefly discuss the numerical simulation method we used.
-We use two different methods for numerically evaluting the average $\overline{Y}$.
-The first one is a brute-force evaluation according to the definition of $Y$.
+We use two different methods for numerically evaluating the average $\overline{Y^{\rm lin}}$.
+The first one is a brute-force evaluation according to the definition of $Y^{\rm lin}$.
 However, this requires a direct enumeration of all $4^{N}$ Pauli strings, which can be time-consuming for larger system sizes.
@@
 (21), there are only four different values for ${{\alpha}_{{{P}_{A}}}}({{a}_{k}},{{a}_{k}},{{a}_{k}},{{a}_{k}})$, depending on whether $P_{A}=I_{A}$ and $a_{k}=I_{A}$.
-Therefore, we only need to evaluate the corresponding $\alpha$’s to efficiently compute the average $\overline{Y}$.
+Therefore, we only need to evaluate the corresponding $\alpha$’s to efficiently compute the average $\overline{Y^{\rm lin}}$.
 Similarly, there are four possible contributions from different $P$ in Eq.
-(20), so we can simply simulate these values to obtain an unbiased estimate for $\overline{Y}$.
+(24), so we can simply simulate these values to obtain an unbiased estimate for $\overline{Y^{\rm lin}}$.